A Complete Bibliography of Publications in Graphical Models

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org, beebe@ieee.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

25 April 2024 Version 2.40

Title word cross-reference

1 [CLL^{+}18, DLP13, GXT18]. 2 [BS04b, CYW04a, CYW04b, CL00, FGV^{+}14, GK04, KADS02, KS04, ME17, VPAM12, WTS17, YL14, ZC18, ZSC^{+}14]. 2.1 [YLW^{+}14]. 2.5 [FGV^{+}14]. 2008 [SCOG09]. 3 [BRC15, BIP00, BLT05, BAA18, BC15, BBML^{+}18, CGPZ21, CVB09, Che22, CK00, CCL22, CWS^{+}20, CL00, CMA14, CLG^{+}23, DXD14, DRD19, FWH13, FWC^{+}19, FW16, FWK20, FCSF16, GWYN19, GXT18, Gol13, GC22, GLXJ14, GCZZ18, HWQ14, HYZ^{+}19, KDCC23, LSA21, LL13, LQJ^{+}14, LJYL16, LZL16, LLW^{+}18, LGG19, LCZG14, LSY23b, MMC23, NKP11, NRW23, OBS05, PKL^{+}23, PS07, PLS^{+}18, PBN^{+}09, RG12, RNDA13, SGHM00, SS19, SSL17, SRML09, SWW^{+}14, TLGS05, TCH07, TCMS04, VRP22, VPAM12, WLT14, WLAT14, WMC^{+}22, WEY06, WTS17, WXZ^{+}16, YWH^{+}21, ZZZL13, ZXY^{+}12, ZZLZ13, ZZZY13, ZbQC^{+}19, ZLS^{+}20, ZZD^{+}23, tHV09]. 4 [Gol13, LXX^{+}19]. 5 [EE19]. ^{1} [BJM19]. A [ZRZ21]. C [BJM19]. C^{1} [BJM18, KP15b]. C^{2} [HFSB14]. ϵ [ME17]. G^{1} [BH14, BFRA12, FH12]. G^{2} [KP11]. H^{2} [TJ12]. kd [GK04]. L [ZLRL19]. L_{2} [ZW03]. λ [CL05]. μ [CW02]. n [RFLSA11, Thü03, UCB13]. T [ZPG18].

-axis [EE19]. -basis [CW02]. -Dimensional [UCB13, Thü03]. -dimensions [Gol13]. -Guarantee [ME17]. -Manifold [KADS02, GK04]. -medial [CL05]. -norm [ZW03]. -smooth [BJM19, BJM18]. -splines [ZPG18]. -system [ZLRL19].

-trees [GK04]. -weighting [ZRZ21].

05 [KSM^{+}06].

1-form [HXS09]. 10th [LV03].

2000 [BSW01]. 2002 [HCS03, Wyv03]. 2003 [Ano04g]. 2008 [Jam09]. 2014 [CDGH15]. 2019 [Ano19]. 2020 [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f]. 2021 [Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f]. 2022 [Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f]. 2023 [Ano23a, Ano23b, Ano23d, Ano23c, Ano23e, Ano23g]. 2024 [Ano24a, Ano24b]. 2D [CBK03].

3D [CBK03, DR03, NFU02, PS03, She03].

4 [FSF07]. 46-56 [BJM19]. 4D [GK03].

5-Axis [GC22].

74 [HQ12b].

8th [GHPW12].

’98 [WP00]. ’99 [KS00]. 9th [CCS23].

Absorption [Sbe00]. abstraction [FWC^{+}19, HD23]. accelerometers [RTKW15]. accessibility [SMKE14]. Accumulative [SSL17]. Accuracy [HS05]. Accurate [GWHH18, VM06, ZHMW22, MV23, PCS19, SZJ^{+}21]. Achieving [ZS09]. Acknowledgement [Ano02f, Ano03f, Ano04i, Ano05h, Ano06g, Ano07g, Ano15g]. Acknowledgment [Ano00b, Ano09h, Ano10f, Ano12e, Ano13e, Ano11f, Ano01f]. Acquiring [ZB05]. acquisition [PS03, TLGS05, ZLX^{+}18]. acting [PHE23]. Active [SO01]. Active-Space [SO01]. adaptation [TCL^{+}20]. Adaptive [FCG01, IM06, KGZ^{+}14, SK21, ZRZ21, ZGLP12, Bay19, HZW^{+}22, LDD14, MIPS14, OBS06b, XHH^{+}20]. Additive [LLL19a, GC22]. adjustment [NSL^{+}21, PT15]. advances [Jia16]. Adversarial [LWG^{+}21, JLW^{+}21]. aerial [ZN13]. aesthetic [KM20]. Affine [HJK02]. Agnostic [FH12]. Algebraic [AGH16, CCF01, AH16, ZL15]. Algorithm [LHS01, Bay19, Buz03, CLL^{+}18, KL19, KKH19, LLH22, Nie17, QHXC12, RFLSA11, SWYH18, TL05, ZL14]. Algorithms [KMP05, QY02, BC15, DR03, GS12, HZL^{+}20, VBN11, VPAM12]. align [CTL15]. aligned [LXZ14]. Alignment [CVB09, GCZZ18]. all-hex [HZW^{+}22]. ambient [BHS18]. AMONG [ZXZ21]. Analogs [LM00]. analogue [CSP20]. analyses [RMCdST19]. Analysis [Kim17, LS01, BBP18, CAF09, CH06, CN18, KL17, KL19, KZD^{+}11, LYKL12, MM19, MR05, SRML09, TJ12, WJG02, WXGZ20, XLX^{+}19]. analysis-suitable [KL19]. analytical [BDC^{+}19]. anatomical [LCZG14]. Angiograms [SGHM00]. angles [WG19]. Animal [FRDC06]. animated [MKS^{+}08]. Animating [GS01, BGA05]. Animation [Jam09, KB01, LK01, BDC^{+}19, DLP13, yKL11, OH06, PHE23, PY08, PKL^{+}23, Ros10, SJ12, UPBS08, WCHZ14, ZCCD06]. animations [FWK20, GVK06, TM07]. Anisotropic [ZSJG14, Bay19, LWZ^{+}18, PDA03]. annotation [MR05]. Announcement [Ano01d, Ano01c, Ano02a, Ano02b]. Anthropomorphic [TGB00]. antisymmetric [HSS15]. appearance [KSK23]. applicability [Tsc20]. Application [BVL02, AO03, BRC15, Che22, DB21, FWH13, KM20, TTF04, kWwZ13, YCKK17]. Applications [CCS23, EPB05, Elb05, HS05, MPVF11]. applied [LWH15]. Applying [KVS15]. Approach [LM00, LS01, dSNJMA22, SGHM00, AFBB17, AD15, CTL15, DXD14, Gus07, LLL19a, LBM04, LJYL16, MCQ05, OBS06a, RCVA11, SBSG23, SMT04, VCT09, VRP22, VN23, WLT14, WXN18, XLX^{+}19, dSGA15]. approaches [ZSL^{+}22]. Approximate [BS04a, HYYZ17]. Approximation [CXC14, CS18, CCS05, MK05, OBS05, VCT09, VM06, ZLAK14]. approximations [AACPMCMJ16, Par23, WHHB12]. April [Ano23a, Ano24a]. arbitrary [BBB11, DMMP03, LBM04, PSF07, UCB13, WBOL07, XTW16]. arc [HFSB14]. arc-length [HFSB14]. architecture [Che22]. arcs [BE11]. area [DGZ12, DB21, Hao20, ZN13]. art [CH06, CC24, CGW^{+}07]. art-directed [CGW^{+}07]. artifacts [BTCH05]. Artistic [RL13]. As-rigid-as-possible [WLL14, CGLX17]. assembling [RG12]. assessing [WXN18]. assessment [FFY^{+}20, HLM16]. atmospheric [LWGP08]. attention [DZSW23, HYZ^{+}19]. attributes [TRS06]. augmentation [LDLS23]. augmented [KP15b, MBH^{+}12, RW20, WS03]. Augmenting [HvBK19]. Author [Ano00a, Ano01a, Ano02g, Ano03e, Ano04h, Ano05g, Ano07f, Ano09a]. Authoring [ZLZ^{+}22, SKMM24]. autoencoder [HYZ^{+}19, ZFLL23]. autoencoder-attention [HYZ^{+}19]. Automated [SLD^{+}23]. Automatic [LCZG14, ZLW^{+}14, ZLS^{+}20, YGL^{+}18]. Autonomous [GVK06, ST07]. auxiliary [ZQ11]. avatar [LL06]. Award [Ano07e]. aware [CYF^{+}23, GMH^{+}20, JLW^{+}21, JSS^{+}14, LWG^{+}23, PHE23, ZFLL23]. axes [WG19]. Axis [CS01, GC22, BJK14, CL05, EE19, FMW^{+}22, LSY^{+}23a, ZC18, ZSC^{+}14].

b [HR23, CXY^{+}09, CL00, GSS00, HL21, KCD14, LLH22, MKH^{+}17, RG12, XWYY10, YZZ^{+}10, ZK05]. B-DNA [RG12]. b-rep [HR23]. B-Spline [CL00, CXY^{+}09, HL21, LLH22, MKH^{+}17, XWYY10, YZZ^{+}10]. B-Splines [GSS00, KCD14, ZK05]. background [VN23]. bag [DXD14]. Balanced [HSS15]. ball [KL14]. barks [VRP22]. Barycentric [ZSZ^{+}20]. Bas [NSL^{+}21, JSZ^{+}21, Nie17, WMR^{+}14, XLX^{+}19, ZMW^{+}14, ZZZY13]. Bas-relief [NSL^{+}21, JSZ^{+}21, XLX^{+}19, ZMW^{+}14, ZZZY13]. bas-reliefs [Nie17, WMR^{+}14]. base [KZW12]. Based [BIP00, BY01, Gle01, GY01, LHS01, ZXZ21, ALP06, BBF^{+}11, BC15, BBP18, BDC^{+}19, BMM^{+}07, BBB11, CMLH19, CGAY13, CBC^{+}07, CSJ13, CM20, CMA14, DRD19, DQ05, DQ04, DLP13, DB21, FML12, FMW^{+}22, FWX^{+}18, GLY^{+}17, GS12, GwGgP18, Gus07, HXS09, HYZ^{+}19, Hub12, JZLZ14, JXCZ13, KZD^{+}11, KK23, LJYL16, LXD19, LXX^{+}19, LXZ14, LWGP08, LYCG08, LV14, MMNG19, MDLH19, MBH^{+}12, MR05, MCQ05, MKB^{+}16, NSZ18, NBPF11, NRW23, Nie16, Nie17, NSL^{+}21, PT15, PFV^{+}11, PBM^{+}11, PD16, PBN^{+}09, QZZ^{+}23, RCVA11, RNDA13, SSHS14, SWHH15, SMT04, SS06, STRD19, TL05, TMT10, TM07, TCMS04, WJG02, WZL^{+}03, WLW06, WLT14, WXN18, WXGZ20, WHHB12, WZtW^{+}14, WTS17, kWwZ15, WLLL22, WLTS15, XXZL23, ZZZL13, ZMW^{+}14, ZZLZ13, ZL14, ZWSH14, ZCW^{+}22, ZLRL19, ZRZ21, ZSL^{+}22, ZSZ^{+}20]. Bases [BJM18, BJM19, GS12]. basic [KP12]. basis [CW02, FSF07, KXCD15, OBS06b, ZK05]. behavior [LL06]. behaviors [GVK06]. bending [TG13]. Bernstein [GS12]. best [ZW03]. between [CXY^{+}09, HKM12, TMT10]. Beyond [AT16]. Bézier [GS12, Hao20, HL22, ZW03, ZZ14]. bi [LLH22]. bi-cubic [LLH22]. Biharmonic [ABCV15, YGL^{+}18]. bilateral [yKL11]. BIM [LXTJ16]. binary [GK03, KWFH15, LWGP08, LDD14, NKP11, VBN11, XTW16]. biquadratic [KP15b]. Bisector [ALSR11, Pet00]. blend [BFRA12]. Blending [CCF01, ZQ11, Elb05, SSHS14]. blends [KP15a]. Blendshape [CO20a, CO20b]. blocks [RG12, RL13]. blossoming [GS12]. blue [MIPS14]. blur [ZWSH14]. blurring [BTCH05]. Board [Ano06a, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano09b, Ano10a, Ano10b, Ano05c, Ano05d, Ano06e, Ano10e, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b, Ano05e, Ano05f, Ano06b, Ano06c, Ano06d, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano14a, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano17a, Ano17b]. body [CN16, JLW^{+}21, KL13, LML^{+}20, LSY^{+}23a, LCZG14, SMT04, UPBS08]. body-centered [CN16]. body-part-aware [JLW^{+}21]. Boor [KL19, ZPG18]. Boor-like [KL19]. Boor-suitable [ZPG18]. bordered [PSF07]. Both [LI00]. Boundaries [LM00]. Boundary [SSP01b, HL22, HZW^{+}22, PS07, VPAM12]. Bounding [BE11, CMA14]. bounds [EE19]. box [Kim17, KK23]. box-spline [KK23]. box-splines [Kim17]. BPA [JLW^{+}21]. BPA-GAN [JLW^{+}21]. branch [GXT18, ZCW^{+}22]. branched [BKL15]. BRDF [GH03, GwGgP18, dSNJMA22]. Breathe [ZCCD06]. brush [XTLP04]. bubbles [ZYP09]. buffer [CLL^{+}18]. Building [SGS01, XTW16, RG12, WZtW^{+}14]. buildings [HLM16]. bump [RCHS18]. BVH [LSY^{+}23a].

C2 [Ano06a, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano09b, Ano10a, Ano10b]. CAD [GG18, KGZ^{+}14, KVS15]. Calibration [KC01]. Call [Ano01b]. Camera [CLG^{+}23, GY05, KC01, KSK23, RP08]. canonical [PLS^{+}18]. capture [GODC07, LML^{+}20, TCMS04, XYZ^{+}23]. Capturing [PCP02]. cardiovascular [LXX^{+}19]. Caricature [CGPZ21]. carrier [SDC04]. Cartoon [LYKL12, CRH05]. cartoon-style [CRH05]. CartoonModes [LYKL12]. categorization [SSL17]. Catmull [SS11, ZZS19]. centered [CN16]. centric [FFY^{+}20, HWQ14, HZL^{+}20]. centripetal [LYCG08]. Centroidal [AdVDI05]. CG [CGW^{+}07]. CGI [WP00]. character [ALP06, BDC^{+}19, CC24, CGW^{+}07]. characteristic [CM20]. characterization [BPG05]. Characterizations [Eva11]. characters [CBC^{+}07]. Christoph [Ano23f]. circle [ALSR11]. circles [BKL15, KKH19]. Cited [Ano07e]. CityGML [GPF19]. clamped [CXY^{+}09]. Clark [SS11, ZZS19]. Class [LFC^{+}18]. Class-sensitive [LFC^{+}18]. classical [ZSL^{+}22]. Classification [RFLSA11, BBML^{+}18, Che22, DXD14, KSS08, SSL17, STRD19]. clay [DC04]. clinical [LXX^{+}19]. Clipmaps [DR23]. cliques [Eva11]. cloning [FSF07, Pan03]. Closed [Thü03, AACPMCMJ16, HA03, KM20, SSHS14]. closely [LML^{+}20]. clothing [CGW^{+}07]. cloud [Che22, DZSW23, LWG^{+}23, MS09, ZML^{+}20, ZCW^{+}22, ZNYL22, ZSL^{+}22]. clouds [CWS^{+}20, DRD19, ES16, GF22, JXCZ13, LDLS23, MLF^{+}12, MB23, NBPF11, Nie16, Nie17, NSL^{+}21, PSL14, PLL12, SS19, STRD19, YCKK17, ZC18, ZN13]. Cluster [dSNJMA22, YXYW00]. clustering [JXC^{+}13]. Cluttered [RKK^{+}00]. coarsening [TCL^{+}20]. Coding [KG01, LB04]. coefficients [ZW03]. Coherence [LI00]. coherent [LYL10]. collections [SSL17]. collision [BHS18, CYF^{+}23, CWM^{+}14, EE19, LSY^{+}23a, TMT10, ZL15]. collision-aware [CYF^{+}23]. collision-free [BHS18]. colon [MMNG19]. Color [PT15, PS00, BS04b]. colour [Gra15]. combinatorial [CN16, CN18, FWWT13, BK03]. combined [NKP11]. Combining [BHS18, SZJ^{+}21, TCMS04, KP12]. commercial [BLT05]. common [KZW12]. Compact [FWWT13, Hub12, LLL19a, LL13, PCS19, ZSC^{+}14]. Compactly [SFB^{+}17, OBS05]. Compactly-supported [SFB^{+}17]. Comparative [FRDC06, BFRA12, FML12]. Comparing [Gle01, SKMM24]. Complete [ZN13, ACS03]. completion [ABCV15, LLL19a, LWG^{+}23, ZFLL23]. complex [MDLH19, RSFdM04]. complexes [CD11, FIF19]. Complexity [PCP02]. Component [ZFLL23, RCVA11]. Component-aware [ZFLL23]. components [ACH^{+}13]. Composing [FWK20]. composite [CWM^{+}14, OBS06a]. comprehensive [kWwZ13]. Compressing [IS02, IA03]. Compression [BIP00, SRK02b, LV14, MB23, NSL^{+}21, Vás11]. Compression-Based [BIP00]. Computation [KM00, SLB^{+}00, AACPMCMJ16, BWW^{+}14, CW02, CM20, MKH^{+}17, WXGZ20]. Computational [BB11, MYC17, HM13]. Computer [Jam09, LV03, AO03, DLP13, ZCCD06]. Computing [AFBB17, CXY^{+}09, DGZ12, FIF19, HKM12, MP03, WJG02, ZC18, ZSC^{+}14, JBK04, QHXC12, WMC^{+}22]. Concepts [DRD19]. condition [CCS05]. Conditions [CL00, HL22]. cones [BGZ16]. Conference [CCS23, LV03, SRK02a, Wyv03, MYC17]. conformal [LSA21, TH14, ZSJG14]. Connected [BBB11, NFU02, RCVA11]. connected-component-labeling-based [RCVA11]. Connectivity [KADS02, KG01, SDC04, ZS09]. Conservative [YG07]. consistent [LZY^{+}14, YWL^{+}19]. Constant [ACWK06]. Constant-volume [ACWK06]. Constrained [LBSP02, WZ14, YRZ18, ZW03]. Constraint [Gle01]. Constraint-Based [Gle01]. constraints [DQ05, HFSB14, KVS15, LB06]. Constructing [KZW12, TCH07]. Construction [HL22, SRK02b, BNR^{+}17, KS02, KGZ^{+}14, MK02]. Constructive [PASS01, SLF^{+}18]. contact [KEK14]. contacts [RTKW15]. Content [JSS^{+}14, Ano13f, FW16]. Content-aware [JSS^{+}14]. context [LGG19]. Continuous [CWM^{+}14, HQ12a, HQ12b, LM00, KEK14, MB23, RCG^{+}09, ZL15]. Contour [BGLSS04, BMM^{+}07]. Contour-based [BMM^{+}07]. Contours [KSS00, SGHM00]. contractible [NPJ14]. contraction [JXC^{+}13]. control [TKPR09, VSR12, ZZS19, ZML^{+}20, ZLZ^{+}22, ZCCD06]. controllable [CGLX17, YL08]. convergence [BE11, MM19]. Convex [EKH01, SR00, BKL15, NPJ14, SZJ^{+}21]. convinced [LLL^{+}19b]. Convolution [Hub12, SWHH15, ZSZ^{+}20]. convolution-based [SWHH15]. Convolutional [LZL16, GXT18, ZLS^{+}20, ZML^{+}20]. Coordinate [LS01, CN16]. Coordinate-Invariant [LS01]. coordinates [ZHM11]. core [GF22, LDD14, TMT10]. corrector [AD15]. correspondence [HLL^{+}23, ZLW^{+}14]. correspondences [BWW^{+}14, WMC^{+}22]. Corrigendum [BJM19, CYW04a, HQ12b]. cost [HLM16]. cost-effective [HLM16]. coupled [JXC^{+}13]. coupling [LWZ^{+}18]. covering [ME17]. CPU [NSL^{+}21]. crack [IO09]. Crease [PSK^{+}02]. create [BHS18]. creating [SBSG23]. Creation [Ros10]. creative [GLXJ14]. Creature [GLXJ14]. cross [LXZ14]. crowd [ACC14, SKMM24, YLG^{+}20]. crowds [BHS18]. Cubic [CL00, BE11, CN16, KK23, LLH22, WJG02, ZK05]. cubical [CM20]. cues [CRH05, CH06, KZW12]. culling [YG07]. Cumuliform [ZML^{+}20]. curl [SFCD21]. Curvature [PSK^{+}02, SSHS14, BJK14, EE19, JYTM14, KWFH15, LR12]. Curvature-based [SSHS14]. Curve [Elb01, JXC^{+}13, LXD19, CW02, VCT09, WWWM12]. Curves [EKH01, AACPMCMJ16, AH16, AMAS16, Far02, GS12, GC22, KM20, KEK14, LL13, LDZ^{+}17, ME17, MK02, SSHS14, Thü03, WJG02, WG15, XWYY10, YZZ^{+}10, ZZ14, tHV09]. Customization [FW16]. customized [FWK20]. cut [PSF07]. cut-graphs [PSF07]. cutaways [dCBM^{+}16]. Cuts [WZ14]. cutting [BGTG04, FML12]. CVM [MYC17]. cyclides [ZK15]. cylinder [AO03]. cylinders [BGZ16].

D [BAA18, CYW04a, BRC15, BIP00, BLT05, BC15, BBML^{+}18, BS04b, CGPZ21, CVB09, CYW04b, Che22, CK00, CLL^{+}18, CCL22, CWS^{+}20, CL00, CMA14, CLG^{+}23, DXD14, DRD19, DLP13, FWH13, FWC^{+}19, FW16, FWK20, FCSF16, FGV^{+}14, GWYN19, GXT18, GC22, GLXJ14, GCZZ18, HWQ14, HYZ^{+}19, KDCC23, KS04, LSA21, LL13, LQJ^{+}14, LJYL16, LXX^{+}19, LZL16, LLW^{+}18, LGG19, LCZG14, LSY23b, ME17, MMC23, NKP11, NRW23, OBS05, PKL^{+}23, PS07, PLS^{+}18, PBN^{+}09, RG12, RFLSA11, RNDA13, SGHM00, SS19, SSL17, SRML09, SWW^{+}14, TLGS05, TCH07, TCMS04, VRP22, VPAM12, WLT14, WLAT14, WMC^{+}22, WEY06, WTS17, WXZ^{+}16, YWH^{+}21, YL14, YLW^{+}14, ZZZL13, ZXY^{+}12, ZZLZ13, ZZZY13, ZbQC^{+}19, ZLS^{+}20, ZZD^{+}23, ZC18, ZSC^{+}14, tHV09]. D-objects [PS07]. D-rotations [RFLSA11].

Dagstuhl [CDGH15, GHPW12]. dashboard [Tsc20]. Data [GKR02, OK07, PS00, YLW^{+}19, ZLX^{+}18, BBF^{+}11, BLSZ20, FWWT13, GWHH18, HFSB14, KL14, KS04, KWFH15, KSS08, LL06, LB04, LDD14, LDLS23, MK05, MMC23, MKB^{+}16, OBS05, OBS06a, VS08, VRP22, YG07, YL14]. Data-driven [YLW^{+}19, MMC23]. Database [dSNJMA22]. dataset [MMC23]. datasets [Kim13, SBSG23]. December [Ano23b]. decimation [GMH^{+}20]. decision [DXD14]. Decomposing [DMMP03]. Decomposition [RFLSA11, BBML^{+}18, GWYN19, JBK04, JSL^{+}20, MB23, TMT10, WSC^{+}12]. decompression [PD16]. decorative [Elb05]. Deep [CMLH19, HLL^{+}23, WMC^{+}22, BLSZ20, Che22, ZNYL22, ZSL^{+}22]. DeepPipes [CWS^{+}20]. Defeating [PB20]. defect [CGAY13]. defect-laden [CGAY13]. defined [AWC06, DR03]. Deformable [BSB14, LKE00, BNR^{+}17, CTL15, CBC^{+}07, DQ04, KMBG09, NÇ10, SBA13, TMT10, XYZ^{+}23]. Deformation [TH14, AWC06, BDC^{+}19, CGLX17, LB06, NSZ18, SWC^{+}21, YGL^{+}18, YLW^{+}19]. deformations [DC04, LF04]. deforming [SOG09, WLTS15]. degree [ZW03]. Delaunay [GYH13]. deleting [Eva11]. denoising [LLL^{+}19b, PCPM23, TH12, XHH^{+}20, ZSL^{+}22]. dense [ZN13]. Dependant [GKR02]. dependency [VBN11]. deposition [ZLRL19]. depth [CMLH19, LLL19a, SRML09]. descent [CO20b]. description [CN18]. descriptor [HLX^{+}22, MPVF11]. design [DQ05, FCSF16, LZX^{+}15, MMC23, SWC^{+}21, TTF04]. design-preserving [SWC^{+}21]. Detail [EB17, TKPR09, KSH18, RCG^{+}09, YXF14, MB23]. Detail-preserving [EB17, TKPR09]. Detailed [SJ12]. Detection [CGPZ21, PSK^{+}02, CLL^{+}18, CWM^{+}14, FMW^{+}22, JZLZ14, JXCZ13, LJYL16, LSY^{+}23a, SZJ^{+}21, TMT10, Wu02, ZL15]. Determination [NÇ10]. determines [ZGLG12]. Deterministic [AO03]. development [ACS03]. device [CTAO20]. devices [Tsc20, WCHZ14]. diagram [CS18, KZW12]. diagrams [AdVDI05, QZZ^{+}23]. diameter [RNDA13]. diameter-based [RNDA13]. diamond [CN18]. difference [ZXY^{+}12]. difference-of-Gaussian [ZXY^{+}12]. different [LSY23b, PBN^{+}09]. differentiable [XXZL23]. Differential [JYTM14, LXD19]. differentiation [GS12]. Digital [LM00, ML00, SR00, SDC04, AACPMCMJ16, Buz03, Eva11]. DigitalSculpture [MCQ05]. Dimension [CD11, And03, BBB11]. Dimension-independent [CD11]. Dimensional [SGS01, UCB13, BSMG05, BWW^{+}14, GWHH18, KCM20, MS09, Thü03]. dimensionality [BLSZ20]. dimensions [BJM18, BJM19, DMMP03, Gol13]. Direct [Att14, JK02, WMC^{+}22]. directed [CGW^{+}07, SZJ^{+}21]. Direction [VBN11, KK23]. Direction-dependency [VBN11]. directional [GWYN19, SLKL11, WS03]. Dirichlet [WZ14]. discoloration [XL23]. discontinuous [KMBG09]. Discrete [And03, LV03, TG13, ZGLG12, ALSR11, CSP20, FIF19, HQ12a, HQ12b, Thü03, VCT09, ZGZ^{+}14]. Discretization [CBK03, BTCH05]. diseases [XL23]. Displaced [JK02]. Display [PS00]. dissemination [Ros10]. dissimilarity [LFC^{+}18]. dissipation [HKL^{+}15]. Distance [CS01, SBA13, BS04a, BBB11, CYW04a, CYW04b, CXY^{+}09, CLG^{+}23, ES16, HKM12, MS10, QHXC12, SWHH15, WXZ^{+}16]. distance-based [BBB11]. distances [MS09]. Distortion [HJK02, NSZ18, SFCD21]. distributed [PB20, RW20]. distribution [ZLX^{+}18]. distributions [MS09]. DNA [RG12]. domain [KZW12, SWYH18, WXRA07]. domains [BJM18, BJM19, ME17, WBOL07]. dominant [PSL14, Wu02]. drawing [WTS17]. drawings [LQJ^{+}14, ZXY^{+}12]. dressed [MMC23]. driven [CYW04a, CYW04b, MMC23, VRP22, YLW^{+}19]. drone [LXH21]. DS [ZPG18]. Dual [Tau02a]. due [BTCH05]. Dupin [ZK15]. Dynamic [DQ05, GZL^{+}20, SGHM00, ALP06, CH06, GLY^{+}17, GY05, MKB^{+}16, SFCD21, VCT09, Vás11, WZL^{+}03, Wu02, XHH^{+}20, ZLX^{+}18]. dynamics [DB21].

ear [SWW^{+}14]. easy [ZCCD06]. eddy [WXN18]. edges [Eva11, NPJ14]. Editing [CO20a, Gle01, ACC14, CO20b, EB17, TPG^{+}23, WXRA07, ZHM11]. Editorial [Ano01d, Ano01c, Ano05c, Ano05d, Ano06a, Ano06e, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano09b, Ano10a, Ano10b, Ano10e, BAA18, CCS23, MYC17, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b, Ano05e, Ano05f, Ano06b, Ano06c, Ano06d, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano14a, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano17a, Ano17b]. Effect [KC01]. effective [HLM16, PSL14]. effects [KMP05, yKL11]. efficiency [HLM16]. Efficient [BDC^{+}19, GK04, KG01, LJYL16, LSY^{+}23a, SLF^{+}18, SLB^{+}00, VPAM12, ZK08, ZXY^{+}12, ZLH13, ZGX^{+}18, GLY^{+}17, SWC^{+}21, XXZL23, ZCS^{+}15, ZCW^{+}22]. eigenfunction [LXZ14]. eigenfunction-based [LXZ14]. elasticity [NÇ10, PDA03]. element [NÇ10, WLTS15]. elements [BPG05, ITF06, MV23]. Elimination [BTCH05]. ellipse [SZJ^{+}21]. Ellipsoidal [DR23]. elliptical [KCM20]. embedding [ZSJG14]. EMD [HWQ14]. emotion [MDLH19]. Empirical [WSC^{+}12]. enabled [HLM16]. Encapsulating [GSS00]. enclosing [KL14]. Encoding [KADS02, CMLH19, CYF^{+}23, FMW^{+}22]. end [Che22]. end-to-end [Che22]. energy [HLM16, YLW^{+}14]. engineering [KVS15, WTS17]. Enhanced [KS02, LQJ^{+}14]. enhancement [VN23, YXF14]. Enhancing [LI00]. enrichment [STRD19]. environment [LHM06]. environments [MV23]. equations [LWH15]. Equiareal [YZC14]. Error [BVL02, GMH^{+}20]. Escher [AT16]. estimated [CLG^{+}23]. Estimating [QY02, RKH05]. Estimation [PSK^{+}02, WS03, LXZ14, LGG19, RP08, RNDA13, ZNYL22, ZZD^{+}23]. Euclidean [AT16, Che22]. Euler [CM20]. evacuation [MDLH19]. Evaluating [SKMM24]. Evaluation [Elb01, KL19, KM20, Kim17, PLS^{+}18, RW20, SLF^{+}18, Tsc20]. Exact [Far02, DLP13, ES16, ZLW^{+}14]. Example [CMB^{+}12, SMT04]. example-based [SMT04]. Example-guided [CMB^{+}12]. examples [HXS09]. exhaustive [ALSR11]. explicit [SOG09]. Exploration [MJ16]. ExploreTree [YWZB17]. exploring [NF06]. Explosion [BY01]. Exponent [ZK01]. expressive [NF06]. Extending [Sta05, ZK05, ZK01]. Extract [LDZ^{+}17]. Extracting [Nie16]. Extraction [BGZ16, FWH13, GWHH18, GK04, HXS09, JXC^{+}13, LDD14, PLL12, VPAM12, WWWM12, YLW^{+}14].

fabrication [YCKK17]. FABRIK [AL11]. facade [JYD^{+}16, WZtW^{+}14]. Face [CGPZ21, FWH13, YWH^{+}21, tHV09]. Faces [RKK^{+}00]. Facial [FSF07, Pan03, PHE23, WCHZ14, tHV09]. fairing [BJK14, LBM04]. Families [PSF07]. Fast [AACPMCMJ16, ES16, MIPS14, Par23, AL11, HYYZ17, PD16, SZJ^{+}21, ZL15]. fat [BE11]. FBA [FSF07]. FCC [Kim13]. Feature [GYH13, LLL^{+}19b, CMLH19, FWH13, HWQ14, KS04, LDZ^{+}17, LDLS23, Nie16, PLL12, WHHB12, XHH^{+}20]. feature-centric [HWQ14]. Feature-convinced [LLL^{+}19b]. Feature-preserving [GYH13, XHH^{+}20]. Features [BMZB02, DIOV06, WLT14, ZNYL22]. February [Ano24b]. FEM [KMBG09]. Few [SGHM00, RTKW15]. fewer [OK07]. fidelity [LWG^{+}23]. field [CAF09, CYW04a, CYW04b, GY05, LB04, LXZ14, WLAT14, WXZ^{+}16]. fields [ABCV15, BS04a, BK03, QHXC12, TCMS04]. Fifth [MYC17]. filleting [Elb05]. films [kWwZ15]. filter [yKL11, Tsc20, ZL15]. Filtering [SO01]. filters [HSS15]. Finite [WBOL07, WLTS15, ITF06]. Fisher [Che22]. fitting [LXD19, MK02, WY11]. flame [ZLX^{+}18]. Flexible [BH14, KMBG09, DLP13]. floorplans [SLD^{+}23]. flow [LWZ^{+}18, MKS^{+}08, TCMS04, WBOL07, YWL^{+}19, ZGZ^{+}14]. fluid [LWZ^{+}18, LWGP08, TKPR09]. fluid-structure [LWZ^{+}18]. fluids [GZL^{+}20]. FoldedGI [CLL^{+}18]. folding [CLL^{+}18]. footage [MR05]. forces [TG13]. forests [DXD14]. Foreword [Ano16g]. form [AGCA06, GC22, HXS09, KP12, KMP05, SMKE14, TL05]. formation [ZML^{+}20, ZZC^{+}21]. forms [KWFH15, PLS^{+}18]. Formulas [GS12]. Fourier [WEY06]. Fourier-interpolated [WEY06]. Fragmented [YLL12, ZL14]. frame [CCJ^{+}18]. frames [Far02]. framework [BGA05, BK03, MDLH19, TH12, TPG^{+}23, VCT09, WXRA07, YLW^{+}14, tHV09]. Free [KP12, AGCA06, BHS18, EE19, GWYN19, GC22, KMP05, SMKE14, TL05]. Free-form [KP12, AGCA06, GC22, KMP05, TL05]. freeform [KEK14]. friction [AD15]. front [TMT10]. front-based [TMT10]. fruit [XL23]. Full [LML^{+}20, UPBS08]. Full-body [LML^{+}20, UPBS08]. function [PFV^{+}11, RKH05, TTF04, TPFA21]. function-based [PFV^{+}11]. Functional [HLL^{+}23, Elb05, Hao20, KZD^{+}11, WMC^{+}22, ZQ11]. Functions [CL00, ZK01, CFG06, FSF07, OBS06b, SWHH15]. fundamental [KWFH15]. fungus [XL23]. fused [ZLRL19]. fusion [HLX^{+}22, LXH21, ZZZL13]. Futurist [CH06]. Fuzzy [GSS00, NFU02]. Fuzzy-connected [NFU02].

gaits [FRDC06]. Galerkin [KMBG09]. GAN [JLW^{+}21]. GANs [CC24]. garment [CYF^{+}23, PKL^{+}23, SWC^{+}21]. gas [GLY^{+}17]. GAT [DZSW23]. Gaussian [KWFH15, TL05, ZXY^{+}12]. GBGVD [QZZ^{+}23]. General [HKZM20, ZSZ^{+}20]. generalised [HD23]. Generalization [HMESI13, PPP19]. Generalized [ALSR11, Elb05, AO03]. generate [VRP22]. generated [BHS18, WG15]. Generating [IO09, LLW^{+}18, PKL^{+}23]. Generation [ZXZ21, ZLRL19, Bul23, CC24, HR23, HL21, Nie17, NSL^{+}21, SLD^{+}23, WXN18, ZMW^{+}14, ZZLZ13, ZZZY13, ZbQC^{+}19]. generative [JLW^{+}21, ZFLL23]. generators [CS18]. Genetic [QY02]. genus [PSF07]. geodesic [HYYZ17, QZZ^{+}23, QHXC12]. Geometric [CDGH15, EKK15, GHPW12, LXH21, Pet00, SSP01a, BBP18, CMLH19, DQ05, EB17, KEK14, KVS15, MS09, MMS^{+}07, PBN^{+}09, XWZ^{+}15, YXF14, ZRZ21]. geometrical [kWwZ13]. Geometrically [DLP13]. Geometry [BRC15, BVL02, LV03, LL13, NSZ18, ZXZ21, Bay19, CLL^{+}18, JYTM14, SLF^{+}18, WXN18]. Geometry-Based [ZXZ21, NSZ18]. gesture [AWC06]. gestures [ZB05].

GIM3D [MMC23]. Global

[BVL02, DC04, GWYN19, WLT14, WSZL13, YLW^{+}14, ZLZ^{+}22]. global-optimal [GWYN19]. global-to-local [ZLZ^{+}22]. Globally [LZY^{+}14]. GMOD [Ros10]. good [BO05, ZS09]. GPGPU [HZL^{+}20]. GPU [BHS18, HKM12, Kim13, MKB^{+}16]. GPU-based [MKB^{+}16]. GPU-generated [BHS18]. GPUs [Kim17]. gradient [HvBK19, WXRA07, ZLH13]. gradients [BS04a]. grammar [GLXJ14]. Graph [NBPF11, RMCdST19, ZCW^{+}22, DZSW23, HA03, HFH16, JXC^{+}13, MLF^{+}12, PKL^{+}23, SZJ^{+}21, YRZ18, ZL14, ZCW^{+}22]. Graph-based [NBPF11, ZCW^{+}22, ZL14]. Graph-PBN [ZCW^{+}22]. graphic [BS04b, JFS11]. Graphical [BJM19, CYW04a, HQ12b, KB01, Ros10, YL14, KS00]. Graphics [Ano04g, BSW01, CCS23, HB05, HCS03, KS00, SRK02a, HLM16, RC06]. graphs [PSF07, ZS09]. greedy [Bay19]. Gregory [FH12]. grid [CN16, CN18, CM20, HD23, IM06, KCM20, MS10, SLKL11]. ground [LHM06, RTKW15]. grouping [AT16]. groups [AT16, Kim17]. growing [Bay19]. Growth [QZZ^{+}23, XL23]. Growth-based [QZZ^{+}23]. GSNet [PKL^{+}23]. Guarantee [ME17]. guidance [LXZ14]. guided [CMB^{+}12, GCZZ18, HFH16, QSS^{+}19, WXZ^{+}16, ZDL^{+}11].

Hair [PCP02, YXYW00, CCL22, JSL^{+}20]. Hairstyle [LK01]. hairstyles [CCL22]. hairy [FJP06, XTLP04]. hand [CLG^{+}23, XYZ^{+}23]. Harmonic [HXS09, ZLS^{+}20]. Hausdorff [CS01, HKM12]. Head [SGS01]. heat [GLY^{+}17, YLL12, ZGLG12]. heat-based [GLY^{+}17]. height [LXH21]. helps [CLG^{+}23]. heptadiagonal [JSL^{+}20]. Heterogeneous [HL21, TPFA21]. hex [HZW^{+}22]. Hexagonal [SWYH18]. hexahedral [ITF06, IA03]. Hierarchical [KCD14, LV14, LKE00, DCL^{+}08, HZL^{+}20, JZLZ14]. hierarchies [JR09]. High [DR23, LWG^{+}23, GZL^{+}20, KZD^{+}11, KSS08, TPG^{+}23]. High-fidelity [LWG^{+}23]. High-performance [DR23]. high-precision [KSS08]. high-resolution [KZD^{+}11, TPG^{+}23]. high-speed [GZL^{+}20]. highly [CLL^{+}18, DLP13]. HMDO [XYZ^{+}23]. HMMs [WZL^{+}03]. hodograph [Far02, HFSB14]. Hoffmann [Ano23f]. holes [QHXC12]. homogeneous [PCPM23, VN23]. Hull [EKH01, SZJ^{+}21]. Human [FFY^{+}20, JLW^{+}21, KB01, RKK^{+}00, DXD14, LL06, SMT04, TCMS04, ZZD^{+}23, ZCCD06]. Human-centric [FFY^{+}20]. humans [MMC23]. Hybrid [LXX^{+}19, TPFA21, HLX^{+}22, LSY^{+}23a, TCL^{+}20, ZFLL23]. HybridTree [AGCA06]. Hyper [ZXZ21]. Hyper-Relations [ZXZ21]. Hyperbolic [CS01, XWYY10, ZK05]. Hypervolume [PASS01].

i.e [FGV^{+}14]. ICP [SK21]. II [NPJ14]. Image [GCZZ18, LI00, LWG^{+}21, BBF^{+}11, BTCH05, CLL^{+}18, CC24, CLG^{+}23, DR03, GWHH18, HR23, KS02, LV03, NRW23, NFU02, PT15, PS03, TL05, WLW06, ZMW^{+}14, ZL14]. image-based [PT15, WLW06]. Image-guided [GCZZ18]. Image-Space [LI00]. image-to-image [CC24]. Imagery [LV03]. Images [GW01, HKZM20, BSMG05, Bay19, BB11, CFG06, GK03, LXX^{+}19, LGG19, NKP11, PBM^{+}11, RKH05, TCH07, VN23, WS03, YLW^{+}14]. imaging [KZD^{+}11, NRW23]. immersive [TCMS04, TPG^{+}23]. Implicit [GA00, SK01, AGCA06, BGA05, BMM^{+}07, LSY23b, MCQ05, TRS06, VCT09, ZQ11]. Implicitly [CYF^{+}23]. Implicitly-proxied [CYF^{+}23]. ImplicitPCA [CYF^{+}23]. Improved [HWQ14, KL14, KP15a, GPF19, KKH19, LM12, MM19, SBSG23]. incenter [HMESI13]. Incremental [GA00, Buz03, KS02, SLF^{+}18]. independent [AFSW03, CD11]. Index [Ano00a, Ano01a, Ano02g, Ano03e, Ano04h, Ano05g, Ano07f, Ano09a]. Indexing [SO01]. Indoor [GW01, FFY^{+}20, GF22, LGG19, STRD19, TPG^{+}23]. inertial [SBSG23]. infinite [Hub12]. information [VN23]. informations [PS03]. Injectivity [CL00]. inpainting [WLLL22]. insect [GODC07]. inspired [FCSF16]. integer [HD23]. integrated [KSS08]. intent [YWZB17]. interacting [LML^{+}20]. Interaction [MKS^{+}08, CL19, DXD14, GLY^{+}17, JFS11, YWL^{+}19]. Interactive [DC04, KSS08, LB06, LZX^{+}15, YWZB17, dCBM^{+}16, AD15, GVK06, KS02, LF04, MCQ05, NFU02]. interest [FWC^{+}19]. interface [LF04]. interference [CLL^{+}18, kWwZ15]. interior [HFH16, VSR12]. International [GP06, LV03, KB01, Wyv03]. interpolated [WEY06]. Interpolating [KS04]. Interpolation [GKR02, BGLSS04, BH14, EKK15, HFSB14, LLXW13, LLH22, OBS05, VSR12, XZWB06]. interpolators [SFB^{+}17]. interpolatory [BFRA12]. Interrogation [Elb01, HS05]. intersecting [Att14]. Intersection [HHS^{+}01]. Intersections [SDC04, JBK04, WJG02, ZZ14]. Interval [SSP01a]. Intrinsic [YRZ18, CYW04a, CYW04b, JXCZ13, QHXC12]. Invariant [LS01, AGH16, MPVF11]. Inverse [AL11, TGB00, UPBS08, OH06]. invertible [ITF06]. investigation [Tsc20]. irregular [VCT09]. Iso [LM00, LDD14]. iso-surface [LDD14]. Iso-Surfaces [LM00]. Isogeometric [NPJ14, BJM18, BJM19, MM19, TJ12, XLX^{+}19]. Isomap [YL14]. isometric [BWW^{+}14]. isometries [LSA21]. isosurface [Kim13]. isosurfaces [GK04, VS08]. isotopic [CCS05]. isotropic [AdVDI05]. Issue [Ano01b, BSW01, CNK01, CDGH15, GHPW12, HB05, KS00, KCOTW06, SCOG09, SRK02a, Tau02b, WP00, Wyv03, Ano06f, BAA18, CCS23, GP06, HCS03, HM13, Jia16, KSM^{+}06, MYC17, PS05]. Isthmus [BC15]. iteration [HKM12, NKP11]. iteration-by-iteration [NKP11]. iterative [AD15, AL11, Nie16].

Jacobi [LLH22]. Jacobi-PIA [LLH22]. Jacobian [CO20b]. Jacobians [TG13]. January [Ano20a, Ano21a, Ano22a, Ano23c]. Jittor [XXZL23]. Joint [LDLS23, KSK23]. Jrender [XXZL23]. July [Ano20b, Ano21b, Ano22b, Ano23d].

kernel [WZL^{+}03, ZGLG12]. kernel-based [WZL^{+}03]. kernels [Hub12, YLL12]. key [YGL^{+}18]. keyframe [TM07]. Kinematics [AL11, TGB00, OH06, UPBS08]. kinetic [AD15]. Knitted [KAD^{+}21].

labeled [MMC23]. labeling [CMLH19, GCZZ18, RCVA11]. labelling [LCZG14]. laden [CGAY13]. lag [PB20]. Laguerre [CS18]. Landmark [CGPZ21, ZLW^{+}14]. landmarks [LCZG14]. lane [KS02]. Laplacian [LV14, ZHM11]. Laplacian-based [LV14]. Large [PSK^{+}02, RW20, Tau02b, WXN18, GF22, KL14, LXTJ16]. Large-eddy [WXN18]. Large-scale [RW20, GF22, LXTJ16]. laser [SRML09]. Layer [GY01]. Layer-Based [GY01]. Layered [QSS^{+}19, GH03]. Layout [ZXZ21, BBM^{+}17]. Layouts [HKZM20, JYD^{+}16]. leaf [QSS^{+}19]. Learning [CWS^{+}20, FWC^{+}19, NRW23, SWC^{+}21, WZL^{+}03, YWH^{+}21, CSJ13, Che22, HLX^{+}22, LXH21, LDLS23, LSY23b, PHE23, YWZB17, ZCW^{+}22, ZNYL22, ZSL^{+}22]. Learning-based [NRW23, ZSL^{+}22]. Learnt [OH06]. leaves [SFCD21]. Legendre [SLB^{+}00]. length [HFSB14]. level [CCL22, EB17, HLX^{+}22, RCG^{+}09, WXN18, ZHMW22, MB23]. level-of-detail [RCG^{+}09]. Levels [HD23]. library [PBN^{+}09, XXZL23]. LiDAR [ZN13]. light [CAF09, LB04, NRW23, WS03]. light-field [CAF09]. Lightweighting [LXTJ16]. like [BNR^{+}17, KL19, ZLRL19]. limb [ZB05]. Limbs [TGB00]. Line [SMKE14, ZZLZ13, CAF09, LQJ^{+}14, WHHB12, ZXY^{+}12]. Line-based [ZZLZ13]. line-space [CAF09]. linear [And03, BHS18, BS04b, Buz03, HD23, LXD19, PDA03]. lines [Buz03, FCSF16, JYTM14, Nie16, RL13]. Liquid [CL19, GLY^{+}17]. Liquid-solid [CL19]. LIR [LWH15]. List [Ano02f, Ano03f, Ano04i, Ano05h, Ano06g, Ano07g, Ano15g]. live [ZB05]. local [DC04, MPVF11, SWW^{+}14, WY11, WLT14, WHHB12, ZLZ^{+}22]. Localized [CO20a, CO20b]. Locating [RKK^{+}00]. location [LM12]. log [Che22]. log-Euclidean [Che22]. logging [MR05]. looking [ME17, RP08]. loop [SS11]. lossless [CMA14]. lossy [MB23]. low [BWW^{+}14, LWG^{+}23, WLLL22]. low-dimensional [BWW^{+}14]. low-resolution [LWG^{+}23]. Lumen [SGHM00]. lumped [NÇ10].

M [Ano23f]. machine [BGTG04]. made [JWL12]. magnetic [KZD^{+}11]. making [WMR^{+}14]. man [JWL12]. man-made [JWL12]. management [HLM16]. Manifold [GHQ06, Gus07, KADS02, ACS03, DMMP03, GK04, SWW^{+}14]. Manifold-based [Gus07]. manifolds [HQ12a, HQ12b]. manipulation [JSZ^{+}21, SMT04, XYZ^{+}23]. Manipulators [CO20a]. manufacturing [GC22]. map [KS02, WXGZ20, YZ20]. Mapping [BIP00, IS02, KCM20, AFSW03, CMLH19, JSS^{+}14, LXH21, MLC^{+}17, RCHS18]. mapping-independent [AFSW03]. maps [DR03, FWWT13, KS04, LLL19a, PT15, PY08, WMC^{+}22, HLL^{+}23]. March [Ano20c, Ano21c, Ano22c]. marching [OK07]. Markerless [XYZ^{+}23]. Masking [RCG^{+}09]. masks [HR23]. Mason [WR05]. mass [JSL^{+}20, MLC^{+}17, MMNG19, NÇ10, dSGA15]. mass-spring [NÇ10, dSGA15]. massive [DRD19, ES16, BAA18]. MAT [BRC15]. matching [BSB14, DIOV06, FWH13, HFH16, KMP05, WTS17, tHV09]. material [LWZ^{+}18, MJ16]. matrices [EKK15]. matrix [AFBB17]. matting [LYL10]. May [Ano20d, Ano21d, Ano22d, Ano23e]. MCCD [TMT10]. means [CFG06]. measure [BBML^{+}18, ZRZ21]. measurement [TH14]. measures [DGZ12, NSZ18]. mechanical [DLP13]. Media [MYC17, HM13]. Medial [CS01, BJK14, CL05, LSY^{+}23a, ZC18, ZSC^{+}14]. merge [DR03]. Mesh [BVL02, JZLZ14, KG01, Tau02a, WSZL13, ACS03, ACH^{+}13, ABCV15, BPG05, FML12, FWWT13, FWX^{+}18, GYH13, GXT18, HDL16, LLL^{+}19b, LLXW13, LLW^{+}18, LVM04, MBH^{+}12, MPVF11, TCL^{+}20, TH12, Vás11, WY11, WLAT14, WLLL22, XHH^{+}20, YLW^{+}19, ZHM11, ZZLZ13, ZZZY13]. Meshes [IS02, KADS02, LBSP02, PSK^{+}02, SRK02b, Tau02b, AGCA06, Att14, BGTG04, BH14, DCL^{+}08, GMH^{+}20, HvBK19, HZW^{+}22, IA03, KDCC23, LBM04, LDZ^{+}17, LV14, MM19, PSF07, QHXC12, RCG^{+}09, SS19, She03, SBA13, SLF^{+}18, UCB13]. meshing [BO05, OBS06a, PSL14, ZSJG14]. Meshless [MMS^{+}07, SOG09]. meta [FW16]. meta-scenes [FW16]. metamorphosis [CYW04a, CYW04b, WXZ^{+}16]. Method [LBSP02, SO01, SLB^{+}00, CMA14, Eva11, PSL14, WZtW^{+}14, Wu02, YL14, ZCS^{+}15, ZNYL22]. Methods [Gle01, NF06, KZD^{+}11, MK05]. metric [FWX^{+}18, GMH^{+}20, Gra15, LFC^{+}18, ZGLG12]. metrics [BBM^{+}17, FFY^{+}20]. Mexican [HQ12b, HQ12a]. micro [Tsc20]. micro-filter [Tsc20]. Microfacet [kWwZ15, GwGgP18]. Microfacet-based [kWwZ15]. microstructures [PFV^{+}11]. mildew [XL23]. Minimal [GK03, ACS03, BGZ16, HL21, RL13, SDC04]. minimising [BKC15]. Minimizing [HJK02, Far02]. minimum [CXY^{+}09, KL14]. Minkowski [MKH^{+}17, MP03, PS07, VM06]. mirroring [TLGS05]. Mix [LSY23b]. Mixing [AGCA06]. MixNet [LSY23b]. mixture [WZ14]. mixtures [LWGP08]. MLS [WHHB12]. mobile [Tsc20, WCHZ14]. modal [LYKL12]. mode [WSC^{+}12]. Model [CGPZ21, GSS00, SSP01b, YXYW00, ZCCD06, And03, AO03, BS04b, DQ04, FGV^{+}14, GLY^{+}17, GZL^{+}20, GWHH18, GH03, GCZZ18, GwGgP18, KGZ^{+}14, LSA21, LWGP08, LYCG08, LGG19, QSS^{+}19, TCH07, WLT14, WZtW^{+}14, WTS17, ZWSH14]. ModelCamera [PBSM06]. modelers [ACS03]. Modeling [Ano01b, BY01, CNK01, EPB05, FJP06, GP06, Gol13, KAD^{+}21, KB01, LK01, PASS01, Wyv03, ZYL^{+}23, ZK01, AGCA06, ACWK06, AD15, BGA05, BB11, CCL22, DLP13, GLXJ14, HWQ14, JSZ^{+}21, LXX^{+}19, MCQ05, SFB^{+}17, XWZ^{+}15, XLX^{+}19, YWZB17, YLG^{+}20, YLL12, CDGH15, GHPW12]. Modelling [LVM04, PFV^{+}11, PBN^{+}09, ZLRL19]. Models [BJM19, CYW04a, GS01, HQ12b, SSP01a, SGS01, BNR^{+}17, BGA05, BMM^{+}07, CVB09, CWM^{+}14, GPF19, GG18, JWL12, KMBG09, KVS15, LQJ^{+}14, LLW^{+}18, MKS^{+}08, NÇ10, Ros10, TMT10, VM06, WZ14, ZGLP12, dSGA15, BAA18, KS00, KB01]. modified [ZGLP12]. molecular [MKB^{+}16, RG12]. Moments [SLB^{+}00]. Momentum [ALP06]. Momentum-based [ALP06]. monocular [YLW^{+}14]. monsters [GLXJ14]. Morphing [GY01, BGA05]. Morphological [KM20, WR05]. Morse [CD11, FIF19, HDL16]. Most [Ano07e]. Motion [Gle01, LS01, MBH^{+}12, PCP02, RTKW15, ALP06, CRH05, CH06, FSF07, GODC07, JLW^{+}21, KL17, KEK14, LB06, LL06, LML^{+}20, Pan03, PY08, TCMS04, ZWSH14, ZB05, ZS09]. Motion-based [MBH^{+}12]. Motions [HJK02, KL13]. moulding [BRC15]. moving [ACH^{+}13, WG19]. MPEG [FSF07]. MPEG-4 [FSF07]. MPU [BMM^{+}07]. Multi [HLX^{+}22, PLL12, TMT10, YXF14, CTAO20, GWYN19, GXT18, GWHH18, KP15a, MJ16, Sta05, TRS06, XYZ^{+}23, ZLZ^{+}22, ZYL^{+}23, ZZC^{+}21]. multi-branch [GXT18]. Multi-core [TMT10]. multi-device [CTAO20]. multi-dimensional [GWHH18]. multi-directional [GWYN19]. multi-level [HLX^{+}22]. multi-material [MJ16]. Multi-scale [HLX^{+}22, PLL12, YXF14, Sta05, TRS06]. multi-style [ZLZ^{+}22, ZYL^{+}23]. multi-surface [KP15a]. multi-UAV [ZZC^{+}21]. multi-view [XYZ^{+}23]. multilayer [kWwZ15]. multilevel [OBS05]. multiple [BSMG05, GMH^{+}20, LML^{+}20, SH05, WS03, ZQ11]. Multiresolution [AMAS16, BBF^{+}11, BMZB02, LS01, PD16, HSS15, UCB13]. Multiscale [FWH13]. multiview [GCZZ18, STRD19, YCKK17].

n [And03]. naive [Buz03]. Natural [FCSF16, LK01, Sta05]. navigation [LXH21]. Near [GWYN19, KADS02]. Near-Optimal [KADS02]. neighborhood [MLF^{+}12]. Net [FMW^{+}22, LXH21]. Network [LWG^{+}21, DZSW23, FMW^{+}22, HYZ^{+}19, HLX^{+}22, PB20, PKL^{+}23, RMCdST19, ZML^{+}20, ZCW^{+}22]. network-distributed [PB20]. Networks [LZL16, BLSZ20, GXT18, HLM16, JLW^{+}21, LSY23b, PPP19, SJ12, ZLS^{+}20]. Neural [KDCC23, BLSZ20, GXT18, ZLS^{+}20, ZML^{+}20, ZHMW22]. Ninth [SRK02a].

Noise [KC01, SRML09, HvBK19, LWG^{+}23, MIPS14, YL08]. noise-aware [LWG^{+}23]. Noisy [GW01, HvBK19, PSK^{+}02, LDZ^{+}17]. Non [PCPM23, PDA03, BNR^{+}17, BS04a, DMMP03, GK03, HLL^{+}23, MPVF11, NPJ14, PLS^{+}18, SLD^{+}23, SLF^{+}18, WMR^{+}14, ZCS^{+}15, ZL15, ZZS19]. non-convex [NPJ14]. Non-homogeneous [PCPM23]. non-incremental [SLF^{+}18]. Non-linear [PDA03]. non-local [MPVF11]. non-manifold [DMMP03]. non-penetration [ZL15]. non-photorealistic [WMR^{+}14]. non-rectangular [SLD^{+}23]. non-rigid [HLL^{+}23, PLS^{+}18, ZCS^{+}15]. non-simple [GK03]. non-stationary [BNR^{+}17, ZZS19]. non-vanishing [BS04a]. nonhomogeneous [LXD19]. Nonlinear [CGPZ21]. NonSelf [KM00]. NonSelf-Occluding [KM00]. Nontriangular [KG01]. norm [ZW03]. Normal [JSZ^{+}21, PSK^{+}02, NSL^{+}21, TH12, ZNYL22, ZQ11]. Note [Ano02e, Ano13d]. Notes [LR12]. Novel [HZL^{+}20, LBSP02, SK01, WLAT14]. November [Ano19, Ano20e, Ano21e, Ano22e]. nucleotides [RG12]. numerical [HKM12, HKL^{+}15]. NURBS [HKM12, KP12, YZC14].

Obituary [Ano23f]. Object

[LI00, ZbQC^{+}19, LV03, ZHMW22]. Object- [LI00]. object-level [ZHMW22]. ObjectFusion [ZHMW22]. Objects [ZXZ21, And03, BS04b, BBB11, BB11, CTL15, CMA14, DMMP03, KMP05, LYKL12, NÇ10, PS07, RNDA13, SOG09, TLGS05, TPFA21, XYZ^{+}23]. obstacles [LHM06]. Occluding [KM00]. occlusion [BTCH05, YG07]. October [Ano23g]. OCTOR [JR09]. Octree [ZZZL13, STRD19, XTW16, ZLRL19]. Octree-based [ZZZL13, STRD19]. octrees [VS08]. odometry [SBSG23]. oil [dCBM^{+}16]. One [Eva11]. Online [SSL17, Par23, YWZB17]. operations [Elb05, EKK15]. operators [ACS03]. optical [IM06]. optics [kWwZ13]. Optimal [ACC14, KADS02, Sbe00, AACPMCMJ16, DB21, GWYN19, KMP05, MLC^{+}17, MMNG19, PD16, ZGX^{+}18]. optimisation [LM12]. Optimised [Vás11]. Optimization [BVL02, GKR02, MK05, CCJ^{+}18, HZW^{+}22, TL05, YLW^{+}14, YLW^{+}19, ZL14, dSGA15]. Optimizing [BBM^{+}17]. optimum [WY11]. orders [CBK03]. organizing [KS04]. orientation [DB21, JWL12, LZL16, XTW16]. Oriented [KAD^{+}21, CCJ^{+}18, STRD19]. ornaments [AT16]. orthogonal [CMA14, MPV13, VPAM12]. Orthogonality [LSA21]. Out-of-core [GF22, LDD14]. outlier [GF22]. overview [PPP19].

Pacific [SRK02a, Ano04g, BSW01, HCS03, KS00]. Packing [HD23]. Painterly [PY08, XTLP04]. painting [FGV^{+}14, LF04]. pants [HDL16]. Paper [Ano07e]. Papers [Ano01b, CDGH15, GHPW12]. Parallel [TCL^{+}20, BC15, CLL^{+}18, ZCW^{+}22, SBA13]. parameter [KSK23, XWYY10, ZLS^{+}20, ZML^{+}20]. parameter-predicting [ZML^{+}20]. parameterization [ALP06, LXZ14, YRZ18, dSGA15]. parameterizations [TJ12, YZC14]. Parameters [KC01, QY02, CYW04a, CYW04b, NÇ10]. Parametric [CGPZ21, AFSW03, CYF^{+}23, JYTM14, ZQ11]. parametrization [WLL14]. part [JLW^{+}21, NPJ14]. partial [BWW^{+}14, KMP05]. particle [GZL^{+}20, GG18]. partition [OBS06b, TRS06]. Partitioning [KM00, GZL^{+}20, LWH15, TL05]. partitions [KCD14]. parts [BBML^{+}18, GC22]. passive [BS04b]. Patch [WLLL22, BJM18, BJM19, BSB14, CMLH19]. Patch-based [WLLL22]. patch-wise [BSB14]. path [DGZ12, KS02, PCPM23]. paths [BHS18, EE19, HYYZ17]. patient [LXX^{+}19]. patient-specific [LXX^{+}19]. pattern [JR09]. patterns [IO09, TLGS05]. PBN [ZCW^{+}22]. PCI [LXX^{+}19]. PDE [DQ05]. PDE-based [DQ05]. PDQ [SBA13]. pedestrian [WXN18]. pedestrian-level [WXN18]. pedestrians [ST07]. penetration [ZL15]. pentahedron [LYCG08]. people [SH05]. Perception [ZWSH14]. Perception-based [ZWSH14]. perfecting [KVS15]. performance [DR23, RW20, TM07, UPBS08]. performance-based [TM07]. performances [PHE23]. periodic [HL21]. periodicity [SDC04]. Perpendicular [ALSR11]. personalized [CSJ13]. persons [LML^{+}20]. perspective [Gol13]. Perspectives [SKMM24]. perturbation [OK07]. Perturbing [ZW03]. PG [SRK02a]. PG2004 [KCOTW06]. phantoms [AO03]. phase [GLY^{+}17, MKS^{+}08]. photo [WSCO^{+}12]. photograph [ZYL^{+}23]. photographs [WMR^{+}14]. photometric [WMR^{+}14]. photorealistic [WMR^{+}14]. PHT [KXCD15]. PHT-splines [KXCD15]. physical [DQ05]. Physically [CBC^{+}07]. Physics [BY01, BDC^{+}19, DLP13, PB20]. Physics-Based [BY01, BDC^{+}19, DLP13]. PIA [LLH22]. Pick [CSP20]. Piecewise [CCF01, SRK02b, CXC14]. pipeline [CTAO20]. pipelines [CWS^{+}20]. Pixel [CC24]. planar [AACPMCMJ16, CW02, KM20, KEK14, SSHS14, ZSZ^{+}20]. Plane [EKH01, LHM06, WJG02]. planes [Buz03]. plant [PBN^{+}09, SFCD21, WTS17]. PlantGL [PBN^{+}09]. plants [FJP06]. Plateau [Hao20]. plausible [Bul23]. plots [KCM20]. plus [MMC23]. Point [DZSW23, KP15b, ZSL^{+}22, AGCA06, BGZ16, CMLH19, CGAY13, Che22, CXY^{+}09, CWS^{+}20, DRD19, ES16, GF22, JXCZ13, LWG^{+}23, LDLS23, MLF^{+}12, MS09, MB23, NBPF11, Nie16, Nie17, NSL^{+}21, PSL14, PLL12, STRD19, Wu02, XTW16, YGL^{+}18, YCKK17, ZCW^{+}22, ZNYL22, ZC18, ZN13]. Point-augmented [KP15b]. point-based [CMLH19]. Points [JK02, Eva11]. Poisson [XZWB06]. Polygon [IS02, KADS02, ZSZ^{+}20]. Polygonal [Tau02b, AACPMCMJ16, DGZ12, LBM04, LVM04, MIPS14, Par23]. Polygonization [GA00]. polygons [Hub12, RSFdM04]. polyhedral [VM06]. Polynomial [DCL^{+}08, XWYY10]. polynomials [CXC14]. pooling [FWX^{+}18]. Pore [LWZ^{+}18]. Pore-scale [LWZ^{+}18]. porosimetry [RCVA11]. porous [HL21, LWZ^{+}18]. Portrait [ZbQC^{+}19, ZYL^{+}23]. pose [CLG^{+}23, DXD14, LGG19, RP08]. position [DB21, ZZD^{+}23]. position-based [DB21]. possible [CGLX17, WLL14]. post [CRH05]. post-production [CRH05]. postproduction [MR05]. pre [KSS08]. pre-integrated [KSS08]. Precise [KEK14]. precision [KSS08]. Precomputing [LL06]. predicting [ZML^{+}20]. predictor [AD15]. predictor-corrector [AD15]. Preface [AHZ14, Ano01e, DHK12, HM13, Jia16, JB14]. preferences [BBM^{+}17]. Preprocess [dSNJMA22]. prescribed [DB21]. presenting [TPG^{+}23]. Preservation [ML00, SS06]. preserving [Eva11, EB17, GYH13, SWC^{+}21, TKPR09, WHHB12, XHH^{+}20]. primitive [AFSW03]. principal [SWW^{+}14]. Printed [CCL22]. printing [GWYN19, LLW^{+}18]. prioritized [LB06]. priors [ZHMW22]. prismatic [GwGgP18]. Probabilities [Sbe00]. Problem [LKE00, AL11, CC24, Hao20, KL14]. Problems [HHS^{+}01, BTCH05, HD23]. Procedural [Bul23, PFV^{+}11, QY02, HvBK19]. Process [KAD^{+}21, WZ14, WTS17]. Process-Oriented [KAD^{+}21]. Processing [Tau02b, DRD19, HWQ14, MPVF11, Ros10, TPG^{+}23, XWZ^{+}15]. product [WG15]. production [CRH05]. programming [HD23]. progressive [Par23, SSL17]. projected [ZGX^{+}18]. Projection [YZZ^{+}10, CSP20, WY11, WXGZ20]. projections [AH16, Gol13]. projector [DB21]. prone [MMNG19]. propagation [HYYZ17, ZRZ21]. Properties [Pet00, SSP01a, CW02, JFS11, JYTM14, RFLSA11, SDC04, TJ12]. Property [IS02]. Provably [BO05]. providing [KKH19]. proxied [CYF^{+}23]. Proximity [GMH^{+}20]. Proximity-aware [GMH^{+}20]. pruning [KL14]. pseudomanifolds [VPAM12]. PU [DZSW23]. PU-GAT [DZSW23]. Publisher [Ano02e, Ano13d]. pyramid [SWYH18, BK03]. Pythagorean [Far02, HFSB14]. Pythagorean-hodograph [Far02]. Python [PBN^{+}09]. Python-based [PBN^{+}09].

quad [BH14, LLXW13, PSL14].

quad-dominant [PSL14].

quadrangulations [HA03]. quadratic [YZZ^{+}10]. Quadric [CCF01, JBK04, CWM^{+}14, GMH^{+}20, WJG02]. quadrilateral [MM19]. qualities [ZB05]. Quality [WY11, BBM^{+}17, FWX^{+}18]. quantum [GS12]. Quasi [Hao20, HL22, TH14]. Quasi-area [Hao20]. quasi-Bézier [HL22]. quasi-conformal [TH14]. Quasiconformal [YZ20]. Quaternion [WG15]. quaternions [Gol11, ZK15]. queries [ES16, FWK20, SBA13]. quintic [HFSB14].

radial [FSF07, OBS06b]. radiance

[RKH05]. radiation [AFBB17]. Radiosity [Sbe00]. Random [Sbe00, DXD14, ME17]. random-looking [ME17]. Range [GW01, PS00]. rank [WLLL22]. rapid [Nie17]. rarity [WSZL13]. rasterization [BBB11]. rasterized [CTAO20]. Rasterizing [RSFdM04]. rate [MM19]. Rational [EKH01, KP11, BFRA12, CW02, WG15, ZZ14]. ray [CTAO20]. ray-tracer [CTAO20]. raycasting [Kim13]. RBFs [OBS05]. re [FCSF16, KGZ^{+}14]. re- [KGZ^{+}14]. re-design [FCSF16]. Real [BIP00, JSL^{+}20, TGB00, WCHZ14, YWL^{+}19, ZZZY13, BHS18, BGTG04, CTAO20, LXX^{+}19, LXTJ16, NSL^{+}21, PHE23, PCPM23, PDA03, YLW^{+}19]. Real-Time [BIP00, TGB00, JSL^{+}20, WCHZ14, ZZZY13, BHS18, BGTG04, CTAO20, LXX^{+}19, LXTJ16, NSL^{+}21, PCPM23, PDA03, YLW^{+}19]. Realistic [XL23, BDC^{+}19, VRP22]. reality [PCPM23, RW20, WS03]. realtime [ZZZL13]. reassembly [ZL14]. Receptive [BK03]. recognition [ALSR11, Buz03, FWH13, MS09, SWW^{+}14]. Recognizing [AH16]. recommendation [Tsc20]. Reconstructing [HKZM20, KL13, TRS06]. Reconstruction [CGPZ21, GW01, JK02, KSS00, LWG^{+}21, PS00, SGHM00, ZK01, BMM^{+}07, CGAY13, CYF^{+}23, CWS^{+}20, DQ04, HYZ^{+}19, KSK23, KL17, KK23, LHM06, OBS06b, RTKW15, VBN11, WSCO^{+}12, WHHB12, YWH^{+}21, ZZZL13, ZZD^{+}23, ZN13]. recovery [LWG^{+}23, WLLL22]. rectangular [SLD^{+}23]. Rectification [SSP01b]. rectilinear [YZ20]. recursive [JR09]. Reducing [HKL^{+}15]. reduction [BLSZ20, ZW03]. Reeb [HA03]. refinement [CD11, KSK23, TCL^{+}20]. reflection [LJYL16]. Refraction [RC06]. Region [ZMW^{+}14, Bay19, FWC^{+}19]. Region-based [ZMW^{+}14]. Regions [KM00, BE11]. Registration [LHS01, YCKK17, BBF^{+}11, BSB14, LZY^{+}14, MLC^{+}17, MMNG19, YRZ18, ZLW^{+}14, ZCS^{+}15]. Regular [SRK02b, Gus07, HD23, KCD14]. regularity [TJ12]. regularization [BJK14]. Related [HHS^{+}01]. Relations [ZXZ21]. reliable [GWHH18]. relief [JSZ^{+}21, NSL^{+}21, XLX^{+}19, ZMW^{+}14, ZZLZ13, ZZZY13, ZbQC^{+}19, ZYL^{+}23]. reliefs [Nie17, WMR^{+}14]. RELIGHT [PCS19]. remeshing [AdVDI05, Gus07, LM12, ZGLP12]. remote [MKB^{+}16]. removal [GF22]. Rendering [BIP00, CRH05, RCG^{+}09, BSMG05, BS04b, GY05, KSH18, KSS08, PCPM23, PD16, RL13, WLW06, kWwZ13, WMR^{+}14, XXZL23, XTLP04, ZWSH14]. rep [HR23]. repair [Att14]. reparameterization [YZZ^{+}10]. Representation [KAD^{+}21, LB04, SSP01b, ZK15, BWW^{+}14, GK04, LL13, LDLS23, PCS19, SS06, TPFA21, UCB13, ZSC^{+}14, LV03]. representations [CAF09, LSY23b, MPV13, NBPF11]. represented [BB11, GC22]. Representing [RJG23]. Resampling [Tau02a]. reservoirs [dCBM^{+}16]. residential [ZN13]. resizing [JZLZ14]. Resolution [FCG01, KSH18, KZD^{+}11, LWG^{+}23, SFB^{+}17, TPG^{+}23]. resonance [KZD^{+}11]. respiration [ZCCD06]. Retrieval [CFG06, LGG19, PLS^{+}18, WLT14]. retroreflective [GwGgP18]. reverse [KVS15, SS11]. review [YLG^{+}20, ZSL^{+}22]. Reviewer [Ano00b, Ano02f, Ano03f, Ano04i, Ano05h, Ano06g, Ano07g, Ano09h, Ano10f, Ano11f, Ano12e, Ano13e, Ano15g, Ano01f]. Revolution [RCHS18, JBK04, WG19]. RFMNet [HLL^{+}23]. RGB [CLG^{+}23, GH03, HKZM20]. RGBN [PBM^{+}11]. RGC [AO03]. RGC-sm [AO03]. rhythms [RMCdST19]. Ricci [ZGZ^{+}14]. Riemannian [ZGLG12]. rigging [CBC^{+}07]. right [AO03]. rigid [CGLX17, HLL^{+}23, LSY^{+}23a, LZY^{+}14, PLS^{+}18, WLL14, ZCS^{+}15]. Rigidity [CGLX17]. rigidly [ACH^{+}13]. Ringed [HHS^{+}01, ZQ11]. road [SJ12]. robot [LHM06]. Robust [BLSZ20, MLC^{+}17, PS03, RNDA13, SK01, WWWM12, ZHM11, ACS03, BWW^{+}14, HR23, MPVF11, ZXY^{+}12, HLL^{+}23]. Room [HKZM20]. rooms [SLD^{+}23]. Rotation [CK00, WEY06, FMW^{+}22, Far02]. rotation-minimizing [Far02]. rotationally [MPVF11]. rotations [Gol13, RFLSA11]. RTI [PCS19]. ruled [MP03].

S.O.M. [BLT05]. salience [BBML^{+}18, SWW^{+}14]. saliency [FWX^{+}18, JZLZ14, WXGZ20, WSZL13]. sampling [BO05, GG18, WZ14]. satellite [WXN18]. satellite-based [WXN18]. SCA [CPOO09]. scaffold [HL21]. Scale [SK21, Bul23, GF22, HLX^{+}22, LWZ^{+}18, LXTJ16, PLL12, RW20, Sta05, TRS06, YXF14]. Scale-Adaptive [SK21]. scales [PBN^{+}09]. scaling [KMP05]. scanline [ZGX^{+}18]. scanners [SRML09]. scanning [BLT05]. scans [LCZG14]. Scattered [GKR02, KS04, MK05, OBS05, OBS06a, XTW16]. scenarios [MDLH19]. Scene [GW01, KL17, RKK^{+}00, CMLH19, FFY^{+}20, HFH16, LGG19]. scenes [FW16, GY05, LXTJ16, TPG^{+}23]. scheme [BNR^{+}17, HYYZ17, HMESI13, LLXW13, MM19, MB23, XHH^{+}20, ZZS19]. schemes [BFRA12]. scissor [AGH16]. Sculpting [FCG01]. segmentability [NPJ14]. Segmentation [ACH^{+}13, CMB^{+}12, DR03, GXT18, KS02, MBH^{+}12, NPJ14, NFU02, WLAT14, ZLS^{+}20]. Segmenting [HDL16]. segments [BHS18]. selected [CDGH15, GHPW12]. selection [JR09, MV23, YGL^{+}18]. Self [ZZ14, Att14, KS04, LLW^{+}18, LDLS23]. self-intersecting [Att14]. Self-intersections [ZZ14]. self-organizing [KS04]. self-supervised [LDLS23]. self-supporting [LLW^{+}18]. Semantic [HYZ^{+}19, FWC^{+}19, FW16, FWK20, RW20, STRD19, YWZB17]. semantically [Bul23]. semantics [DRD19, HS05]. semi [Gus07, SBSG23]. semi-regular [Gus07]. semi-synthetic [SBSG23]. seminar [CDGH15, GHPW12]. sensitive [LFC^{+}18]. sensor [HLM16]. sentiment [RJG23]. September [Ano20f, Ano21f, Ano22f]. sequence [WZL^{+}03]. sequences [ACH^{+}13, PS03, RG12, SH05]. Sequential [UPBS08]. Service [STRD19]. Service-oriented [STRD19]. services [RW20]. set [ACS03, EB17, XTW16]. Sets [GW01, SR00, AGCA06, BGZ16, CGAY13, GK03, KL14, WSCO^{+}12]. Shape [GP06, LYCG08, Wyv03, BSMG05, BBML^{+}18, BSB14, CYW04a, CYW04b, CGLX17, FWC^{+}19, FCSF16, GG18, HS05, HLL^{+}23, HLX^{+}22, KSK23, KP15a, LLL19a, LFC^{+}18, PLS^{+}18, SFB^{+}17, SSL17, TH14, WXZ^{+}16, XZWB06, XWYY10, ZZS19, ZFLL23, ZSC^{+}14, ZLAK14]. shape-interrogation [HS05]. shapes [BKL15, KP12, LL13, LZL16, MJ16, MPV13, WMC^{+}22]. shared [SWC^{+}21]. Sharp [BMZB02, WHHB12, ZNYL22]. SharpNet [ZNYL22]. Shear [CK00]. shears [AGH16, WEY06]. sheeting [GwGgP18]. ship [YCKK17]. shockwaves [SGTL09]. short [PS03]. shortest [ES16]. shrink [Nie16]. shrinkage [XL23]. SIFT [DIOV06]. Silhouette [LHS01, TCMS04]. Silhouette-Based [LHS01, TCMS04]. silhouettes [LYCG08, MBH^{+}12]. similarity [TH14]. Simple [BKL15, BPG05, Eva11, GK03, GH03]. simplicial [FIF19]. Simplification [BVL02, KSS00, BBP18, CD11, CMA14, DGZ12, LLW^{+}18, VS08, WR05, ZWSH14]. simulating [LGG19]. Simulation [LWGP08, ZYP09, CTAO20, GZL^{+}20, HKL^{+}15, JSL^{+}20, KMBG09, LWZ^{+}18, LXX^{+}19, LSY^{+}23a, LZX^{+}15, MDLH19, PDA03, SGTL09, kWwZ15, XL23, YLG^{+}20, ZLX^{+}18]. simulations [AD15, PB20, SKMM24, WXN18, WBOL07]. simultaneous [LXH21]. simultaneously [WMC^{+}22]. Single [LWG^{+}21, CLG^{+}23, LLL19a, LVM04, NRW23, RKH05, WLAT14, WZtW^{+}14, YWH^{+}21, ZMW^{+}14, ZYL^{+}23]. singular [TJ12]. six [KK23]. six-direction [KK23]. size [CFG06]. Skeletal [MPV13, WXGZ20, AGCA06]. Skeleton [CYW04a, CYW04b, JXCZ13, LQJ^{+}14, SS06, WXZ^{+}16, HXS09, JXC^{+}13, SS19, WWWM12]. Skeleton-based [JXCZ13]. Skeleton-driven [CYW04a, CYW04b]. Skeleton-enhanced [LQJ^{+}14]. Skeleton-guided [WXZ^{+}16]. skeletonization [TTF04]. skeletons [BGLSS04, ZSZ^{+}20]. Sketch [FML12, PBM^{+}11, ZDL^{+}11, WLT14, YWH^{+}21, YLW^{+}14]. Sketch-based [FML12, PBM^{+}11, WLT14]. sketches [XLX^{+}19]. Sketching [CO20a]. Skin [XL23, BDC^{+}19, CGW^{+}07]. Skinning [She03, KKH19, PKL^{+}23]. skins [BKL15]. skull [YLL12]. Skyline [BBP18]. Skyline-based [BBP18]. SLAM [ZHMW22]. slice [YG07]. slice-wise [YG07]. slow [yKL11]. slow-in [yKL11]. slow-out [yKL11]. sm [AO03]. small [Bul23, MV23]. small-scale [Bul23]. Smart [CCS23]. SMI [Wyv03, PS05, SCOG09]. smoke [HKL^{+}15]. Smooth [SS11, SWHH15, BJM18, BJM19, KKH19, Nie16, SFB^{+}17, ZZC^{+}21]. Smoothing [LBSP02, GYH13, NKP11, WY11]. software [BLT05]. solar [BBP18]. Solid [EPB05, SSP01a, CL19, GLY^{+}17, SS06, SLF^{+}18, ZDL^{+}11, ZLH13, ZGLP12]. solid-liquid-gas [GLY^{+}17]. solids [HL21, NPJ14]. solution [BLT05]. Solutions [LKE00, MMC23]. solver [AL11]. Some [HHS^{+}01]. sonar [CTAO20]. Sorting [WSCO^{+}12]. sound [CL19, ZRZ21]. sources [WS03]. Space [LI00, SO01, BTCH05, CAF09, Gra15, LSA21, LWH15, NSZ18, NSL^{+}21, SWC^{+}21, TH14, Thü03, WG15, YWZB17, YL14, ZLS^{+}20]. spaces [Eva11]. Sparse [MS10, OBS06b, AFBB17, EKK15, LHM06, LLW^{+}18]. sparsity [XWZ^{+}15]. Spatial [SO01, BBM^{+}17, Far02, HFSB14, ZZD^{+}23]. spatial-temporal [ZZD^{+}23]. spatially [WZ14]. spatio [WXRA07]. spatio-temporal [WXRA07]. Special [Ano01b, Ano06f, BSW01, BAA18, CNK01, CDGH15, GP06, GHPW12, HB05, HCS03, KS00, KCOTW06, KSM^{+}06, MYC17, SCOG09, SRK02a, Tau02b, WP00, Wyv03, CCS23, HM13, Jia16, PS05]. specific [DXD14, LXX^{+}19]. specified [YL14]. Spectral [WLAT14]. Speed [LI00, GZL^{+}20]. speeds [NFU02]. SPH [GZL^{+}20, HZL^{+}20]. SPH-centric [HZL^{+}20]. sphere [BNR^{+}17]. sphere-like [BNR^{+}17]. spheres [KKH19, ZQ11]. Spherical [CS18, AMAS16, KCM20, TPG^{+}23, WLL14]. SPIDER [TPG^{+}23]. Spiral [BE11]. Splat [CGAY13]. Splat-based [CGAY13]. Splatting [LI00]. Spline [CL00, HJK02, KM00, XLX^{+}19, CXY^{+}09, HL21, KK23, LLH22, MKH^{+}17, XWYY10, YZZ^{+}10, ZSC^{+}14]. Splines [GSS00, BJM18, BJM19, DCL^{+}08, GHQ06, HFSB14, KCD14, KXCD15, KL19, KP11, KP12, Kim17, KGZ^{+}14, ZK05, ZPG18, ZQ11]. Split [DR03]. Split-and-merge [DR03]. Splitting [SOG09]. SPM [KSM^{+}06]. spring [JSL^{+}20, NÇ10, dSGA15]. Stabilized [CO20b]. Stable [YL08]. STAG [CCS23]. stage [CTL15]. stance [NF06, RJG23]. standard [And03, Buz03]. state [AO03, BGTG04]. static [AD15, GY05]. stationary [BNR^{+}17, ZZS19]. step [MV23]. stereo [WMR^{+}14]. Stick [WLW06]. stippling [SLKL11]. Stitching [LHS01]. Stochastic [SGHM00, AO03]. Stokes [LWH15]. straight [BGLSS04]. strand [CCL22]. strand-level [CCL22]. strategies [HZL^{+}20]. strategy [FWX^{+}18]. stream [BKC15]. streaming [MB23]. streams [KLV06]. street [PPP19]. Stress [CCJ^{+}18]. Stress-oriented [CCJ^{+}18]. Stretch [BKC15, Bay19]. Stretch-minimising [BKC15]. strict [ZK08]. striped [TLGS05]. Strips [RCG^{+}09]. strokes [Par23]. structural [CCJ^{+}18, VN23]. Structure [HFH16, LXZ14, SLKL11, FWWT13, LWZ^{+}18, LZX^{+}15, QSS^{+}19, YG07, ZFLL23, ZLRL19]. Structure-aligned [LXZ14]. structure-guided [QSS^{+}19]. structured [NRW23]. structured-light [NRW23]. structures [BBF^{+}11, CCJ^{+}18, CMA14, LVM04, WWWM12]. studies [FRDC06]. studio [RP08]. study [AT16, BFRA12, FML12]. style [CRH05, KDCC23, PHE23, ZLZ^{+}22, ZYL^{+}23]. style-aware [PHE23]. stylization [LYKL12]. sub [BBML^{+}18]. sub-parts [BBML^{+}18]. Subdivision [BMZB02, JK02, BNR^{+}17, DQ04, HMESI13, KP15b, LLXW13, MM19, MK05, MCQ05, MMS^{+}07, SS11, UCB13, ZZS19]. subdivision-based [DQ04, MCQ05]. subject [DXD14, HFSB14]. suboptimal [GYH13]. Subset [JR09]. suitable [KL19, ZPG18]. sum [MKH^{+}17, MP03, PS07, VM06]. sunken [ZZLZ13]. Superquadrics [ZK01]. supervised [LDLS23]. Supine [MMNG19]. Supplementary [Ano13f]. support [GWYN19, Hub12, RCHS18, ZLRL19]. support-free [GWYN19]. supported [OBS05, SFB^{+}17]. supporting [LLW^{+}18, LZX^{+}15]. Surface [CM20, GSS00, JK02, KZD^{+}11, LKE00, MK02, PS00, TH12, AdVDI05, BMM^{+}07, CGAY13, CXY^{+}09, DQ05, DQ04, EB17, GYH13, HDL16, HWQ14, IO09, JXC^{+}13, KP15a, KCM20, LF04, LXD19, LXZ14, LDD14, MLC^{+}17, MPVF11, MCQ05, OBS06b, RKH05, SOG09, TL05, VSR12, WJG02, WY11, WHHB12, ZGZ^{+}14, ZSJG14]. Surface-based [CM20, KZD^{+}11]. Surfaces [BMZB02, CCF01, Elb01, GA00, HHS^{+}01, KSS00, KM00, LM00, ML00, Pet00, SK01, WG19, AGH16, AGCA06, AFSW03, BRC15, BKC15, BO05, CTL15, FH12, GWHH18, HKM12, HL22, HA03, HL21, Hub12, JBK04, JYTM14, KP15b, LLH22, MJ16, MK05, MIPS14, MKH^{+}17, MP03, PS07, SMKE14, TRS06, WSC^{+}12, WG15, WLTS15, XWYY10, YZZ^{+}10, YZC14, YXF14, YRZ18, ZQ11, ZSZ^{+}20]. surgery [LXX^{+}19, PDA03]. Survey [XWZ^{+}15, RJG23]. SVD [DIOV06]. SVD-matching [DIOV06]. sweepers [ACWK06, AWC06]. Swendsen [WZ14]. Swept [AWC06]. Swirling [ACWK06]. Swirling-sweepers [ACWK06]. symmetric [BC15, HSS15, WMC^{+}22]. symmetric/antisymmetric [HSS15]. Symmetrization [JYD^{+}16]. symmetry [AT16, JXCZ13, Kim17, LJYL16, WXGZ20]. Symposium [CPOO09, Jam09]. synthesis [AO03, CL19, CH06, FFY^{+}20, GODC07, HFH16, OH06, Sta05, SH05, SRML09, WS03, WZL^{+}03, WZtW^{+}14, ZK08, ZLH13]. synthetic [SBSG23]. system [AGCA06, CN16, CGW^{+}07, GODC07, JSL^{+}20, LXD19, LWH15, ZLRL19]. systematic [VN23]. systems [BKL15, DLP13, ZLW^{+}14].

T [DCL^{+}08, KL19]. T-meshes [DCL^{+}08]. T-splines [KL19]. tactile [FGV^{+}14]. TAD [FMW^{+}22]. TAD-Net [FMW^{+}22]. technique [GG18, TM07]. Techniques [GPF19, GKR02, SK01, TGB00, BTCH05, DRD19, MV23, PPP19]. Teichmüller [MLC^{+}17]. temperature [ZLX^{+}18]. temporal [ACH^{+}13, RMCdST19, WXRA07, ZZD^{+}23]. Temporally [LYL10]. tensor [LR12, PLL12, TH12]. Terrain [KSH18, ZLZ^{+}22]. tessellation [CXC14]. tessellations [CS18, RSFdM04]. Tetra [JFS11]. Tetra-trees [JFS11]. tetrahedra [OK07]. Tetrahedral [ITF06, BGTG04, BPG05, ZGLP12].

Textiles [KAD^{+}21]. Texture [BIP00, LHS01, EB17, JSS^{+}14, PT15, ZK08]. textured [BGA05]. textures [Sta05, WLW06, ZLH13]. Texturing [QY02, QSS^{+}19, ZDL^{+}11]. THB [KGZ^{+}14]. THB-splines [KGZ^{+}14]. their [JYTM14, TG13]. theorem [CSP20]. Theory [EPB05, GwGgP18, HDL16]. thickness [RNDA13]. Thinning [NKP11, BC15, Nie16]. Three [BSMG05, MS09, SGS01]. Three-dimensional [BSMG05, MS09]. Tiles [ZK08]. Time [BIP00, TGB00, BHS18, BGTG04, CTAO20, JSL^{+}20, LXX^{+}19, LXTJ16, NSL^{+}21, PCPM23, PDA03, WCHZ14, YXF14, YLW^{+}19, ZZZY13]. time-varying [YXF14]. timing [TM07]. tomographic [VBN11]. tomography [IM06]. tool [EE19, MR05, SKMM24]. tool-paths [EE19]. Tools [CCS23]. tooth [FMW^{+}22, ZLS^{+}20]. TopoKnit [KAD^{+}21]. Topological [HA03, LM00, SSP01a, TTF04, VS08, CN18, DR03]. Topology [KAD^{+}21, ML00, SR00, WTS17, BNR^{+}17, DQ04, Eva11, LV03, SS06, UCB13]. topology-preserving [Eva11]. Total [BJK14]. towns [Bul23]. tracer [CTAO20]. tracing [GG18, PCPM23]. tracking [SOG09, WLTS15]. trademark [CFG06]. traffic [CSJ13, SJ12, YWL^{+}19]. trait [YWZB17]. trajectories [KL13]. trajectory [BHS18]. transfer [EB17, JLW^{+}21, KDCC23, SS19, SWC^{+}21, TTF04, YGL^{+}18]. Transfinite [VSR12]. Transform [CS01, LSY^{+}23a, ZSC^{+}14]. transformation [FMW^{+}22, KLV06, ZZC^{+}21]. Transformations [CK00, Eva11, ZC18]. transformer [ZZD^{+}23]. transforms [HQ12a, HQ12b, MS10]. transition [GLY^{+}17]. transitions [KKH19]. translation [CC24]. transport [MLC^{+}17, MMNG19]. transpose [CO20b]. traversal [Vás11]. Tree [LWG^{+}21, BBF^{+}11, LVM04, VRP22, YWZB17, ZLRL19]. tree-like [ZLRL19]. trees [GK04, JFS11]. triangle [AGCA06, DB21, LLXW13, LV14, QHXC12]. triangle/quad [LLXW13]. triangles [OK07]. triangular [KCD14, LDZ^{+}17, SWYH18, SLF^{+}18]. triangulated [HA03]. triangulation [AFSW03, Bay19, GYH13]. trigonometric [ZK05]. Trimmed [KM00]. triply [HL21]. trivariate [HL21]. trunk [TCH07]. tubular [LZX^{+}15]. Two [HHS^{+}01, MV23, BJM18, BJM19, CTL15, KCM20, LWGP08, MKS^{+}08, TCH07, WG15]. two-dimensional [KCM20]. two-fluid [LWGP08]. two-patch [BJM18, BJM19]. two-phase [MKS^{+}08]. two-stage [CTL15]. Two-step [MV23].

UAV [ZZC^{+}21]. Uncertainty [GSS00, GWHH18, RJG23]. Understanding [Gol11]. Unified [KSK23, LBM04, ZK05, ZGZ^{+}14]. Uniform [CL00, XWYY10, ZK05]. unity [OBS06b, TRS06]. unknown [DQ04]. unlimited [KSH18]. unordered [MB23]. Unorganized [JK02, MK02, PSL14, WSCO^{+}12]. unstructured [MM19, PLL12]. Unsupervised [JWL12, PHE23, HLL^{+}23]. Untangling [HZW^{+}22]. Upright [LZL16, JWL12]. upsampling [DZSW23, LWG^{+}23]. Urban [BAA18, AFBB17, BBP18, SJ12, WSCO^{+}12, WXN18, YG07, ZN13]. Use [WMR^{+}14]. user [BBM^{+}17, YL14]. user-specified [YL14]. Using [BVL02, CK00, EE19, GKR02, LI00, SO01, SGHM00, BBM^{+}17, BSB14, BMM^{+}07, CGPZ21, CO20b, CYW04a, CYW04b, CC24, DIOV06, DQ05, EKK15, FWX^{+}18, GMH^{+}20, GWHH18, GK04, GC22, HDL16, HKM12, HA03, HL21, HD23, HFSB14, JLW^{+}21, KMBG09, KCM20, KS04, KWFH15, LSY^{+}23a, LXZ14, LLW^{+}18, LVM04, MLC^{+}17, ME17, MBH^{+}12, PY08, QSS^{+}19, RG12, RTKW15, STRD19, SWW^{+}14, SFCD21, TMT10, TLGS05, VS08, VN23, WLT14, YZZ^{+}10, YG07, YL14, YLL12, ZK08, ZHM11, ZXY^{+}12, ZGX^{+}18, ZLS^{+}20, ZML^{+}20, ZQ11, ZK15]. utilising [Tsc20].

validating [ZB05]. vanishing [BS04a]. variance [ZRZ21]. variation [BJK14]. variational [TRS06]. varying [SFB^{+}17, YXF14]. vascular [WWWM12]. Vector [PSK^{+}02, Che22, ZNYL22]. vehicle [GVK06]. ventilation [WXN18]. Vertex [KLV06, LM12, ZZD^{+}23]. Vertex-transformation [KLV06]. Very [SGHM00, RTKW15]. Vessel [SGHM00]. vestibular [ZLW^{+}14]. VGF [LXH21]. VGF-Net [LXH21]. via [FWC^{+}19, GYH13, GWYN19, GXT18, GG18, GCZZ18, HFH16, HLX^{+}22, HZW^{+}22, LWZ^{+}18, LWG^{+}21, MS09, PKL^{+}23, WY11, WLLL22]. Video [CSJ13, CH06, HB05, CRH05, FRDC06, LYL10, LYKL12, SH05, TCMS04, WXRA07]. Video-based [CSJ13]. Videoshop [WXRA07]. view [LHM06, LLL19a, LJYL16, SH05, XYZ^{+}23]. view-based [LJYL16]. Virtual [SH05, XTLP04, DC04, PCPM23, RP08, RCVA11, YWL^{+}19]. Visibility [KM00]. Visible [MLF^{+}12]. Vision [HB05, MR05]. vision-based [MR05]. Visual [LHM06, LXH21, MYC17, SGTL09, FWX^{+}18, HM13, RMCdST19, SBSG23]. Visual-Geometric [LXH21]. visualisations [Tsc20]. Visualization [LKE00, SK01, YL14, BLSZ20, DRD19, JYTM14, LXTJ16, MMNG19, MKB^{+}16, Ros10, YG07]. visualizations [RJG23]. Visually [ZZC^{+}21]. Volume [Ano00a, Ano01a, Ano01b, Ano02g, Ano03e, Ano03f, Ano04h, Ano04i, Ano05g, Ano05h, Ano06g, Ano07f, Ano07g, Ano09a, Ano15g, CK00, CNK01, FCG01, KK23, LBSP02, ACWK06, FWWT13, IA03, KSS08, LDD14, RC06, TTF04, WBOL07, XL23]. volumes [WEY06]. Volumetric [GS01, BJM18, BJM19, PFV^{+}11, VS08, ZCS^{+}15]. Voronoi [AdVDI05, CS18, CXC14, KZW12, QZZ^{+}23]. Voting [PSK^{+}02, PLL12]. voxel [KWFH15]. Voxelization [SK01, ZGX^{+}18]. Voxels [SR00]. VR [MV23].

Walk [Sbe00]. WallNet [HKZM20]. Wang [ZK08]. warping [PBM^{+}11]. wave [kWwZ13]. wavelet [HQ12a, HQ12b]. web [DRD19, FWK20, MKB^{+}16, PD16, PCS19]. web-based [DRD19, PD16]. Web3D [HLM16, Jia16, LXTJ16]. Weft [KAD^{+}21]. Weft-Knitted [KAD^{+}21]. weight [YLW^{+}19]. weighted [MJ16, TL05, YL14]. weighting [FWX^{+}18, ZRZ21]. weighting-based [FWX^{+}18]. whole [KL13]. whole-body [KL13]. wireframe [LLW^{+}18]. wise [BSB14, YG07]. Within [ML00, BK03]. without [CS18, RSFdM04]. words [DXD14]. workflow [HR23]. Workshop [KB01]. Wrinkle [SFCD21, CGW^{+}07]. wrist [KL13].

Yarns [KAD^{+}21].

Zometool [ZLAK14].

References

Aguilera-Aguilera:2016:FCO

[AACPMCMJ16] E. J. Aguilera-Aguilera, A. Carmona-Poyato, F. J. Madrid-Cuevas, and M. J. Marín-Jiménez. Fast computation of optimal polygonal approximations of digital planar closed curves. Graphical Models, 84(??):15–27, March 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070316000175.

Argudo:2015:BFM

[ABCV15] Oscar Argudo, Pere Brunet, Antoni Chica, and Àlvar Vinacua. Biharmonic fields and mesh completion. Graphical Models, 82(??):137–148, November 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000296.

Allain:2014:OCE

[ACC14] Pierre Allain, Nicolas Courty, and Thomas Corpetti. Optimal crowd editing. Graphical Models, 76(1):1–16, January 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000271.

Arcila:2013:STM

[ACH^{+}13] Romain Arcila, Cédric Cagniart, Franck Hétroy, Edmond Boyer, and Florent Dupont. Segmentation of temporal mesh sequences into rigidly moving components. Graphical Models, 75(1):10–22, January 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000720.

Akleman:2003:MCS

[ACS03] Ergun Akleman, Jianer Chen, and Vinod Srinivasan. A minimal and complete set of operators for the development of robust manifold mesh modelers. Graphical Models, 65(5): 286–304, September 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Angelidis:2006:SSC

[ACWK06] Alexis Angelidis, Marie-Paule Cani, Geoff Wyvill, and Scott King. Swirlingsweepers: Constant-volume modeling. Graphical Models, 68(4):324–332, July 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070306000038.

Arikatla:2015:IPC

[AD15] Venkata S. Arikatla and Suvranu De. An iterative predictor-corrector approach for modeling static and kinetic friction in interactive simulations. Graphical Models, 82(??):29–42, November 2015. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000491.

Alliez:2005:CVD

[AdVDI05] Pierre Alliez, Éric Colin de Verdière, Olivier Devillers, and Martin Isenburg. Centroidal Voronoi diagrams for isotropic surface remeshing. Graphical Models, 67(3):204–231, May 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Aguerre:2017:CUR

[AFBB17] J. P. Aguerre, E. Fernández, G. Besuievsky, and B. Beckers. Computing urban radiation: a sparse matrix approach. Graphical Models, 91(??):1–11, May 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300486.

Attene:2003:MIP

[AFSW03] Marco Attene, Bianca Falcidieno, Michela Spagnuolo, and Geoff Wyvill. A mapping-independent primitive for the triangulation of parametric surfaces. Graphical Models, 65(5): 260–273, September 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Allegre:2006:HMS

[AGCA06] Rémi Allègre, Eric Galin, Raphaëlle Chaine, and Samir Akkouche. The HybridTree: Mixing skeletal implicit surfaces, triangle meshes, and point sets in a free-form modeling system. Graphical Models, 68(1):42– 64, January 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070305000561.

Alcazar:2016:ASI

[AGH16] Juan G. Alcázar, Ron Goldman, and Carlos Hermoso. Algebraic surfaces invariant under scissor shears. Graphical Models, 87(??):23–34, September 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316300273.

Alcazar:2016:RPA

[AH16] Juan Gerardo Alcázar and Carlos Hermoso. Recognizing projections of algebraic curves. Graphical Models, 87(??):1–10, September 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070316300194.

Alliez:2014:P

[AHZ14] Pierre Alliez, Ying He, and Yongjie Zhang. Preface. Graphical Models, 76(5):239, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000423.

Aristidou:2011:FFI

[AL11] Andreas Aristidou and Joan Lasenby. FABRIK: a fast, iterative solver for the inverse kinematics problem. Graphical Models, 73(5):243–260, September 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070311000178.

Abe:2006:MBP

[ALP06] Yeuhi Abe, C. Karen Liu, and Zoran Popović. Momentum-based parameterization of dynamic character motion. Graphical Models, 68(2):194–211, March 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000299.

Andres:2011:GPB

[ALSR11] Eric Andres, Gaëlle Largeteau-Skapin, and Marc Rodríguez. Generalized perpendicular bisector and exhaustive discrete circle recognition. Graphical Models, 73(6):354–364, November 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070311000233.

Alderson:2016:MSC

[AMAS16] Troy Alderson, Ali Mahdavi-Amiri, and Faramarz Samavati. Multiresolution on spherical curves. Graphical Models, 86(??):13–24, July 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316300078.

Andres:2003:DLO

[And03] Eric Andres. Discrete linear objects in dimension n: the standard model. Graphical Models, 65(1–3):92–111, May 2003. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2000:AIVe

[Ano00a] Anonymous. Author index for volume 62. Graphical Models, 62(6):447, November 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2000.
0535;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0535/pdf.

Anonymous:2000:RA

[Ano00b] Anonymous. Reviewer acknowledgment. Graphical Models, 62(6):445, November 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2000.
0534;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0534/pdf.

Anonymous:2001:AIVe

[Ano01a] Anonymous. Author index for volume 63. Graphical Models, 63(6):480, November 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2001:CPS

[Ano01b] Anonymous. Call for papers: Special issue on volume modeling. Graphical Models, 63(1):61–62, January 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/

                                        doi/10.1006/gmod.2001.
                                        0538;   http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.2001.0538/pdf.

Anonymous:2001:EAb

[Ano01c]  Anonymous.  Editorial

announcement. Graphical Models, 63(2):63–64, March 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/
                                        doi/10.1006/gmod.2001.
                                        0545;   http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.2001.0545/pdf.

Anonymous:2001:EAa

[Ano01d] Anonymous. Editorial announcement. Graphical Models, 63(2):??, March 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2001:P

[Ano01e]  Anonymous.  Preface.

Graphical Models, 63(2):??, March 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2001:RA

[Ano01f] Anonymous. REVIEWER ACKNOWLEDGMENT. Graphical Models, 63(6): 479, November 2001. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2002:Aa

[Ano02a] Anonymous. Announcement. Graphical Models, 64(5):333, September 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2002:Ab

[Ano02b] Anonymous. Announcement. Graphical Models, 64(6):396, November 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2002:EBa

[Ano02c]  Anonymous.  Editorial

Board. Graphical Models, 64(5):C2, September 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2002:EBb

[Ano02d]  Anonymous.  Editorial

Board. Graphical Models, 64(6):C2, November 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2002:PN

[Ano02e] Anonymous. Publisher’s note. Graphical Models, 64 (5):iii–iv, September 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2002:RAL

[Ano02f] Anonymous. Reviewer acknowledgement list. Graphical Models, 64(6):397, November 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2002:VAI

[Ano02g] Anonymous. Volume author index. Graphical Models, 64(6):398, November 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2003:EBj

[Ano03a]  Anonymous.  Editorial

Board. Graphical Models, 65(1–3):C2, May 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2003:EBk

[Ano03b]  Anonymous.  Editorial

Board. Graphical Models, 65(4):C2, July 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2003:EBl

[Ano03c]  Anonymous.  Editorial

Board. Graphical Models, 65(5):C2, September 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2003:EBm

[Ano03d]  Anonymous.  Editorial

Board. Graphical Models, 65(6):CO2, November 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2003:VAIe

[Ano03e] Anonymous. Volume author index. Graphical Models, 65(6):405, November 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2003:VRA

[Ano03f] Anonymous. Volume reviewer acknowledgement list. Graphical Models, 65 (6):406, November 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:EBk

[Ano04a]  Anonymous.  Editorial

Board. Graphical Models, 66(1):CO2, January 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:EBl

[Ano04b]  Anonymous.  Editorial

Board. Graphical Models, 66(2):CO2, March 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:EBm

[Ano04c]  Anonymous.  Editorial

Board. Graphical Models, 66(3):CO2, May 2004. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:EBn

[Ano04d]  Anonymous.  Editorial

Board. Graphical Models, 66(4):CO2, July 2004. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:EBo

[Ano04e]  Anonymous.  Editorial

Board. Graphical Models, 66(5):CO2, September 2004. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:EBp

[Ano04f]  Anonymous.  Editorial

Board. Graphical Models, 66(6):CO2, November 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:PG

[Ano04g] Anonymous. Pacific Graphics 2003. Graphical Models, 66(6):331–332, November 2004. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:VAIe

[Ano04h] Anonymous. Volume author index. Graphical Models, 66(6):440, November 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:VRA

[Ano04i] Anonymous. Volume reviewer acknowledgement list. Graphical Models, 66 (6):439, November 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2005:EBj

[Ano05a]  Anonymous.  Editorial

Board. Graphical Models, 67(1):CO2, January 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2005:EBk

[Ano05b]  Anonymous.  Editorial

Board. Graphical Models, 67(2):CO2, March 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2005:EBl

[Ano05c]  Anonymous.  Editorial

board. Graphical Models, 67(3):CO2, May 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2005:EBm

[Ano05d]  Anonymous.  Editorial

board. Graphical Models, 67(4):CO2, July 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2005:EBn

[Ano05e]  Anonymous.  Editorial

Board. Graphical Models, 67(5):C2, September 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000469.

Anonymous:2005:EBo

[Ano05f]  Anonymous.  Editorial

Board. Graphical Models, 67(6):C2, November 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000603.

Anonymous:2005:VAIe

[Ano05g] Anonymous. Volume author index. Graphical Models, 67(6):622, November 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000640.

Anonymous:2005:VRA

[Ano05h] Anonymous. Volume reviewer acknowledgement list. Graphical Models, 67(6):621, November 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000639.

Anonymous:2006:CEB

[Ano06a] Anonymous. C2 — editorial board. Graphical Models, 68(5–6):??, September/ November 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000701.

Anonymous:2006:EBh

[Ano06b]  Anonymous.  Editorial

Board. Graphical Models, 68(1):C2, January 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000858.

Anonymous:2006:EBi

[Ano06c]  Anonymous.  Editorial

Board. Graphical Models, 68(2):C2, March 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S1524070306000063.

Anonymous:2006:EBj

[Ano06d]  Anonymous.  Editorial

Board. Graphical Models, 68(3):C2, May 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1524070306000257.

Anonymous:2006:EBk

[Ano06e]  Anonymous.  Editorial

board. Graphical Models, 68(4):??, July 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1524070306000452.

Anonymous:2006:SI

[Ano06f] Anonymous. Special issue. Graphical Models, 68(2):65, March 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070305000354.

Anonymous:2006:VRA

[Ano06g] Anonymous. Volume reviewer acknowledgement list. Graphical Models, 68(5–6):496, September/ November 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000786.

Anonymous:2007:CEBd

[Ano07a] Anonymous. C2 — editorial board. Graphical Models, 69(1):??, January 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000889.

Anonymous:2007:CEBe

[Ano07b] Anonymous. C2 — editorial board. Graphical Models, 69(2):??, March 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070307000057.

Anonymous:2007:CEBf

[Ano07c] Anonymous. C2 — editorial board. Graphical Models, 69(3–4):??, May/ July 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070307000185.

Anonymous:2007:CEBg

[Ano07d] Anonymous. C2 — editorial board. Graphical Models, 69(5–6):??, September/ November 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407030700029X.

Anonymous:2007:MCPb

[Ano07e] Anonymous. Most Cited Paper Award. Graphical Models, 69(3–4):159, May/ July 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070307000136.

Anonymous:2007:VAI

[Ano07f] Anonymous. Volume author index. Graphical Models, 69(5–6):276, September/November 2007. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070307000331.

Anonymous:2007:VRA

[Ano07g] Anonymous. Volume reviewer acknowledgement list. Graphical Models, 69(5–6):275, September/ November 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407030700032X.

Anonymous:2008:CEBk

[Ano08a] Anonymous. C2 — editorial board. Graphical Models, 70(1–2):??, January/ March 2008. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070308000076.

Anonymous:2008:CEBl

[Ano08b] Anonymous. C2 — editorial board. Graphical Models, 70(3):??, May 2008. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070308000106.

Anonymous:2008:CEBm

[Ano08c] Anonymous. C2 — editorial board. Graphical Models, 70(4):??, July 2008. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070308000118.

Anonymous:2008:CEBn

[Ano08d] Anonymous. C2 — editorial board. Graphical Models, 70(5):??, September 2008. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407030800012X.

Anonymous:2008:CEBo

[Ano08e] Anonymous. C2 — editorial board. Graphical Models, 70(6):??, November 2008. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070308000131.

Anonymous:2009:AIV

[Ano09a] Anonymous. Author index for volume 71. Graphical Models, 71(6):242, November 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000459.

Anonymous:2009:CEBl

[Ano09b] Anonymous. C2 — editorial board. Graphical Models, 71(4):??, July 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000228.

Anonymous:2009:EBa

[Ano09c]  Anonymous.  Editorial

Board. Graphical Models, 71(1):C2, January 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000083.

Anonymous:2009:EBb

[Ano09d]  Anonymous.  Editorial

Board. Graphical Models, 71(2):C2, March 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070309000113.

Anonymous:2009:EBc

[Ano09e]  Anonymous.  Editorial

Board. Graphical Models, 71(3):C2, May 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070309000162.

Anonymous:2009:EBd

[Ano09f]  Anonymous.  Editorial

Board. Graphical Models, 71(5):??, September 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000241.

Anonymous:2009:EBe

[Ano09g]  Anonymous.  Editorial

Board. Graphical Models, 71(6):??, November 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000629.

Anonymous:2009:RA

[Ano09h] Anonymous. Reviewer acknowledgment. Graphical Models, 71(6):240–241, November 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070309000307.

Anonymous:2010:CEBl

[Ano10a] Anonymous. C2 — editorial board. Graphical Models, 72(3):??, May 2010. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070310000093.

Anonymous:2010:CEBm

[Ano10b] Anonymous. C2 — editorial board. Graphical Models, 72(4):i–ii, July 2010. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070310000111.

Anonymous:2010:EBa

[Ano10c]  Anonymous.  Editorial

Board. Graphical Models, 72(1):??, January 2010. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070310000044.

Anonymous:2010:EBb

[Ano10d]  Anonymous.  Editorial

Board. Graphical Models, 72(2):??, March 2010. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S152407031000007X.

Anonymous:2010:EBc

[Ano10e]  Anonymous.  Editorial

board. Graphical Models, 72(6):i–ii, November 2010. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031000038X.

Anonymous:2010:RA

[Ano10f] Anonymous. Reviewer acknowledgment. Graphical Models, 72(6):74–75, November 2010. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070310000378.

Anonymous:2011:EBl

[Ano11a]  Anonymous.  Editorial

Board. Graphical Models, 73(1):i–ii, January 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070310000445.

Anonymous:2011:EBm

[Ano11b]  Anonymous.  Editorial

Board. Graphical Models, 73(2):i–ii, March 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1524070311000038.

Anonymous:2011:EBn

[Ano11c]  Anonymous.  Editorial

Board. Graphical Models, 73(3):i–ii, May 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1524070311000130.

Anonymous:2011:EBo

[Ano11d]  Anonymous.  Editorial

Board. Graphical Models, 73(5):i–ii, September 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070311000269.

Anonymous:2011:EBp

[Ano11e]  Anonymous.  Editorial

Board. Graphical Models, 73(6):i–ii, November 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070311000300.

Anonymous:2011:RA

[Ano11f] Anonymous. Reviewer acknowledgment. Graphical Models, 73(6):376, November 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070311000440.

Anonymous:2012:EBj

[Ano12a]  Anonymous.  Editorial

Board. Graphical Models, 74(2):i–ii, March 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070312000070.

Anonymous:2012:EBk

[Ano12b]  Anonymous.  Editorial

Board. Graphical Models, 74(4):i–ii, July 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070312000392.

Anonymous:2012:EBl

[Ano12c]  Anonymous.  Editorial

Board. Graphical Models, 74(5):i–ii, September 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000409.

Anonymous:2012:EBp

[Ano12d]  Anonymous.  Editorial

Board. Graphical Models, 74(6):i–ii, November 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000434.

Anonymous:2012:RAb

[Ano12e] Anonymous. Reviewer acknowledgment. Graphical Models, 74(6):374, November 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000641.

Anonymous:2013:EBf

[Ano13a]  Anonymous.  Editorial

Board. Graphical Models, 75(1):i–ii, January 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070313000040.

Anonymous:2013:EBg

[Ano13b]  Anonymous.  Editorial

Board. Graphical Models, 75(2):i–ii, March 2013. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1524070313000106.

Anonymous:2013:EBm

[Ano13c]  Anonymous.  Editorial

Board. Graphical Models, 75(3):i–ii, May 2013. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S152407031300009X.

Anonymous:2013:PNa

[Ano13d] Anonymous. Publisher’s note. Graphical Models, 75 (1):1, January 2013. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070313000052.

Anonymous:2013:RAb

[Ano13e] Anonymous. Reviewer acknowledgment. Graphical Models, 75(6):371, November 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000568.

Anonymous:2013:SCI

[Ano13f] Anonymous. Supplementary content. Graphical Models, 75(2):??, March 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/S152407031200077X#
MMCvFirst.

Anonymous:2014:EBa

[Ano14a]  Anonymous.  Editorial

Board. Graphical Models, 76(3):i–ii, May 2014. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070314000381.

Anonymous:2014:EBb

[Ano14b]  Anonymous.  Editorial

Board. Graphical Models, 76(4):i–ii, July 2014. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070314000393.

Anonymous:2014:EBc

[Ano14c]  Anonymous.  Editorial

Board. Graphical Models, 76(5):i–ii, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000435.

Anonymous:2015:EBa

[Ano15a]  Anonymous.  Editorial

Board. Graphical Models, 77(??):i–ii, January 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000757.

Anonymous:2015:EBb

[Ano15b]  Anonymous.  Editorial

Board. Graphical Models, 78(??):i–ii, March 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000089.

Anonymous:2015:EBc

[Ano15c]  Anonymous.  Editorial

Board. Graphical Models, 79(??):ii–iii, May 2015. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070315000156.

Anonymous:2015:EBd

[Ano15d]  Anonymous.  Editorial

Board. Graphical Models, 80(??):ii–iii, July 2015. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070315000351.

Anonymous:2015:EBe

[Ano15e]  Anonymous.  Editorial

Board. Graphical Models, 81(??):ii–iii, September 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070315000466.

Anonymous:2015:EBf

[Ano15f]  Anonymous.  Editorial

Board. Graphical Models, 82(??):ii–iii, November 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070315000557.

Anonymous:2015:VRA

[Ano15g] Anonymous. Volume reviewer acknowledgement list. Graphical Models, 78(??):60–61, March 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070315000090.

Anonymous:2016:EBa

[Ano16a]  Anonymous.  Editorial

Board. Graphical Models, 83(??):ii–iii, January 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316000126.

Anonymous:2016:EBb

[Ano16b]  Anonymous.  Editorial

Board. Graphical Models, 84(??):ii–iii, March 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070316000230.

Anonymous:2016:EBc

[Ano16c]  Anonymous.  Editorial

Board. Graphical Models, 85(??):ii–iii, May 2016. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070316300145.

Anonymous:2016:EBd

[Ano16d]  Anonymous.  Editorial

Board. Graphical Models, 86(??):ii–iii, July 2016. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070316300236.

Anonymous:2016:EBe

[Ano16e]  Anonymous.  Editorial

Board. Graphical Models, 87(??):ii–iii, September 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070316300327.

Anonymous:2016:EBf

[Ano16f]  Anonymous.  Editorial

Board. Graphical Models, 88(??):ii–iii, November 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316300467.

Anonymous:2016:F

[Ano16g]  Anonymous.  Foreword.

Graphical Models, 83(??): 1, January 2016. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316000084.

Anonymous:2017:EBg

[Ano17a]  Anonymous.  Editorial

Board. Graphical Models, 89(??):ii–iii, January 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070317300073.

Anonymous:2017:EBh

[Ano17b]  Anonymous.  Editorial

Board. Graphical Models, 90(??):ii–iii, March 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070317300176.

Anonymous:2019:N

[Ano19]  Anonymous.  November

2019. Graphical Models, 106(??):??, November 2019. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2020:Ja

[Ano20a] Anonymous. January 2020.

Graphical Models, 107(??): ??, January 2020. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2020:Jb

[Ano20b] Anonymous. July 2020.

Graphical Models, 110(??): ??, July 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2020:Ma

[Ano20c] Anonymous. March 2020.

Graphical Models, 108(??): ??, March 2020. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2020:Mb

[Ano20d] Anonymous. May 2020.

Graphical Models, 109(??): ??, May 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2020:N

[Ano20e]  Anonymous.  November

2020. Graphical Models, 112(??):??, November 2020. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2020:S

[Ano20f] Anonymous. September 2020. Graphical Models, 111 (??):??, September 2020. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2021:Ja

[Ano21a] Anonymous. January 2021. Graphical Models, 113(??): ??, January 2021. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2021:Jb

[Ano21b] Anonymous. July 2021. Graphical Models, 116(??): ??, July 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2021:Ma

[Ano21c] Anonymous. March 2021. Graphical Models, 114(??): ??, March 2021. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2021:Mb

[Ano21d] Anonymous. May 2021. Graphical Models, 115(??): ??, May 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2021:N

[Ano21e]  Anonymous.  November

2021. Graphical Models, 118(??):??, November 2021. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2021:S

[Ano21f] Anonymous. September 2021. Graphical Models, 117 (??):??, September 2021. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2022:Ja

[Ano22a] Anonymous. January 2022.

Graphical Models, 119(??): ??, January 2022. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2022:Jb

[Ano22b] Anonymous. July 2022.

Graphical Models, 122(??): ??, July 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2022:Ma

[Ano22c] Anonymous. March 2022.

Graphical Models, 120(??): ??, March 2022. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2022:Mb

[Ano22d] Anonymous. May 2022. Graphical Models, 121(??): ??, May 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2022:N

[Ano22e]  Anonymous.  November

2022. Graphical Models, 124(??):??, November 2022. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2022:S

[Ano22f] Anonymous. September 2022. Graphical Models, 123 (??):??, September 2022. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2023:A

[Ano23a] Anonymous. April 2023. Graphical Models, 126(??): ??, April 2023. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2023:D

[Ano23b]  Anonymous.  December

2023. Graphical Models, 130(??):??, December 2023. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2023:Ja

[Ano23c] Anonymous. January 2023.

Graphical Models, 125(??): ??, January 2023. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2023:J

[Ano23d] Anonymous. July 2023.

Graphical Models, 128(??): ??, July 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2023:M

[Ano23e] Anonymous. May 2023.

Graphical Models, 127(??): ??, May 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2023:OCM

[Ano23f]  Anonymous.  Obituary:

Christoph M. Hoffmann. Graphical Models, 129(??): ??, October 2023. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000164.

Anonymous:2023:O

[Ano23g] Anonymous. October 2023.

Graphical Models, 129(??): ??, October 2023. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2024:A

[Ano24a] Anonymous. April 2024. Graphical Models, 132(??): ??, April 2024. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2024:F

[Ano24b]  Anonymous.  February

2024. Graphical Models, 131(??):??, February 2024. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Azencot:2003:DSS

[AO03] Jacques Azencot and Maciej Orkisz. Deterministic and stochastic state model of right generalized cylinder (RGC-sm): application in computer phantoms synthesis. Graphical Models, 65(6):323–350, November 2003. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Adanova:2016:BSG

[AT16] V. Adanova and S. Tari. Beyond symmetry groups: a grouping study on Escher’s Euclidean ornaments. Graphical Models, 83(??):15–27, January 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000399.

Attene:2014:DRS

[Att14]  Marco Attene.  Direct

repair of self-intersecting meshes. Graphical Models, 76(6):658–668, November 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000496.

Angelidis:2006:SSD

[AWC06] Alexis Angelidis, Geoff Wyvill, and Marie-Paule Cani. Sweepers: Swept deformation defined by gesture. Graphical Models, 68(1):2–14, January 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407030500055X.

Beckers:2018:ESI

[BAA18] Benoit Beckers, Pierre Alliez, and Daniel Aliaga. Editorial for special issue on “Massive 3D Urban Models”. Graphical Models, 95(??):27–28, January 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300541.

Bayazit:2019:GRG

[Bay19] Ulug Bayazit. A greedy region growing algorithm for anisotropic stretch adaptive triangulation of geometry images. Graphical Models, 106(??):Article 101045, November 2019. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300360.

Brimkov:2011:CMO

[BB11] Valentin E. Brimkov and Reneta P. Barneva. Computational modeling of objects represented in images. Graphical Models, 73(6):311–312, November 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070311000208.

Brimkov:2011:CDB

[BBB11] Valentin E. Brimkov, Reneta P. Barneva, and Boris Brimkov. Connected distance-based rasterization of objects in arbitrary dimension. Graphical Models, 73(6): 323–334, November 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070311000191.

Bardera:2011:MIR

[BBF^{+}11] Anton Bardera, Imma Boada, Miquel Feixas, Jaume Rigau, and Mateu Sbert. Multiresolution image registration based on tree data structures. Graphical Models, 73(4):111–126, July 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031100004X.

Bahrehmand:2017:OLU

[BBM^{+}17] Arash Bahrehmand, Thomas Batard, Ricardo Marques, Alun Evans, and Josep Blat. Optimizing layout using spatial quality metrics and user preferences. Graphical Models, 93(??):25–38, September 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300577.

Blanc-Beyne:2018:SMS

[BBML^{+}18] Thibault Blanc-Beyne, Géraldine Morin, Kathryn Leonard, Stefanie Hahmann, and Axel Carlier. A salience measure for 3D shape decomposition and sub-parts classification. Graphical Models, 99(??):22–30, September 2018. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300353.

Besuievsky:2018:SBG

[BBP18] Gonzalo Besuievsky, Benoit Beckers, and Gustavo Patow. Skyline-based geometric simplification for urban solar analysis. Graphical Models, 95(??):42–50, January 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070317300516.

Bertrand:2015:IBP

[BC15] Gilles Bertrand and Michel Couprie. Isthmus based parallel and symmetric 3D thinning algorithms. Graphical Models, 80(??):1– 15, July 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070315000181.

Bian:2019:ERC

[BDC^{+}19] Shaojun Bian, Zhigang Deng, Ehtzaz Chaudhry, Lihua You, Xiaosong Yang, Lei Guo, Hassan Ugail, Xiaogang Jin, Zhidong Xiao, and Jian Jun Zhang. Efficient and realistic character animation through analytical physics-based skin deformation. Graphical Models, 104(??):Article 101035, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300268.

Barton:2011:SFA

[BE11] Michael Barton and Gershon Elber. Spiral fat arcs

— bounding regions with cubic convergence. Graphical Models, 73(2):50–57, March 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070310000214.

Boschiroli:2012:RBI

[BFRA12] Maria Boschiroli, Christoph Fünfzig, Lucia Romani, and Gudrun Albrecht. G^{1} rational blend interpolatory schemes: a comparative study. Graphical Models, 74(1):29–49, January 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031100066X.

Barbier:2005:FMA

[BGA05] Aurélien Barbier, Eric Galin, and Samir Akkouche. A framework for modeling, animating, and morphing textured implicit models. Graphical Models, 67(3): 166–188, May 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Barequet:2004:CIS

[BGLSS04] Gill Barequet, Michael T. Goodrich, Aya Levi-Steiner, and Dvir Steiner. Contour interpolation by straight skeletons. Graphical Models, 66(4):245–260, July 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Bielser:2004:SMR

[BGTG04] D. Bielser, P. Glardon, M. Teschner, and M. Gross. A state machine for realtime cutting of tetrahedral meshes. Graphical Models, 66(6):398–417, November 2004. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Buse:2016:ECC

[BGZ16] Laurent Busé, André Galligo, and Jiajun Zhang. Extraction of cylinders and cones from minimal point sets. Graphical Models, 86(??):1–12, July 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S152407031630008X.

Bonneau:2014:FIQ

[BH14] Georges-Pierre Bonneau and Stefanie Hahmann. Flexible G^{1} interpolation of quad meshes. Graphical Models, 76(6):669–681, November 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000484.

Barut:2018:CGG

[BHS18] Oner Barut, Murat Haciomeroglu, and Ebru A. Sezer. Combining GPUgenerated linear trajectory segments to create collisionfree paths for real-time ambient crowds. Graphical Models, 99(??):31–45, September 2018. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300341.

Bajaj:2000:CBT

[BIP00] Chandrajit Bajaj, Insung Ihm, and Sanghun Park. Compression-based 3D texture mapping for real-time rendering. Graphical Models, 62(6):391–410, November 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2000.
0532;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0532/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0532/ref.

Buchegger:2014:TCV

[BJK14] Florian Buchegger, Bert Jüttler, and Mario Kapl. Total curvature variation fairing for medial axis regularization. Graphical Models, 76(6):633–647, November 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070314000460.

Birner:2018:BDS

[BJM18] Katharina Birner, Bert Jüttler, and Angelos Mantzaflaris. Bases and dimensions of C^{1}-smooth isogeometric splines on volumetric twopatch domains. Graphical Models, 99(??):46–56, September 2018. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070318300377.

Birner:2019:CBD

[BJM19] Katharina Birner, Bert Jüttler, and Angelos Mantzaflaris. Corrigendum to “Bases and dimensions of C ^{1}- smooth isogeometric splines on volumetric two-patch domains” [graphical models, 99 (2018), 46-56]. Graphical Models, 105 (??):Article 101014, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300420.

Brun:2003:RFW

[BK03] Luc Brun and Walter Kropatsch. Receptive fields within the Combinatorial Pyramid framework. Graphical Models, 65(1–3): 23–42, May 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Barton:2015:SMS

[BKC15] Michael Barton, Jirí Kosinka, and Victor M. Calo. Stretchminimising stream surfaces. Graphical Models, 79(??): 12–22, May 2015. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000041.

Bastl:2015:SBS

[BKL15] Bohumír Bastl, Jiří Kosinka, and Miroslav Lávicka. Simple and branched skins of systems of circles and convex shapes. Graphical Models, 78(??):1–9, March 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000745.

Becker:2020:RDR

[BLSZ20] Martin Becker, Jens Lippel, André Stuhlsatz, and Thomas Zielke. Robust dimensionality reduction for data visualization with deep neural networks. Graphical Models, 108(??):Article 101060, March 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300035.

Baumberg:2005:MCS

[BLT05] Adam Baumberg, Alex Lyons, and Richard Taylor. 3D S.O.M. —- a commercial software solution to 3D scanning. Graphical Models, 67(6):476–495, November 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407030500010X.

Braude:2007:CBS

[BMM^{+}07] Ilya Braude, Jeffrey Marker, Ken Museth, Jonathan Nissanov, and David Breen. Contour-based surface reconstruction using MPU implicit models. Graphical Models, 69(2):139–157, March 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000683.

Biermann:2002:SFM

[BMZB02] Henning Biermann, Ioana M.

Martin, Denis Zorin, and Fausto Bernardini. Sharp features on multiresolution subdivision surfaces. Graphical Models, 64(2): 61–77, March 2002. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Badoual:2017:NSS

[BNR^{+}17] Anaïs Badoual, Paola Novara, Lucia Romani, Daniel Schmitter, and Michael Unser. A non-stationary subdivision scheme for the construction of deformable models with sphere-like topology. Graphical Models, 94(??):38–51, November 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300607.

Boissonnat:2005:PGS

[BO05] Jean-Daniel Boissonnat and Steve Oudot. Provably good sampling and meshing of surfaces. Graphical Models, 67(5):405–451, September 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070305000056.

Bloch:2005:NCS

[BPG05] Isabelle Bloch, Jérémie Pescatore, and Line Garnero. A new characterization of simple elements in a tetrahedral mesh. Graphical Models, 67(4):260–284, July 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Ba:2015:GMA

[BRC15] Wenlan Ba, Ning Ren, and Lixin Cao. Geometry of 3D MAT and its application to moulding surfaces. Graphical Models, 82(??):1–12, November 2015. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070315000429.

Biswas:2004:ADF

[BS04a] Arpan Biswas and Vadim Shapiro. Approximate distance fields with nonvanishing gradients. Graphical Models, 66(3):133–159, May 2004. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Braquelaire:2004:CMR

[BS04b] Achille Braquelaire and Robert Strandh. A color model for rendering linear passive graphic 2D objects. Graphical Models, 66(2):71–88, March 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Bonarrigo:2014:DRU

[BSB14] Francesco Bonarrigo, Alberto Signoroni, and Mario Botsch. Deformable registration using patch-wise shape matching. Graphical Models, 76(5):554–565, September 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000307.

Bartesaghi:2005:TDS

[BSMG05] Alberto Bartesaghi, Guillermo Sapiro, Tom Malzbender, and Dan Gelb. Threedimensional shape rendering from multiple images. Graphical Models, 67(4): 332–346, July 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Barsky:2001:SIP

[BSW01] Brian Barsky, Yoshihisa Shinagawa, and Wenping Wang. Special issue on Pacific Graphics 2000. Graphical Models, 63(4):211, July 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Barsky:2005:EAD

[BTCH05] Brian A. Barsky, Michael J. Tobias, Derrick P. Chu, and Daniel R. Horn. Elimination of artifacts due to occlusion and discretization problems in image space blurring techniques. Graphical Models, 67(6):584–599, November 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070305000093.

Bulbul:2023:PGS

[Bul23] Abdullah Bulbul. Procedural generation of semantically plausible smallscale towns. Graphical Models, 126(??):??, April 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070323000012.

Buzer:2003:LIA

[Buz03] Lilian Buzer. A linear incremental algorithm for naive and standard digital lines and planes recognition. Graphical Models, 65(1–3):61–76, May 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Balmelli:2002:MOU

[BVL02] Laurent Balmelli, Martin Vetterli, and Thomas M.

Liebling. Mesh optimization using global error with application to geometry simplification. Graphical Models, 64(3–4):230–257, May 2002. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Brunton:2014:LDR

[BWW^{+}14] Alan Brunton, Michael Wand, Stefanie Wuhrer, Hans-Peter Seidel, and Tino Weinkauf. A lowdimensional representation for robust partial isometric correspondences computation. Graphical Models, 76(2):70–85, March 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031300060X.

Bashforth:2001:PBE

[BY01] Byron Bashforth and Yee-Hong Yang. Physicsbased explosion modeling. Graphical Models, 63(1):21–44, January 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2000.
0536;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0536/pdf;
http://www.idealibrary.

                                        com/links/doi/10.1006/
                                        gmod.2000.0536/ref.

Camahort:2009:LSA

[CAF09] Emilio Camahort, Francisco Abad, and Don Fussell. A line-space analysis of light-field representations. Graphical Models, 71(5):169–183, September 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070309000101.

Capell:2007:PBR

[CBC^{+}07] Steve Capell, Matthew Burkhart, Brian Curless, Tom Duchamp, and Zoran Popović. Physically based rigging for deformable characters. Graphical Models, 69(1):71–87, January 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070306000634.

Couprie:2003:DO

[CBK03] Michel Couprie, Gilles Bertrand, and Yukiko Kenmochi. Discretization in 2D and 3D orders. Graphical Models, 65(1–3):77–91, May 2003. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Coutinho:2024:PAC

[CC24] Flávio Coutinho and Luiz Chaimowicz. Pixel art character generation as an image-to-image translation problem using GANs. Graphical Models, 132(??): ??, April 2024. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070324000018.

Chen:2001:BQS

[CCF01] Changsong Chen, Falai Chen, and Yuyu Feng. Blending quadric surfaces with piecewise algebraic surfaces. Graphical Models, 63(4):212–227, July 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Chai:2018:SOS

[CCJ^{+}18] Shuangming Chai, Baiyu Chen, Mengyu Ji, Zhouwang Yang, Manfred Lau, Xiao-Ming Fu, and Ligang Liu. Stress-oriented structural optimization for frame structures. Graphical Models, 97(??):80–88, May 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300110.

Chen:2022:PHM

[CCL22] Han Chen, Minghai Chen, and Lin Lu. 3D printed hair modeling from strandlevel hairstyles. Graphical Models, 121(??):??, May 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070322000121.

Chazal:2005:CIA

[CCS05] Frédéric Chazal and David Cohen-Steiner. A condition for isotopic approximation. Graphical Models, 67(5):390–404, September 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000044.

Cabiddu:2023:ESI

[CCS23] Daniela Cabiddu, Gianmarco Cherchi, and Teseo Schneider. Editorial special issue on the 9th Smart Tools and Applications in Graphics Conference (STAG 2022). Graphical Models, 130(??):??, December 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000334.

Comic:2011:DIS

[CD11] Lidija Comić and Leila De Floriani. Dimensionindependent simplification and refinement of Morse complexes. Graphical Models, 73(5):261–285, September 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070311000154.

Chen:2015:SIS

[CDGH15] Falai Chen, Tor Dokken, Thomas Grandine, and Stefanie Hahmann. Special issue of selected papers from the 2014 Dagstuhl seminar on Geometric Modeling. Graphical Models, 82(??):43, November 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000508.

Cerri:2006:RTI

[CFG06] A. Cerri, M. Ferri, and D. Giorgi. Retrieval of trademark images by means of size functions. Graphical Models, 68(5–6):451– 471, September/November 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000592.

Campos:2013:SBS

[CGAY13] Ricard Campos, Rafael Garcia, Pierre Alliez, and Mariette Yvinec. Splatbased surface reconstruction from defect-laden point sets. Graphical Models, 75(6):346–361, November 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070313000258.

Chen:2017:RCR

[CGLX17] Shu-Yu Chen, Lin Gao, Yu-Kun Lai, and Shihong Xia. Rigidity controllable as-rigid-as-possible shape deformation. Graphical Models, 91(??):13–21, May 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070317300139.

Cai:2021:LDF

[CGPZ21] Hongrui Cai, Yudong Guo, Zhuang Peng, and Juyong Zhang. Landmark detection and 3D face reconstruction for caricature using a nonlinear parametric model. Graphical Models, 115(??): ??, May 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070321000084.

Cutler:2007:ADW

[CGW^{+}07] Lawrence D. Cutler, Reid Gershbein, Xiaohuan Corina Wang, Cassidy Curtis, Erwan Maigret, Luca Prasso, and Peter Farson. An art-directed wrinkle system for CG character clothing and skin. Graphical Models, 69(5– 6):219–230, September/ November 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000622.

Collomosse:2006:VMA

[CH06] J. P. Collomosse and P. M.

Hall. Video motion analysis for the synthesis of dynamic cues and futurist art. Graphical Models, 68 (5–6):402–414, September/ November 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000403.

Chekir:2022:DAL

[Che22] Amira Chekir. A deep architecture for log-Euclidean Fisher vector end-to-end learning with application to 3D point cloud classification. Graphical Models, 123(??):??, September 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070322000406.

Chen:2000:VRU

[CK00] Baoquan Chen and Arie Kaufman. 3D volume rotation using shear transformations. Graphical Models, 62(4):308–322, July 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2000.
0525;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0525/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0525/ref.

Choi:2000:ICU

[CL00] Yongchoel Choi and Seungyong Lee. Injectivity conditions of 2D and 3D uniform cubic B-spline functions. Graphical Models, 62(6):411–427, November 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2000.
0531;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0531/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0531/ref.

Chazal:2005:MA

[CL05] Frédéric Chazal and André Lieutier. The “λ-medial axis”. Graphical Models, 67(4):304–331, July 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Cheng:2019:LSI

[CL19] Haonan Cheng and Shiguang Liu. Liquid-solid interaction sound synthesis. Graphical Models, 103 (??):Article 101028, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300190.

Cui:2023:CDH

[CLG^{+}23] Yuan Cui, Moran Li, Yuan Gao, Changxin Gao, Fan Wu, Hao Wen, Jiwei Li, and Nong Sang. Camera distance helps 3D hand pose estimated from a single RGB image. Graphical Models, 127(??):??, May 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000097.

Chen:2018:FHP

[CLL^{+}18] Shuangmin Chen, Bangquan Liu, Taijun Liu, Xiaokang Yu, Shiqing Xin, Ying He, and Changhe Tu. FoldedGI: a highly parallel algorithm for interference detection by folding a geometry image into a 1D buffer. Graphical Models, 100(??):26–32, November 2018. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300298.

Comic:2020:SBC

[CM20] Lidija Comić and Paola Magillo. Surface-based computation of the Euler characteristic in the cubical grid. Graphical Models, 112(??):Article 101093, November 2020. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407032030031X.

Cruz-Matias:2014:NLO

[CMA14] Irving Cruz-Matías and Dolors Ayala. A new lossless orthogonal simplification method for 3D objects based on bounding structures. Graphical Models, 76(4):181–201, July 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000046.

Chica:2012:EGS

[CMB^{+}12] Antoni Chica, Eva Monclús, Pere Brunet, Isabel Navazo, and Àlvar Vinacua. Exampleguided segmentation. Graphical Models, 74(6):302–310, November 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000094.

Cai:2019:DPB

[CMLH19] Jun-Xiong Cai, Tai-Jiang Mu, Yu-Kun Lai, and Shi-Min Hu. Deep pointbased scene labeling with depth mapping and geometric patch feature encoding. Graphical Models, 104(??):Article 101033, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300244.

Comic:2016:CCS

[CN16] Lidija Comić and Benedek Nagy. A combinatorial coordinate system for the body-centered cubic grid. Graphical Models, 87(??):11–22, September 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070316300261.

Comic:2018:DDG

[CN18] Lidija Comić and Benedek Nagy. A description of the diamond grid for topological and combinatorial analysis. Graphical Models, 100(??):33–50, November 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070318300389.

Chen:2001:SIV

[CNK01] Min Chen, Gregory M. Nielson, and Arie E. Kaufman. Special issue on volume modeling. Graphical Models, 63(6):385–386, November 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Cetinaslan:2020:SML

[CO20a] Ozan Cetinaslan and Verónica Orvalho. Sketching manipulators for localized blendshape editing. Graphical Models, 108(??):Article 101059, March 2020. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1524070320300023.

Cetinaslan:2020:SBE

[CO20b] Ozan Cetinaslan and Verónica Orvalho. Stabilized blendshape editing using localized Jacobian transpose descent. Graphical Models, 112(??):Article 101091, November 2020. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300308.

Cani:2009:SS

[CPOO09] Marie-Paule Cani, Fred Pighin, James F. O’Brien, and Carol O’Sullivan. SCA 2006 Symposium. Graphical Models, 71(6):197, November 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000277.

Collomosse:2005:RCS

[CRH05] J. P. Collomosse, D. Rowntree, and P. M. Hall. Rendering cartoon-style motion cues in post-production video. Graphical Models, 67(6):549–564, November 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000111.

Choi:2001:HHD

[CS01] Sung Woo Choi and Hans-Peter Seidel. Hyperbolic Hausdorff distance for medial axis transform. Graphical Models, 63(5):369–384, September 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Chaidee:2018:SLV

[CS18] Supanut Chaidee and Kokichi Sugihara. Spherical Laguerre Voronoi diagram approximation to tessellations without generators. Graphical Models, 95(??):1– 13, January 2018. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070317300632.

Chao:2013:VBP

[CSJ13] Qianwen Chao, Jingjing Shen, and Xiaogang Jin. Video-based personalized traffic learning. Graphical Models, 75(6):305–317, November 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070313000234.

Ceko:2020:DPA

[CSP20] Matthew Ceko, Imants Svalbe, and Timothy Petersen. A discrete projection analogue to Pick’s theorem. Graphical Models, 109(??):Article 101066, May 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300114.

Cerqueira:2020:RRT

[CTAO20] Rômulo Cerqueira, Tiago Trocoli, Jan Albiez, and Luciano Oliveira. A rasterized ray-tracer pipeline for realtime, multi-device sonar simulation. Graphical Models, 111(??):Article 101086, September 2020. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300278.

Cao:2015:TSA

[CTL15] Van-Toan Cao, Trung-Thien Tran, and Denis Laurendeau. A two-stage approach to align two surfaces of deformable objects. Graphical Models, 82(??):13–28, November 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000405.

Chaouch:2009:AM

[CVB09] Mohamed Chaouch and Anne Verroust-Blondet. Alignment of 3D models. Graphical Models, 71(2):63– 76, March 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000046.

Chen:2002:BPR

[CW02] Falai Chen and Wenping Wang. The μ-basis of a planar rational curve– properties and computation. Graphical Models, 64(6):368–381, November 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Choi:2014:CCD

[CWM^{+}14] Yi-King Choi, Wenping Wang, Bernard Mourrain, Changhe Tu, Xiaohong Jia, and Feng Sun. Continuous collision detection for composite quadric models. Graphical Models, 76(5):566–579, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000125.

Cheng:2020:DLP

[CWS^{+}20] Lili Cheng, Zhuo Wei, Mingchao Sun, Shiqing Xin, Andrei Sharf, Yangyan Li, Baoquan Chen, and Changhe Tu. DeepPipes: Learning 3D pipelines reconstruction from point clouds. Graphical Models, 111(??):Article 101079, September 2020. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300230.

Chen:2014:APP

[CXC14] Zhonggui Chen, Yanyang Xiao, and Juan Cao. Approximation by piecewise polynomials on Voronoi tessellation. Graphical Models, 76(5):522–531, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000320.

Chen:2009:CMD

[CXY^{+}09] Xiao-Diao Chen, Gang Xu, Jun-Hai Yong, Guozhao Wang, and Jean-Claude Paul. Computing the minimum distance between a point and a clamped Bspline surface. Graphical Models, 71(3):107–112, May 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000071.

Chen:2023:IIP

[CYF^{+}23] Lan Chen, Jie Yang, Hongbo Fu, Xiaoxu Meng, Weikai Chen, Bo Yang, and Lin Gao. Implicit-PCA : Implicitly-proxied parametric encoding for collision-aware garment reconstruction. Graphical Models, 129(??):??, October 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070323000255.

Che:2004:CSD

[CYW04a] WuJun Che, XunNian Yang, and GuoZhao Wang. Corrigendum to “Skeletondriven 2D distance field metamorphosis using intrinsic shape parameters” [Graphical Models 66 (2004) 102–126]. Graphical Models, 66(4):261, July 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Che:2004:SDD

[CYW04b] WuJun Che, XunNian Yang, and GuoZhao Wang. Skeleton-driven 2D distance field metamorphosis using intrinsic shape parameters. Graphical Models, 66(2): 102–126, March 2004. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Duque:2021:OTP

[DB21] Carlos Arango Duque and Adrien Bartoli. An optimal triangle projector with prescribed area and orientation, application to position-based dynamics. Graphical Models, 118(??):??, November 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000229.

Dewaele:2004:IGL

[DC04] Guillaume Dewaele and Marie-Paule Cani. Interactive global and local deformations for virtual clay. Graphical Models, 66(6): 352–369, November 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

deCarvalho:2016:ICO

[dCBM^{+}16] Felipe Moura de Carvalho, Emilio Vital Brazil, Ricardo Guerra Marroquim, Mario Costa Sousa, and Antonio Oliveira. Interactive cutaways of oil reservoirs. Graphical Models, 84(??):1–14, March 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070316000187.

Deng:2008:PSH

[DCL^{+}08] Jiansong Deng, Falai Chen, Xin Li, Changqi Hu, Weihua Tong, Zhouwang Yang, and Yuyu Feng. Polynomial splines over hierarchical T-meshes. Graphical Models, 70(4):76–86, July 2008. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070308000039.

Daneshpajouh:2012:CPP

[DGZ12] Shervin Daneshpajouh, Mohammad Ghodsi, and Alireza Zarei. Computing polygonal path simplification under area measures. Graphical Models, 74(5):283–289, September 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000264.

Deng:2012:P

[DHK12] Jiansong Deng, Kai Hormann, and Misha Kazhdan. Preface. Graphical Models, 74(4):75, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000355.

Delponte:2006:SMU

[DIOV06] Elisabetta Delponte, Francesco Isgrò, Francesca Odone, and Alessandro Verri. SVDmatching using SIFT features. Graphical Models, 68 (5–6):415–431, September/ November 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000579.

Duan:2013:GEP

[DLP13] Ye Duan, Dong Li, and P. Frank Pai. Geometrically exact physics-based modeling and computer animation of highly flexible 1D mechanical systems. Graphical Models, 75(2):56–68, March 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000039.

DeFloriani:2003:DNM

[DMMP03] Leila De Floriani, Mostefa M.

Mesmoudi, Franco Morando, and Enrico Puppo. Decomposing non-manifold objects in arbitrary dimensions. Graphical Models, 65 (1–3):2–22, May 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Duan:2004:SBD

[DQ04] Ye Duan and Hong Qin.

A subdivision-based deformable model for surface reconstruction of unknown topology. Graphical Models, 66(4):181–202, July 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Du:2005:DPB

[DQ05] Haixia Du and Hong Qin. Dynamic PDE-based surface design using geometric and physical constraints. Graphical Models, 67(1):43– 71, January 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Damiand:2003:SMA

[DR03] Guillaume Damiand and Patrick Resch. Split-and-merge algorithms defined on topological maps for 3D image segmentation. Graphical Models, 65(1–3): 149–167, May 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Dimitrijevic:2023:HPE

[DR23] Aleksandar Dimitrijević and Dejan Rancić. Highperformance ellipsoidal clipmaps. Graphical Models, 130(??): ??, December 2023. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1524070323000395.

Discher:2019:CTW

[DRD19] Sören Discher, Rico Richter, and Jürgen Döllner. Concepts and techniques for web-based visualization and processing of massive 3D point clouds with semantics. Graphical Models, 104(??):Article 101036, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031930027X.

daSilva:2015:NOA

[dSGA15] Josildo Pereira da Silva, Gilson A. Giraldi, and Antônio L. Apolinário, Jr. A new optimization approach for massspring models parameterization. Graphical Models, 81(??):1–17, September 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000375.

Nunes:2022:APC

[dSNJMA22] Mislene da Silva Nunes, Methanias Colaço Júnior, Gastão Florêncio Miranda, Jr., and Beatriz Trinchão Andrade. An approach to preprocess and cluster a BRDF database. Graphical Models, 119(??):??, January 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407032100028X.

Deng:2014:BWA

[DXD14] Jingjing Deng, Xianghua Xie, and Ben Daubney. A bag of words approach to subject specific 3D human pose interaction classification with random decision forests. Graphical Models, 76(3):162–171, May 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070313000337.

Deng:2023:PGP

[DZSW23] Xuan Deng, Cheng Zhang, Jian Shi, and Zizhao Wu. PU-GAT: Point cloud upsampling with graph attention network. Graphical Models, 130(??):??, December 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070323000310.

Eyiyurekli:2017:DPL

[EB17] Manolya Eyiyurekli and David E. Breen. Detailpreserving level set surface editing and geometric texture transfer. Graphical Models, 93(??):39–52, September 2017. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070317300565.

Ezair:2019:UCB

[EE19] Ben Ezair and Gershon Elber. Using curvature bounds towards collision free 5-axis toolpaths. Graphical Models, 103(??):Article 101022, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031930013X.

Elber:2001:CHR

[EKH01] Gershon Elber, Myung-Soo Kim, and Hee-Seok Heo. The convex hull of rational plane curves. Graphical Models, 63(3):151–162, May 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2001.
0546;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2001.0546/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.2001.0546/ref.

Emiris:2015:GOU

[EKK15] Ioannis Z. Emiris, Tatjana Kalinka, and Christos Konaxis. Geometric operations using sparse interpolation matrices. Graphical Models, 82(??):99–109, November 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070315000260.

Elber:2001:CEI

[Elb01]  Gershon Elber.  Curve

evaluation and interrogation on surfaces. Graphical Models, 63(3):197–210, May 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/
                                        doi/10.1006/gmod.2001.
                                        0541;   http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.2001.0541/pdf;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.2001.0541/ref.

Elber:2005:GFB

[Elb05] Gershon Elber. Generalized filleting and blending operations toward functional and decorative applications. Graphical Models, 67(3):189–203, May 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Elber:2005:SMT

[EPB05] Gershon Elber, Nick Patrikalakis, and Pere Brunet. Solid modeling theory and applications. Graphical Models, 67(5):371–372, September 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407030500007X.

Eriksson:2016:FES

[ES16] David Eriksson and Evan Shellshear. Fast exact shortest distance queries for massive point clouds. Graphical Models, 84(??): 28–37, March 2016. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070316000266.

Evako:2011:CSP

[Eva11] Alexander V. Evako. Characterizations of simple points, simple edges and simple cliques of digital spaces: One method of topology-preserving transformations of digital spaces by deleting simple points and edges. Graphical Models, 73(1):1–9, January 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070310000159.

Farouki:2002:ERM

[Far02] Rida T. Farouki. Exact rotation-minimizing frames for spatial Pythagoreanhodograph curves. Graphical Models, 64(6):382–395, November 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Ferley:2001:RAV

[FCG01] Eric Ferley, Marie-Paule Cani, and Jean-Dominique Gascuel. Resolution adaptive volume sculpting. Graphical Models, 63(6): 459–478, November 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Fu:2016:NLI

[FCSF16] Qiang Fu, Xiaowu Chen, Xiaoyu Su, and Hongbo Fu. Natural lines inspired 3D shape re-design. Graphical Models, 85(??):1–10, May 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316000151.

Fu:2020:HCM

[FFY^{+}20] Qiang Fu, Hongbo Fu, Hai Yan, Bin Zhou, Xiaowu Chen, and Xueming Li. Human-centric metrics for indoor scene assessment and synthesis. Graphical Models, 110(??):Article 101073, July 2020. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070320300175.

Furferi:2014:PTM

[FGV^{+}14] Rocco Furferi, Lapo Governi, Yary Volpe, Luca Puggelli, Niccolò Vanni, and Monica Carfagni. From 2D to 2.5D, i.e., from painting to tactile model. Graphical Models, 76(6):706–723, November 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000526.

Farin:2012:AGS

[FH12] Gerald Farin and Dianne Hansford. Agnostic G^{1} Gregory surfaces. Graphical Models, 74(6):346–350, November 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000367.

Fugacci:2019:CDM

[FIF19] Ulderico Fugacci, Federico Iuricich, and Leila De Floriani. Computing discrete Morse complexes from simplicial complexes. Graphical Models, 103 (??):Article 101023, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300141.

Fuhrer:2006:MHP

[FJP06] Martin Fuhrer, Henrik Wann Jensen, and Przemyslaw Prusinkiewicz. Modeling hairy plants. Graphical Models, 68(4):333–342, July 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S152407030600004X.

Fan:2012:SBM

[FML12] Lubin Fan, Min Meng, and Ligang Liu. Sketch-based mesh cutting: a comparative study. Graphical Models, 74(6):292–301, November 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000082.

Fan:2022:TNT

[FMW^{+}22] Yeying Fan, Qian Ma, Guangshun Wei, Zhiming Cui, Yuanfeng Zhou, and Wenping Wang. TAD-Net: tooth axis detection network based on rotation transformation encoding. Graphical Models, 121(??): ??, May 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070322000157.

Favreau:2006:AGV

[FRDC06] Laurent Favreau, Lionel Reveret, Christine Depraz, and Marie-Paule Cani. Animal gaits from video: Comparative studies. Graphical Models, 68(2):212–234, March 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000330.

Fratarcangeli:2007:FMC

[FSF07] Marco Fratarcangeli, Marco Schaerf, and Robert Forchheimer. Facial motion cloning with radial basis functions in MPEG-4 FBA. Graphical Models, 69(2):106–118, March 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000658.

Flotynski:2016:CCS

[FW16] Jakub Flotyński and Krzysztof Walczak. Customization of 3D content with semantic meta-scenes. Graphical Models, 88(??):23–39, November 2016. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070316300182.

Fang:2019:LSA

[FWC^{+}19] Haiyue Fang, Xiaogang Wang, Zheyuan Cai, Yahao Shi, Xun Sun, Shilin Wu, and Bin Zhou. Learning semantic abstraction of shape via 3D region of interest. Graphical Models, 105(??):Article 101038, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300293.

Fadaifard:2013:MFE

[FWH13] Hadi Fadaifard, George Wolberg, and Robert Haralick. Multiscale 3D feature extraction and matching with an application to 3D face recognition. Graphical Models, 75(4):157–176, July 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000076.

Flotynski:2020:CCW

[FWK20] Jakub Floty’nski, Krzysztof Walczak, and Marcin Krzyszkowski. Composing customized web 3D animations with semantic queries. Graphical Models, 107(??):Article 101052, January 2020. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300426.

Feng:2013:CCM

[FWWT13] Xin Feng, Yuanzhen Wang, Yanlin Weng, and Yiying Tong. Compact combinatorial maps: a volume mesh data structure. Graphical Models, 75(3):149–156, May 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000690.

Feng:2018:NMV

[FWX^{+}18] Xiang Feng, Wanggen Wan, Richard Yi Da Xu, Stuart Perry, Song Zhu, and Zexin Liu. A new mesh visual quality metric using saliency weighting-based pooling strategy. Graphical Models, 99(??):1–12, September 2018. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031830033X.

Galin:2000:IPI

[GA00] Eric Galin and Samir Akkouche. Incremental polygonization of implicit surfaces. Graphical Models, 62(1):19–39, January 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.

                                        idealibrary.com/links/
                                        artid/gmod.1999.0514/
                                        production;   http://
                                        www.idealibrary.com/links/
                                        artid/gmod.1999.0514/
                                        production/pdf;   http:
                                        //www.idealibrary.com/
                                        links/artid/gmod.1999.
                                        0514/production/ref.

Gunpinar:2022:AAM

[GC22] Erkan Gunpinar and Serhat Cam. 4 and 5axis additive manufacturing of parts represented using free-form 3D curves. Graphical Models, 120(??): ??, March 2022. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070322000145.

Guo:2018:IGM

[GCZZ18] Kan Guo, Xiaowu Chen, Bin Zhou, and Qinping Zhao. Image-guided 3D model labeling via multiview alignment. Graphical Models, 96(??):30–37, March 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070318300043.

Ge:2022:COR

[GF22] Linlin Ge and Jieqing Feng. Out-of-core outlier removal for large-scale indoor point clouds. Graphical Models, 122(??):??, July 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070322000182.

Gunpinar:2018:SST

[GG18] Erkan Gunpinar and Serkan Gunpinar. A shape sampling technique via particle tracing for CAD models. Graphical Models, 96(??):11–29, March 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300031.

Granier:2003:SLR

[GH03] Xavier Granier and Wolfgang Heidrich. A simple layered RGB BRDF model. Graphical Models, 65(4): 171–184, July 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Grandine:2012:SIS

[GHPW12] Thomas Grandine, Stefanie Hahmann, Jörg Peters, and Wenping Wang. Special issue of selected papers from the 8th Dagstuhl seminar on Geometric Modeling. Graphical Models, 74(6):291, November 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000422.

Gu:2006:MS

[GHQ06] Xianfeng Gu, Ying He, and Hong Qin. Manifold splines. Graphical Models, 68(3):237–254, May 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S152407030600021X.

Gau:2003:MNS

[GK03] C. J. Gau and T. Yung Kong. Minimal non-simple sets in 4D binary images. Graphical Models, 65(1–3): 112–130, May 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Gress:2004:ERE

[GK04] Alexander Greß and Reinhard Klein. Efficient representation and extraction of 2-manifold isosurfaces using kd-trees. Graphical Models, 66(6):370–397, November 2004. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Greiner:2002:SDI

[GKR02] Günther Greiner, Andreas Kolb, and Angela Riepl.

Scattered data interpolation using data dependant optimization techniques. Graphical Models, 64(1):1– 18, January 2002. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Gleicher:2001:CCB

[Gle01] Michael Gleicher. Comparing constraint-based motion editing methods. Graphical Models, 63(2):107–134, March 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2001.
0549;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2001.0549/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.2001.0549/ref.

Guo:2014:CGC

[GLXJ14] Xuekun Guo, Juncong Lin, Kai Xu, and Xiaogang Jin. Creature grammar for creative modeling of 3D monsters. Graphical Models, 76(5):376–389, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000265.

Gao:2017:EHB

[GLY^{+}17] Yang Gao, Shuai Li, Lipeng Yang, Hong Qin, and Aimin Hao. An efficient heat-based model for solid-liquid-gas phase transition and dynamic interaction. Graphical Models, 94(??):14–24, November 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300589.

Ghazanfarpour:2020:PAM

[GMH^{+}20] Anahid Ghazanfarpour, Nicolas Mellado, Chems E. Himeur, Loïc Barthe, and Jean-Pierre Jessel. Proximityaware multiple meshes decimation using quadric error metric. Graphical Models, 109(??):Article 101062, May 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300059.

Gibson:2007:SCS

[GODC07] D. P. Gibson, D. J. Oziem, C. J. Dalton, and N. W. Campbell. A system for the capture and synthesis of insect motion. Graphical Models, 69(5–6):231– 245, September/November 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000671.

Goldman:2011:UQ

[Gol11] Ron Goldman. Understanding quaternions. Graphical Models, 73(2):21–49, March 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070310000172.

Goldman:2013:MPP

[Gol13] Ron Goldman. Modeling perspective projections in 3-dimensions by rotations in 4-dimensions. Graphical Models, 75(2):41–55, March 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000707.

Giannini:2006:SIS

[GP06] Franca Giannini and Alexander Pasko. Special issue: Shape modeling international 2004. Graphical Models, 68(1):1, January 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000585.

Goebbels:2019:TIC

[GPF19] Steffen Goebbels and Regina Pohle-Fröhlich. Techniques for improved CityGML models. Graphical Models, 106(??):Article 101044, November 2019. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300359.

Gravesen:2015:MCS

[Gra15] Jens Gravesen. The metric of colour space. Graphical Models, 82(??):77–86, November 2015. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000247.

Gagvani:2001:AVM

[GS01] Nikhil Gagvani and Deborah Silver. Animating volumetric models. Graphical Models, 63(6):443–458, November 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Goldman:2012:FAQ

[GS12] Ron Goldman and Plamen Simeonov. Formulas and algorithms for quantum differentiation of quantum Bernstein bases and quantum Bézier curves based on quantum blossoming. Graphical Models, 74(6):326–334, November 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000240.

Gallo:2000:FBS

[GSS00] Giovanni Gallo, Michela Spagnuolo, and Salvatore Spinello. Fuzzy B-splines: a surface model encapsulating uncertainty. Graphical Models, 62(1):40–55, January 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
artid/gmod.1999.0512/
production;   http://
www.idealibrary.com/links/
artid/gmod.1999.0512/
production/pdf;   http:
//www.idealibrary.com/
links/artid/gmod.1999.
0512/production/ref.

Guskov:2007:MBA

[Gus07] Igor Guskov. Manifoldbased approach to semiregular remeshing. Graphical Models, 69(1):1–18, January 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000385.

Go:2006:ABI

[GVK06] Jared Go, Thuc D. Vu, and James J. Kuffner. Autonomous behaviors for interactive vehicle animations. Graphical Models, 68(2):90–112, March 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070305000342.

Gregor:2001:ISR

[GW01] J. Gregor and R. T. Whitaker. Indoor scene reconstruction from sets of noisy range images. Graphical Models, 63(5):304–332, September 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Guo:2018:RBM

[GwGgP18] Jie Guo, Yan wen Guo, and Jin gui Pan. A retroreflective BRDF model based on prismatic sheeting and microfacet theory. Graphical Models, 96(??):38–46, March 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070318300018.

Gillmann:2018:ARE

[GWHH18] Christina Gillmann, Thomas Wischgoll, Bernd Hamann, and Hans Hagen. Accurate and reliable extraction of surfaces from image data using a multi-dimensional uncertainty model. Graphical Models, 99(??):13–21, September 2018. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300365.

Gao:2019:NSF

[GWYN19] Yisong Gao, Lifang Wu, Dong-Ming Yan, and Liangliang Nan. Near supportfree multi-directional 3D printing via global-optimal decomposition. Graphical Models, 104(??):Article 101034, ???? 2019. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070319300256.

George:2018:MSM

[GXT18] David George, Xianghua Xie, and Gary KL Tam. 3D mesh segmentation via multi-branch 1D convolutional neural networks. Graphical Models, 96(??):1– 10, March 2018. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031830002X.

Gong:2001:LBM

[GY01] Minglun Gong and Yee-Hong Yang. Layer-based morphing. Graphical Models, 63(1):45–59, January 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/
                                        doi/10.1006/gmod.2000.
                                        0537;   http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.2000.0537/pdf;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.2000.0537/ref.

Gong:2005:CFR

[GY05] Minglun Gong and Yee-Hong Yang. Camera field rendering for static and dynamic scenes. Graphical Models, 67(2):29, March 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Gao:2013:FPS

[GYH13] Zhanheng Gao, Zeyun Yu, and Michael Holst. Featurepreserving surface mesh smoothing via suboptimal Delaunay triangulation. Graphical Models, 75(1):23– 38, January 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000756.

Gao:2020:DPP

[GZL^{+}20] Yang Gao, Zhong Zheng, Jin Li, Shuai Li, Aimin Hao, and Hong Qin. Dynamic particle partitioning SPH model for high-speed fluids simulation. Graphical Models, 109(??):Article 101061, May 2020. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300047.

Hetroy:2003:TQC

[HA03] Franck Hétroy and Dominique Attali. Topological quadrangulations of closed triangulated surfaces using the Reeb graph. Graphical Models, 65(1–3):131–148, May 2003. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Hao:2020:QAF

[Hao20]  Yong-Xia Hao.  Quasi-

area functional for the Plateau–Bézier problem. Graphical Models, 112(??): Article 101095, November 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300321.

Hall:2005:SIV

[HB05] Peter Hall and Brian Barsky. Special issue: Vision, video and graphics 2003. Graphical Models, 67(6):475, November 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070305000305.

Hu:2003:SIP

[HCS03] Shi-Min Hu, Sabine Coquillart, and Heung-Yeung Shum. Special issue on Pacific Graphics 2002. Graphical Models, 65(4):169–170, July 2003. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Hua:2023:PPG

[HD23] Hao Hua and Benjamin Dillenburger. Packing problems on generalised regular grid: Levels of abstraction using integer linear programming. Graphical Models, 130(??):??, December 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070323000358.

Hajij:2016:SSM

[HDL16] Mustafa Hajij, Tamal Dey, and Xin Li. Segmenting a surface mesh into pants using Morse theory. Graphical Models, 88(??):12–21, November 2016. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316300376.

Huang:2016:SGI

[HFH16] Shi-Sheng Huang, Hongbo Fu, and Shi-Min Hu. Structure guided interior scene synthesis via graph matching. Graphical Models, 85(??):46–55, May 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070316300030.

Huard:2014:ISD

[HFSB14] Mathieu Huard, Rida T.

Farouki, Nathalie Sprynski, and Luc Biard. C^{2} interpolation of spatial data subject to arc-length constraints using Pythagorean– hodograph quintic splines. Graphical Models, 76(1):30– 42, January 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000325.

Heo:2001:ITR

[HHS^{+}01] Hee-Seok Heo, Sung Je Hong, Joon-Kyung Seong, Myung-Soo Kim, and Gershon Elber. The intersection of two ringed surfaces and some related problems. Graphical Models, 63(4): 228–244, July 2001. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Hyun:2002:MDA

[HJK02] Dae-Eun Hyun, Bert Jüttler, and Myung-Soo Kim. Minimizing the distortion of affine spline motions. Graphical Models, 64(2):128–144, March 2002. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Huang:2015:RND

[HKL^{+}15] Zhanpeng Huang, Ladislav Kavan, Weikai Li, Pan Hui, and Guanghong Gong. Reducing numerical dissipation in smoke simulation. Graphical Models, 78(??):10–25, March 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070314000769.

Hanniel:2012:CHD

[HKM12] Iddo Hanniel, Adarsh Krishnamurthy, and Sara Mc-Mains. Computing the Hausdorff distance between NURBS surfaces using numerical iteration on the GPU. Graphical Models, 74(4):255–264, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000343.

Huang:2020:WRG

[HKZM20] Jiahui Huang, Zheng-Fei Kuang, Fang-Lue Zhang, and Tai-Jiang Mu. Wall-Net: Reconstructing general room layouts from RGB images. Graphical Models, 111(??):Article 101076, September 2020. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070320300205.

Hu:2021:HPS

[HL21] Chuanfeng Hu and Hongwei Lin. Heterogeneous porous scaffold generation using trivariate B-spline solids and triply periodic minimal surfaces. Graphical Models, 115(??):??, May 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000102.

Hao:2022:CQB

[HL22] Yong-Xia Hao and Ting Li. Construction of quasi-Bézier surfaces from boundary conditions. Graphical Models, 123(??):??, September 2022. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070322000352.

Hu:2023:RRD

[HLL^{+}23] Ling Hu, Qinsong Li, Shengjun Liu, Dong-Ming Yan, Haojun Xu, and Xinru Liu. RFMNet: Robust Deep Functional Maps for unsupervised nonrigid shape correspondence. Graphical Models, 129(??): ??, October 2023. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S152407032300019X.

Hamza-Lup:2016:WGE

[HLM16] Felix G. Hamza-Lup and Marcel Maghiar. Web3D graphics enabled through sensor networks for costeffective assessment and management of energy efficiency in buildings. Graphical Models, 88(??):66–74, November 2016. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316300042.

Huang:2022:MSM

[HLX^{+}22] Xinwei Huang, Nannan Li, Qing Xia, Shuai Li, Aimin Hao, and Hong Qin. Multi-scale and multi-level shape descriptor learning via a hybrid fusion network. Graphical Models, 119(??): ??, January 2022. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000266.

Hu:2013:PSI

[HM13] Shi-Min Hu and Ralph R.

Martin. Preface of special issue on computational visual media. Graphical Models, 75(3):103, May 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000088.

Hernandez-Mederos:2013:GIS

[HMESI13] Victoria Hernández-Mederos, Jorge Estrada-Sarlabous, and Ioannis Ivrissimtzis. Generalization of the incenter subdivision scheme. Graphical Models, 75(2):79– 89, March 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000781.

Hou:2012:CDM

[HQ12a] Tingbo Hou and Hong Qin.

Continuous and discrete Mexican hat wavelet transforms on manifolds. Graphical Models, 74(4):221–232, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/S1524070312000306.

See corrigendum [HQ12b].

Hou:2012:CSD

[HQ12b] Tingbo Hou and Hong Qin. Corrigendum to “Continuous and discrete Mexican hat wavelet transforms on manifolds” [Graphical Models 74 (2012) 221– 232]. Graphical Models, 74(6):373, November 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000410.

See [HQ12a].

Hafez:2023:RWB

[HR23] Omar M. Hafez and Mark M. Rashid. A robust workflow for b-rep generation from image masks. Graphical Models, 128(??):??, July 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407032300005X.

Hoffmann:2005:ASS

[HS05] C. M. Hoffmann and N. F. Stewart. Accuracy and semantics in shape-interrogation applications. Graphical Models, 67(5):373–389, September 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000032.

Hasan:2015:BMS

[HSS15] Mahmudul Hasan, Faramarz F. Samavati, and Mario C. Sousa. Balanced multiresolution for symmetric/antisymmetric filters. Graphical Models, 78(??):36–59, March 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031500003X.

Hubert:2012:CSB

[Hub12] Evelyne Hubert. Convolution surfaces based on polygons for infinite and compact support kernels. Graphical Models, 74(1):1– 13, January 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070311000257.

Hettinga:2019:NGM

[HvBK19] Gerben J. Hettinga, Rowan van Beckhoven, and Jirí Kosinka. Noisy gradient meshes: Augmenting gradient meshes with procedural noise. Graphical Models, 103(??):Article 101024, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070319300153.

Hu:2014:IFC

[HWQ14] Jianping Hu, Xiaochao Wang, and Hong Qin. Improved, feature-centric EMD for 3D surface modeling and processing. Graphical Models, 76(5):340–354, September 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070314000137.

He:2009:HFB

[HXS09] Ying He, Xian Xiao, and Hock-Soon Seah. Harmonic 1-form based skeleton extraction from examples. Graphical Models, 71(2):49– 62, March 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070308000349.

Han:2017:FPS

[HYYZ17] Xiaoguang Han, Hongchuan Yu, Yizhou Yu, and Jianjun Zhang. A fast propagation scheme for approximate geodesic paths. Graphical Models, 91(??):22–29, May 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300115.

Hu:2019:SBA

[HYZ^{+}19] Fei Hu, Long Ye, Wei Zhong, Li Fang, Yun Tie, and Qin Zhang. Semantic based autoencoderattention 3D reconstruction network. Graphical Models, 106(??):Article 101050, November 2019. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300414.

Huang:2020:NHS

[HZL^{+}20] Kemeng Huang, Zipeng Zhao, Chen Li, Changbo Wang, and Hong Qin. Novel hierarchical strategies for SPH-centric algorithms on GPGPU. Graphical Models, 111(??):Article 101088, September 2020. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407032030028X.

Huang:2022:UAH

[HZW^{+}22] Qing Huang, Wen-Xiang Zhang, Qi Wang, Ligang Liu, and Xiao-Ming Fu. Untangling all-hex meshes via adaptive boundary optimization. Graphical Models, 121(??):??, May 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070322000133.

Isenburg:2003:CHV

[IA03] Martin Isenburg and Pierre Alliez. Compressing hexahedral volume meshes. Graphical Models, 65(4): 239–257, July 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Ihrke:2006:AGO

[IM06] Ivo Ihrke and Marcus Magnor. Adaptive grid optical tomography. Graphical Models, 68(5–6):484– 495, September/November 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070306000610.

Iben:2009:GSC

[IO09] Hayley N. Iben and James F. O’Brien. Generating surface crack patterns. Graphical Models, 71(6):198–208, November 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070309000058.

Isenburg:2002:CPM

[IS02] Martin Isenburg and Jack Snoeyink. Compressing the property mapping of polygon meshes. Graphical Models, 64(2):114–127, March 2002. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Irving:2006:THI

[ITF06] G. Irving, J. Teran, and R. Fedkiw. Tetrahedral and hexahedral invertible finite elements. Graphical Models, 68(2):66–89, March 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000317.

James:2009:SCA

[Jam09] Doug James. Symposium on Computer Animation 2008. Graphical Models, 71 (4):125, July 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000186.

Ju:2014:P

[JB14] Tao Ju and Hujun Bao.

Preface. Graphical Models, 76(3):115, May 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070314000058.

Jia:2004:QDC

[JBK04] Jinyuan Jia, George Baciu, and Ki-Wan Kwok. Quadric decomposition for computing the intersections of surfaces of revolution. Graphical Models, 66(5):303–330, September 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Jimenez:2011:TTP

[JFS11] Juan José Jiménez, Francisco Ramón Feito, and Rafael Jesús Segura. Tetratrees properties in graphic interaction. Graphical Models, 73(5):182–201, September 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070311000105.

Jia:2016:PSI

[Jia16] Jinyuan Jia. Preface of the special issue advances in Web3D. Graphical Models, 88(??):22, November 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316300388.

Jeong:2002:DRD

[JK02] Won-Ki Jeong and Chang-Hun Kim. Direct reconstruction of a displaced subdivision surface from unorganized points. Graphical Models, 64(2):78–93, March 2002. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Jiang:2021:BGH

[JLW^{+}21] Jinfeng Jiang, Guiqing Li, Shihao Wu, Huiqian Zhang, and Yongwei Nie. BPA-GAN: Human motion transfer using bodypart-aware generative adversarial networks. Graphical Models, 115(??):??, May 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000126.

Jang:2009:OSS

[JR09] Justin Jang and Jarek Rossignac. OCTOR: Subset selection in recursive pattern hierarchies. Graphical Models, 71(2):92–106, March 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070308000350.

Jiang:2020:RTH

[JSL^{+}20] Jianwei Jiang, Bin Sheng, Ping Li, Lizhuang Ma, Xin Tong, and Enhua Wu. Realtime hair simulation with heptadiagonal decomposition on mass spring system. Graphical Models, 111(??): Article 101077, September 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070320300217.

Jin:2014:CAT

[JSS^{+}14] Yao Jin, Zeyun Shi, Jun Sun, Jin Huang, and Ruofeng Tong. Contentaware texture mapping. Graphical Models, 76(3): 152–161, May 2014. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1524070313000581.

Ji:2021:NMB

[JSZ^{+}21] Zhongping Ji, Xianfang Sun, Yu-Wei Zhang, Weiyin Ma, and Mingqiang Wei. Normal manipulation for bas-relief modeling. Graphical Models, 114(??):Article 101099, March 2021. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1524070321000047.

Jin:2012:UUO

[JWL12] Yong Jin, Qingbiao Wu, and Ligang Liu. Unsupervised upright orientation of man-made models. Graphical Models, 74(4):99–108, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000148.

Jiang:2013:CSE

[JXC^{+}13] Wei Jiang, Kai Xu, Zhi-Quan Cheng, Ralph R. Martin, and Gang Dang. Curve skeleton extraction by coupled graph contraction and surface clustering. Graphical Models, 75(3):137–148, May 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000732.

Jiang:2013:SBI

[JXCZ13] Wei Jiang, Kai Xu, Zhi-Quan Cheng, and Hao Zhang. Skeleton-based intrinsic symmetry detection on point clouds. Graphical Models, 75(4):177–188, July 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000118.

Jiang:2016:SFL

[JYD^{+}16] Haiyong Jiang, Dong-Ming Yan, Weiming Dong, Fuzhang Wu, Liangliang Nan, and Xiaopeng Zhang. Symmetrization of facade layouts. Graphical Models, 85(??):11–21, May 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316000163.

Joo:2014:DGP

[JYTM14] Han Kyul Joo, Tatsuya Yazaki, Masahito Takezawa, and Takashi Maekawa. Differential geometry properties of lines of curvature of parametric surfaces and their visualization. Graphical Models, 76(4):224–238, July 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S152407031400040X.

Jia:2014:MRB

[JZLZ14] Shixiang Jia, Caiming Zhang, Xuemei Li, and Yuanfeng Zhou. Mesh resizing based on hierarchical saliency detection. Graphical Models, 76(5):355–362, September 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070314000198.

Kapllani:2021:TPO

[KAD^{+}21] Levi Kapllani, Chelsea Amanatides, Genevieve Dion, Vadim Shapiro, and David E. Breen. TopoKnit: a process-oriented representation for modeling the topology of yarns in weftknitted textiles. Graphical Models, 118(??):??, November 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000199.

Khodakovsky:2002:NOC

[KADS02] Andrei Khodakovsky, Pierre Alliez, Mathieu Desbrun, and Peter Schröder. Nearoptimal connectivity encoding of 2-manifold polygon meshes. Graphical Models, 64(3–4):147–168, May 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Ko:2001:IWH

[KB01] Hyeong-Seok Ko and Norman I. Badler. The International Workshop on Human Modeling and Animation in Graphical Models. Graphical Models, 63(2):65, March 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2001.
0550;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2001.0550/pdf.

Kopparapu:2001:ENC

[KC01] Sunil Kopparapu and Peter Corke. The effect of noise on camera calibration parameters. Graphical Models, 63(5):277–303, September 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Kang:2014:HBS

[KCD14] Hongmei Kang, Falai Chen, and Jiansong Deng. Hierarchical B-splines on regular triangular partitions. Graphical Models, 76(5):289–300, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070314000095.

Kita:2020:MTD

[KCM20] Naoki Kita, Grégoire Cliquet, and Kazunori Miyata. Mapping two-dimensional plots to a spherical surface using elliptical grid mapping. Graphical Models, 109(??):Article 101067, May 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070320300126.

Ko:2006:SIP

[KCOTW06] Hyeong-Seok Ko, Daniel Cohen-Or, Demetri Terzopoulos, and Joe Warren. Special issue: PG2004. Graphical Models, 68(4): 323, July 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000026.

Kang:2023:NST

[KDCC23] Hongyuan Kang, Xiao Dong, Juan Cao, and Zhonggui Chen. Neural style transfer for 3D meshes. Graphical Models, 129(??):??, October 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000280.

Kim:2014:PCC

[KEK14] Yong-Joon Kim, Gershon Elber, and Myung-Soo Kim. Precise continuous contact motion for planar freeform geometric curves. Graphical Models, 76(5):580–592, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000332.

Kronrod:2001:ECN

[KG01] Boris Kronrod and Craig Gotsman. Efficient coding of nontriangular mesh connectivity. Graphical Models, 63(4):263–275, July 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Kiss:2014:ACM

[KGZ^{+}14] Gábor Kiss, Carlotta Giannelli, Urska Zore, Bert Jüttler, David Großmann, and Johannes Barner. Adaptive CAD model (re- )construction with THBsplines. Graphical Models, 76(5):273–288, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070314000241.

Kim:2013:GIR

[Kim13] Minho Kim. GPU isosurface raycasting of FCC datasets. Graphical Models, 75(2):90–101, March 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S152407031200077X.

Kim:2017:ASG

[Kim17] Minho Kim. Analysis of symmetry groups of boxsplines for evaluation on GPUs. Graphical Models, 93(??):14–24, September 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300553.

Kim:2023:VRB

[KK23] Hyunjun Kim and Minho Kim. Volume reconstruction based on the sixdirection cubic box-spline. Graphical Models, 125(??): ??, January 2023. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070322000443.

Kruppa:2019:ISA

[KKH19] Kinga Kruppa, Roland Kunkli, and Miklós Hoffmann. An improved skinning algorithm for circles and spheres providing smooth transitions. Graphical Models, 101(??):27–37, January 2019. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300444.

Kim:2013:RWB

[KL13] Hyejin Kim and Sung-Hee Lee. Reconstructing whole-body motions with wrist trajectories. Graphical Models, 75(6):328–345, November 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031300026X.

Kallberg:2014:IPL

[KL14] Linus Källberg and Thomas Larsson. Improved pruning of large data sets for the minimum enclosing ball problem. Graphical Models, 76(6):609–619, November 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000447.

Kang:2017:SRA

[KL17] Changgu Kang and Sung-Hee Lee. Scene reconstruction and analysis from motion. Graphical Models, 94(??):25–37, November 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300619.

Kang:2019:BLE

[KL19] Hongmei Kang and Xin Li. de Boor-like evaluation algorithm for analysissuitable T-splines. Graphical Models, 106(??):Article 101042, November 2019. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070319300335.

Kim:2006:VTS

[KLV06] Youngmin Kim, Chang Ha Lee, and Amitabh Varshney. Vertex-transformation streams. Graphical Models, 68(4):371–383, July 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000373.

Krishnan:2000:PTS

[KM00] Shankar Krishnan and Dinesh Manocha. Partitioning trimmed spline surfaces into nonself-occluding regions for visibility computation. Graphical Models, 62(4):283–307, July 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2000.
0526;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0526/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0526/ref.

Kato:2020:MEC

[KM20] Takeo Kato and Taishi Matsumoto. Morphological evaluation of closed planar curves and its application to aesthetic evaluation. Graphical Models, 109(??):Article 101064, May 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070320300072.

Kaufmann:2009:FSD

[KMBG09] Peter Kaufmann, Sebastian Martin, Mario Botsch, and Markus Gross. Flexible simulation of deformable models using discontinuous Galerkin FEM. Graphical Models, 71(4):153–167, July 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070309000125.

Ko:2005:AOP

[KMP05] K. H. Ko, T. Maekawa, and N. M. Patrikalakis. Algorithms for optimal partial matching of free-form objects with scaling effects. Graphical Models, 67(2):29, March 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Karciauskas:2011:RS

[KP11] Kestutis Karčiauskas and Jörg Peters. Rational G^{2} splines. Graphical Models, 73(5):286–295, September 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031100021X.

Karciauskas:2012:FFS

[KP12] Keştutis Karciauskas and Jörg Peters. Free-form splines combining NURBS and basic shapes. Graphical Models, 74(6):351–360, November 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000379.

Karciauskas:2015:ISM

[KP15a] Kestutis Karciauskas and Jörg Peters. Improved shape for multi-surface blends. Graphical Models, 82(??):87–98, November 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000259.

Karciauskas:2015:PAB

[KP15b] Keştutis Karciauskas and Jörg Peters. Pointaugmented biquadratic C^{1} subdivision surfaces. Graphical Models, 77(??):18–26, January 2015. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031400054X.

Kim:2000:SIP

[KS00] Myung-Soo Kim and Hans-Peter Seidel. Special issue on Pacific Graphics ’99 in Graphical Models. Graphical Models, 62(6):389, November 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/
                                        doi/10.1006/gmod.2000.
                                        0533;   http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.2000.0533/pdf.

Kang:2002:ELI

[KS02] Hyung Woo Kang and Sung Yong Shin. Enhanced lane: interactive image segmentation by incremental path map construction. Graphical Models, 64(5):282–303, September 2002. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Knopf:2004:ISD

[KS04] George K. Knopf and Archana Sangole. Interpolating scattered data using 2D self-organizing feature maps. Graphical Models, 66(1):50–69, January 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Kang:2018:TRU

[KSH18] HyeongYeop Kang, Yeram Sim, and JungHyun Han. Terrain rendering with unlimited detail and resolution. Graphical Models, 97(??):64–79, May 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300109.

Kaltheuner:2023:USA

[KSK23] Julian Kaltheuner, Patrick Stotko, and Reinhard Klein. Unified shape and appearance reconstruction with joint camera parameter refinement. Graphical Models, 129(??):??, October 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000231.

Kobbelt:2006:SIS

[KSM^{+}06] Leif Kobbelt, Vadim Shapiro, Botsch Mario, Cazals Frederic, Cohen-Or Danny, Hoppe Hugues, Hu Shimin, Jüttler Bert, Kim Myung-Soo, O’Brien James, Puppo Enrico, Velho Luiz, Wang Wenping, and Zeilfelder Frank. Special issue on SPM 05. Graphical Models, 68(3):235–236, May 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070306000233.

Klein:2000:RSS

[KSS00] Reinhard Klein, Andreas Schilling, and Wolfgang Straßer. Reconstruction and simplification of surfaces from contours. Graphical Models, 62(6):429–443, November 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/
                                        doi/10.1006/gmod.2000.
                                        0530;   http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.2000.0530/pdf;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.2000.0530/ref.

Kye:2008:ICP

[KSS08] Heewon Kye, Byeong-Seok Shin, and Yeong Gil Shin. Interactive classification for pre-integrated volume rendering of high-precision volume data. Graphical Models, 70(6):125–132, November 2008. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070308000088.

Kovacs:2015:AGC

[KVS15] István Kovács, Tamás Várady, and Péter Salvi. Applying geometric constraints for perfecting CAD models in reverse engineering. Graphical Models, 82(??):44–57, November 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000211.

Kronenberger:2015:GCU

[KWFH15] Markus Kronenberger, Oliver Wirjadi, Johannes Freitag, and Hans Hagen. Gaussian curvature using fundamental forms for binary voxel data. Graphical Models, 82(??):123–136, November 2015. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000284.

Wu:2013:CGO

[kWwZ13] Fu kun Wu and Chang wen Zheng. A comprehensive geometrical optics application for wave rendering. Graphical Models, 75(6):318–327, November 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000246.

Wu:2015:MBI

[kWwZ15] Fu kun Wu and Chang wen Zheng. Microfacet-based interference simulation for multilayer films. Graphical Models, 78(??):26–35, March 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070315000028.

Kang:2015:NBP

[KXCD15] Hongmei Kang, Jinlan Xu, Falai Chen, and Jiansong Deng. A new basis for PHTsplines. Graphical Models, 82(??):149–159, November 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070315000302.

Khan:2011:SBA

[KZD^{+}11] Rez Khan, Qin Zhang, Shayan Darayan, Sankari Dhandapani, Sucharit Katyal, Clint Greene, Chandra Bajaj, and David Ress. Surface-based analysis methods for high-resolution functional magnetic resonance imaging. Graphical Models, 73(6):313–322, November 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070310000421.

Kwok:2012:CCB

[KZW12] Tsz-Ho Kwok, Yunbo Zhang, and Charlie C. L. Wang. Constructing common base domain by cues from Voronoi diagram. Graphical Models, 74(4):152–163, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000197.

Lelescu:2004:RCL

[LB04] Dan Lelescu and Frank Bossen. Representation and coding of light field data. Graphical Models, 66(4): 203–225, July 2004. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

LeCallennec:2006:IMD

[LB06] Benoît Le Callennec and Ronan Boulic. Interactive motion deformation with prioritized constraints. Graphical Models, 68(2):175–193, March 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000263.

Li:2004:UAF

[LBM04] Guiqing Li, Hujun Bao, and Weiyin Ma. A unified approach for fairing arbitrary polygonal meshes. Graphical Models, 66(3):160–179, May 2004. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Liu:2002:NVC

[LBSP02] Xinguo Liu, Hujun Bao, Heung-Yeung Shum, and Qunsheng Peng. A novel volume constrained smoothing method for meshes. Graphical Models, 64(3–4):169–182, May 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Lovato:2014:ALA

[LCZG14] Christian Lovato, Umberto Castellani, Carlo Zancanaro, and Andrea Giachetti. Automatic labelling of anatomical landmarks on 3D body scans. Graphical Models, 76(6):648–657, November 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000459.

Lobello:2014:CAI

[LDD14] Ricardo Uribe Lobello, Florent Dupont, and Florence Denis. Out-of-core adaptive iso-surface extraction from binary volume data. Graphical Models, 76(6):593–608, November 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000411.

Lu:2023:JDF

[LDLS23] Zhuheng Lu, Yuewei Dai, Weiqing Li, and Zhiyong Su. Joint data and feature augmentation for selfsupervised representation learning on point clouds. Graphical Models, 129(??): ??, October 2023. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000188.

Liu:2017:EFC

[LDZ^{+}17] Hao Liu, Ning Dai, Baojiang Zhong, Tao Li, and Jun Wang. Extract feature curves on noisy triangular meshes. Graphical Models, 93(??):1–13, September 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300498.

Lawrence:2004:PII

[LF04] Jason Lawrence and Thomas Funkhouser. A painting interface for interactive surface deformations. Graphical Models, 66(6):418–438, November 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Li:2018:CSS

[LFC^{+}18] Manyi Li, Noa Fish, Lili Cheng, Changhe Tu, Daniel Cohen-Or, Hao Zhang, and Baoquan Chen. Classsensitive shape dissimilarity metric. Graphical Models, 98(??):33–42, July 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300328.

Liu:2019:MRP

[LGG19] Mingming Liu, Jie Guo, and Yanwen Guo. 3D model retrieval and pose estimation for indoor images by simulating scene context. Graphical Models, 103(??):Article 101032, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300232.

Laganiere:2006:VRG

[LHM06] R. Laganière, H. Hajjdiab, and A. Mitiche. Visual reconstruction of ground plane obstacles in a sparse view robot environment. Graphical Models, 68(3):282–293, May 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000166.

Lensch:2001:SBA

[LHS01] Hendrik P. A. Lensch, Wolfgang Heidrich, and Hans-Peter Seidel. A silhouettebased algorithm for texture registration and stitching. Graphical Models, 63(4): 245–262, July 2001. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Lee:2000:ESS

[LI00] Rae Kyoung Lee and Insung Ihm. On enhancing the speed of splatting using both object- and imagespace coherence. Graphical Models, 62(4):263–282, July 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2000.
0524;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0524/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0524/ref.

Li:2016:ERS

[LJYL16] Bo Li, Henry Johan, Yuxiang Ye, and Yijuan Lu. Efficient 3D reflection symmetry detection: a viewbased approach. Graphical Models, 83(??):2–14, January 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000417.

Lee:2001:NHM

[LK01] Doo-Won Lee and Hyeong-Seok Ko. Natural hairstyle modeling and animation. Graphical Models, 63(2):67– 85, March 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.idealibrary.com/
links/doi/10.1006/gmod.
2001.0547;   http://www.
idealibrary.com/links/
doi/10.1006/gmod.2001.
0547/pdf;   http://www.
idealibrary.com/links/
doi/10.1006/gmod.2001.
0547/ref.

Lurig:2000:HSD

[LKE00] Christoph Lürig, Leif Kobbelt, and Thomas Ertl. Hierarchical solutions for the deformable surface problem in visualization. Graphical Models, 62(1):2–18, January 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
artid/gmod.1999.0515/
production;   http://
www.idealibrary.com/links/
artid/gmod.1999.0515/
production/pdf;   http:

//www.idealibrary.com/
links/artid/gmod.1999.
0515/production/ref.

Lee:2006:PAB

[LL06] Jehee Lee and Kang Hoon Lee. Precomputing avatar behavior from human motion data. Graphical Models, 68(2):158–174, March 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000275.

Li:2013:GCC

[LL13] Guo Li and Ligang Liu.

Geometry curves: a compact representation for 3D shapes. Graphical Models, 75(5):265–278, September 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000179.

Liu:2022:JPA

[LLH22] Chengzhi Liu, Juncheng Li, and Lijuan Hu. Jacobi-PIA algorithm for bicubic B-spline interpolation surfaces. Graphical Models, 120(??):??, March 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407032200011X.

Lai:2019:ADM

[LLL19a] Po Kong Lai, Weizhe Liang, and Robert Laganière. Additive depth maps, a compact approach for shape completion of single view depth maps. Graphical Models, 104(??):Article 101030, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300219.

Li:2019:FCM

[LLL^{+}19b] Tao Li, Wei Liu, Hao Liu, Jun Wang, and Ligang Liu. Feature-convinced mesh denoising. Graphical Models, 101(??):17–26, January 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300456.

Liu:2018:GSS

[LLW^{+}18] Xiuping Liu, Liping Lin, Jun Wu, Weiming Wang, Baocai Yin, and Charlie C. L. Wang. Generating sparse self-supporting wireframe models for 3D printing using mesh simplification. Graphical Models, 98(??):14–23, July 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300286.

Lin:2013:NIS

[LLXW13] Shujin Lin, Xiaonan Luo, Songhua Xu, and Jianmin Wang. A new interpolation subdivision scheme for triangle/quad mesh. Graphical Models, 75(5):247–254, September 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031300012X.

Lachaud:2000:CAD

[LM00] Jacques-Olivier Lachaud and Annick Montanvert. Continuous analogs of digital boundaries: a topological approach to isosurfaces. Graphical Models, 62(3):129–164, May 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2000.
0522;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0522/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0522/ref.

Lai:2012:VLO

[LM12] Yu-Kun Lai and Ralph R.

Martin. Vertex location optimisation for improved remeshing. Graphical Models, 74(4):233–243, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000318.

Li:2020:FBM

[LML^{+}20] Kun Li, Yali Mao, Yunke Liu, Ruizhi Shao, and Yebin Liu. Full-body motion capture for multiple closely interacting persons. Graphical Models, 110(??):Article 101072, July 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070320300163.

Li:2014:SEL

[LQJ^{+}14] Zhong Li, Shengwei Qin, Xiaogang Jin, Zeyun Yu, and Jiao Lin. Skeletonenhanced line drawings for 3D models. Graphical Models, 76(6):620–632, November 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070314000472.

Lehmann:2012:NCT

[LR12] Nicole Lehmann and Ulrich Reif. Notes on the curvature tensor. Graphical Models, 74(6):321–325, November 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000239.

Lee:2001:CIA

[LS01] Jehee Lee and Sung Yong Shin. A coordinateinvariant approach to multiresolution motion analysis. Graphical Models, 63(2):87–105, March 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2001.
0548;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2001.0548/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.2001.0548/ref.

Lavor:2021:OIC

[LSA21] Carlile Lavor, Michael Souza, and José Luis Aragón. Orthogonality of isometries in the conformal model of the 3D space. Graphical Models, 114(??):Article 101100, March 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000059.

Li:2023:ECD

[LSY^{+}23a] Xingxin Li, Shibo Song, Junfeng Yao, Hanyin Zhang, Rongzhou Zhou, and Qingqi Hong. Efficient collision detection using hybrid medial axis transform and BVH for rigid body simulation. Graphical Models, 128(??): ??, July 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000103.

Lyu:2023:MMD

[LSY23b] Bowen Lyu, Li-Yong Shen, and Chun-Ming Yuan. MixNet: Mix different networks for learning 3D implicit representations. Graphical Models, 129(??): ??, October 2023. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000206.

Lachaud:2003:ICD

[LV03] Jacques-Olivier Lachaud and Anne Vialard. 10th International Conference on Discrete Geometry for Computer Imagery: Discrete topology and geometry for image and object representation. Graphical Models, 65(1–3):1, May 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Lobaz:2014:HLB

[LV14] P. Lobaz and L. Vása. Hierarchical Laplacian-based compression of triangle meshes. Graphical Models, 76(6):682–690, November 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000502.

Lluch:2004:MTS

[LVM04] J. Lluch, R. Vivó, and C. Monserrat. Modelling tree structures using a single polygonal mesh. Graphical Models, 66(2):89–101, March 2004. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Liu:2021:SIT

[LWG^{+}21] Zhihao Liu, Kai Wu, Jianwei Guo, Yunhai Wang, Oliver Deussen, and Zhanglin Cheng. Single image tree reconstruction via adversarial network. Graphical Models, 117(??):??, September 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000205.

Li:2023:HFP

[LWG^{+}23] Ren-Wu Li, Bo Wang, Lin Gao, Ling-Xiao Zhang, and Chun-Peng Li. Highfidelity point cloud completion with low-resolution recovery and noise-aware upsampling. Graphical Models, 126(??):??, April 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000048.

Liu:2008:SAB

[LWGP08] Shiguang Liu, Zhangye Wang, Zheng Gong, and Qunsheng Peng. Simulation of atmospheric binary mixtures based on two-fluid model. Graphical Models, 70(6):117–124, November 2008. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070308000064.

Linden:2015:LSP

[LWH15] Sven Linden, Andreas Wiegmann, and Hans Hagen. The LIR space partitioning system applied to the Stokes equations. Graphical Models, 82(??):58–66, November 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000223.

Li:2018:PSF

[LWZ^{+}18] Chen Li, Changbo Wang, Shenfan Zhang, Sheng Qiu, and Hong Qin. Porescale flow simulation in anisotropic porous material via fluid-structure coupling. Graphical Models, 95(??):14–26, January 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300644.

Li:2019:CSF

[LXD19] Chong-Jun Li, Lin-Lin Xie, and Wen-Bin Du. Curve and surface fitting based on the nonhomogeneous linear differential system. Graphical Models, 103(??):Article 101026, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300177.

Liu:2021:VNV

[LXH21] Yilin Liu, Ke Xie, and Hui Huang. VGF-Net: Visualgeometric fusion learning for simultaneous drone navigation and height mapping. Graphical Models, 116 (??):??, July 2021. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070321000138.

Liu:2016:LWV

[LXTJ16] Xiaojun Liu, Ning Xie, Kai Tang, and Jinyuan Jia. Lightweighting for Web3D visualization of large-scale BIM scenes in real-time. Graphical Models, 88(??):40–56, November 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316300170.

Li:2019:HCM

[LXX^{+}19] Shuai Li, Zhijun Xie, Qing Xia, Aimin Hao, and Hong Qin. Hybrid 4D cardiovascular modeling based on patient-specific clinical images for real-time PCI surgery simulation. Graphical Models, 101(??):1–7, January 2019. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070318300432.

Liao:2014:SAG

[LXZ14] Tao Liao, Guoliang Xu, and Yongjie Jessica Zhang. Structure-aligned guidance estimation in surface parameterization using eigenfunctionbased cross field. Graphical Models, 76(6):691–705, November 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000514.

Liu:2008:SSB

[LYCG08] Xin Liu, Hongxun Yao, Xilin Chen, and Wen Gao. Shape from silhouettes based on a centripetal pentahedron model. Graphical Models, 70(6):133–148, November 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407030800009X.

Lee:2012:CCS

[LYKL12] Sun-Young Lee, Jong-Chul Yoon, Ji-Yong Kwon, and In-Kwon Lee. Cartoon-Modes: Cartoon stylization of video objects through modal analysis. Graphical Models, 74(2):51–60, March 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000033.

Lee:2010:TCV

[LYL10] Sun-Young Lee, Jong-Chul Yoon, and In-Kwon Lee. Temporally coherent video matting. Graphical Models, 72(3):25–33, May 2010. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070310000081.

Liu:2016:UOS

[LZL16] Zishun Liu, Juyong Zhang, and Ligang Liu. Upright orientation of 3D shapes with convolutional networks. Graphical Models, 85(??):22–29, May 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316000199.

Luo:2015:IDS

[LZX^{+}15] Ran Luo, Lifeng Zhu, Weiwei Xu, Patrick Kelley, Vanessa Svihla, and Yin Yang. Interactive design and simulation of tubular supporting structure. Graphical Models, 80(??): 16–30, July 2015. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070315000193.

Liu:2014:GCR

[LZY^{+}14] Yuan Liu, Wen Zhou, Zhouwang Yang, Jiansong Deng, and Ligang Liu. Globally consistent rigid registration. Graphical Models, 76(5):542–553, September 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000290.

Martens:2023:DSC

[MB23] Jan Martens and Jörg Blankenbach. A decomposition scheme for continuous Level of Detail, streaming and lossy compression of unordered point clouds. Graphical Models, 130(??):??, December 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000383.

Marras:2012:MBM

[MBH^{+}12] Stefano Marras, Michael M.

Bronstein, Kai Hormann, Riccardo Scateni, and Roberto Scopigno. Motionbased mesh segmentation using augmented silhouettes. Graphical Models, 74(4):164–172, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000215.

McDonnell:2005:DSB

[MCQ05] Kevin T. McDonnell, Yu-Sung Chang, and Hong Qin. DigitalSculpture: a subdivision-based approach to interactive implicit surface modeling. Graphical Models, 67(4):347–369, July 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Mao:2019:EBS

[MDLH19] Yan Mao, Xuemei Du, Yongjian Li, and Wu He. An emotion based simulation framework for complex evacuation scenarios. Graphical Models, 102(??): 1–9, March 2019. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070319300013.

Machchhar:2017:GCD

[ME17] Jinesh Machchhar and Gershon Elber. ϵ-guarantee of a covering of 2D domains using random-looking curves. Graphical Models, 89(??):1– 13, January 2017. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316300510.

Medeiros:2014:FAB

[MIPS14] Esdras Medeiros, Lis Ingrid, Sinésio Pesco, and Claudio Silva. Fast adaptive blue noise on polygonal surfaces. Graphical Models, 76(1):17–29, January 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000313.

MacLachlan:2016:EMM

[MJ16] Lynne MacLachlan and Iestyn Jowers. Exploration of multi-material surfaces as weighted shapes. Graphical Models, 83(??):28–36, January 2016. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000387.

Maekawa:2002:SCF

[MK02] T. Maekawa and K. H.

Ko. Surface construction by fitting unorganized curves. Graphical Models, 64(5): 316–332, September 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Marinov:2005:OMS

[MK05] Martin Marinov and Leif Kobbelt. Optimization methods for scattered data approximation with subdivision surfaces. Graphical Models, 67(5):452–473, September 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000068.

Mwalongo:2016:GBR

[MKB^{+}16] Finian Mwalongo, Michael Krone, Michael Becher, Guido Reina, and Thomas Ertl. GPU-based remote visualization of dynamic molecular data on the web. Graphical Models, 88(??):57–65, November 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070316300066.

Mizrahi:2017:MSC

[MKH^{+}17] Jonathan Mizrahi, Sijoon Kim, Iddo Hanniel, Myung Soo Kim, and Gershon Elber. Minkowski sum computation of Bspline surfaces. Graphical Models, 91(??):30–38, May 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300127.

Mihalef:2008:ITP

[MKS^{+}08] Viorel Mihalef, Samet Kadioglu, Mark Sussman, Dimitris Metaxas, and Vassilios Hurmusiadis. Interaction of two-phase flow with animated models. Graphical Models, 70(3):33–42, May 2008. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070307000264.

Malgouyres:2000:TPW

[ML00] Rémy Malgouyres and Alexandre Lenoir. Topology preservation within digital surfaces. Graphical Models, 62(2):71–84, March 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.1999.
0517;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.1999.0517/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.1999.0517/ref.

Ma:2017:RSR

[MLC^{+}17] Ming Ma, Na Lei, Wei Chen, Kehua Su, and Xianfeng Gu. Robust surface registration using optimal mass transport and Teichmüller mapping. Graphical Models, 90(??):13–23, March 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300024.

Ma:2012:VNG

[MLF^{+}12] Teng Ma, Xiang Long, Lu Feng, Pei Luo, and Zhuangzhi Wu. Visible neighborhood graph of point clouds. Graphical Models, 74(4):184–196, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000276.

Ma:2019:SSU

[MM19] Yue Ma and Weiyin Ma. A subdivision scheme for unstructured quadrilateral meshes with improved convergence rate for isogeometric analysis. Graphical Models, 106(??):Article 101043, November 2019. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070319300347.

Musoni:2023:GPL

[MMC23] Pietro Musoni, Simone Melzi, and Umberto Castellani. GIM3D plus: a labeled 3D dataset to design data-driven solutions for dressed humans. Graphical Models, 129(??):??, October 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070323000176.

Ma:2019:SPC

[MMNG19] Ming Ma, Joseph Marino, Saad Nadeem, and Xianfeng Gu. Supine to prone colon registration and visualization based on optimal mass transport. Graphical Models, 104(??):Article 101031, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300220.

Moenning:2007:MGS

[MMS^{+}07] C. Moenning, F. Mémoli, G. Sapiro, N. Dyn, and N. A. Dodgson. Meshless geometric subdivision. Graphical Models, 69(3–4):160–179, May/ July 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000956.

Muhlthaler:2003:CMS

[MP03] Heidrun Mühlthaler and Helmut Pottmann. Computing the Minkowski sum of ruled surfaces. Graphical Models, 65(6):369–384, November 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Martinez:2013:SRO

[MPV13] Jonàs Martínez, Núria Pla, and Marc Vigo. Skeletal representations of orthogonal shapes. Graphical Models, 75(4):189–207, July 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070313000155.

Maximo:2011:RRI

[MPVF11] A. Maximo, R. Patro, A. Varshney, and R. Farias. A robust and rotationally invariant local surface descriptor with applications to non-local mesh processing. Graphical Models, 73(5):231–242, September 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070311000166.

Mateer:2005:VBP

[MR05] J. W. Mateer and J. A. Robinson. A vision-based postproduction tool for footage logging, analysis, and annotation. Graphical Models, 67(6):565–583, November 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070305000123.

Mahmoudi:2009:TDP

[MS09] Mona Mahmoudi and Guillermo Sapiro. Three-dimensional point cloud recognition via distributions of geometric distances. Graphical Models, 71(1):22–31, January 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070308000313.

Michikawa:2010:SGD

[MS10] Takashi Michikawa and Hiromasa Suzuki. Sparse grid distance transforms. Graphical Models, 72(4):35– 45, July 2010. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031000010X.

Molina:2023:TST

[MV23] Elena Molina and Pere-Pau Vázquez. Two-step techniques for accurate selection of small elements in VR environments. Graphical Models, 128(??):??, July 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000139.

Mitra:2017:ESI

[MYC17] Niloy J. Mitra, Yizhou Yu, and Ming-Ming Cheng. Editorial special issue on the Fifth Computational Visual Media conference (CVM 2017). Graphical Models, 91 (??):12, May 2017. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070317300462.

Natali:2011:GBR

[NBPF11] Mattia Natali, Silvia Biasotti, Giuseppe Patanè, and Bianca Falcidieno. Graphbased representations of point clouds. Graphical Models, 73(5):151–164, September 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070311000099.

Natsupakpong:2010:DEP

[NÇ10] Suriya Natsupakpong and M. Cenk Çavusoğlu. Determination of elasticity parameters in lumped element (mass-spring) models of deformable objects. Graphical Models, 72(6):61–73, November 2010. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070310000147.

Neff:2006:MEE

[NF06] Michael Neff and Eugene Fiume. Methods for exploring expressive stance. Graphical Models, 68(2):133–157, March 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000329.

Nyul:2002:FCI

[NFU02] László G. Nyúl, Alexandre X. Falcão, and Jayaram K. Udupa. Fuzzyconnected 3D image segmentation at interactive speeds. Graphical Models, 64(5):259–281, September 2002. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Nie:2016:EFL

[Nie16] Jianhui Nie. Extracting feature lines from point clouds based on smooth shrink and iterative thinning. Graphical Models, 84(??):38–49, March 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070316300054.

Nie:2017:ARG

[Nie17] Jianhui Nie. An algorithm for the rapid generation of bas-reliefs based on point clouds. Graphical Models, 94(??):1–13, November 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300590.

Nemeth:2011:TCI

[NKP11] Gábor Németh, Péter Kardos, and Kálmán Palágyi. Thinning combined with iteration-by-iteration smoothing for 3D binary images. Graphical Models, 73(6):335–345, November 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070311000063.

Nguyen:2014:ISP

[NPJ14] Dang-Manh Nguyen, Michael Pauley, and Bert Jüttler. Isogeometric segmentation. Part II: On the segmentability of contractible solids with non-convex edges. Graphical Models, 76(5):426–439, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000204.

Nguyen:2023:LBI

[NRW23] Andrew-Hieu Nguyen, Olivia Rees, and Zhaoyang Wang. Learning-based 3D imaging from single structuredlight image. Graphical Models, 126(??):??, April 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000024.

Nie:2021:BRG

[NSL^{+}21] Jianhui Nie, Wenkai Shi, Ye Liu, Hao Gao, Feng Xu, and Zhaochen Zhang. Bas-relief generation from point clouds based on normal space compression with real-time adjustment on CPU. Graphical Models, 113(??):Article 101096, January 2021. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000011.

Naitsat:2018:GBD

[NSZ18] Alexander Naitsat, Emil Saucan, and Yehoshua Y. Zeevi. Geometry-based distortion measures for space deformation. Graphical Models, 100(??):12–25, November 2018. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300407.

Ohtake:2005:SDI

[OBS05] Yutaka Ohtake, Alexander Belyaev, and Hans-Peter Seidel. 3D scattered data interpolation and approximation with multilevel compactly supported RBFs. Graphical Models, 67(3):150–165, May 2005.

CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Ohtake:2006:CAM

[OBS06a] Yutaka Ohtake, Alexander Belyaev, and Hans-Peter Seidel. A composite approach to meshing scattered data. Graphical Models, 68(3):255–267, May 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070306000221.

Ohtake:2006:SSR

[OBS06b] Yutaka Ohtake, Alexander Belyaev, and Hans-Peter Seidel. Sparse surface reconstruction with adaptive partition of unity and radial basis functions. Graphical Models, 68(1):15–24, January 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070305000548.

Ong:2006:LIK

[OH06] Eng-Jon Ong and Adrian Hilton. Learnt inverse kinematics for animation synthesis. Graphical Models, 68 (5–6):472–483, September/ November 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000609.

Oh:2007:DPF

[OK07] Seungtaik Oh and Bon Ki Koo. Data perturbation for fewer triangles in marching tetrahedra. Graphical Models, 69(3–4):211–218, May/ July 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070307000148.

Pandzic:2003:FMC

[Pan03] Igor S. Pandzic. Facial motion cloning. Graphical Models, 65(6):385–404, November 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Parvez:2023:FPP

[Par23] Mohammad Tanvir Parvez.

Fast progressive polygonal approximations for online strokes. Graphical Models, 129(??):??, October 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000309.

Pasko:2001:CHM

[PASS01] Alexander Pasko, Valery Adzhiev, Benjamin Schmitt, and Christophe Schlick. Constructive hypervolume modeling. Graphical Models, 63(6):413–442, November 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Peitso:2020:DLN

[PB20] Loren Peitso and Don Brutzman. Defeating lag in network-distributed physics simulations. Graphical Models, 111(??):Article 101075, September 2020. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070320300199.

Pereira:2011:SBW

[PBM^{+}11] Thiago Pereira, Emilio Vital Brazil, Ives Macêdo, Mario Costa Sousa, Luiz Henrique de Figueiredo, and Luiz Velho. Sketch-based warping of RGBN images. Graphical Models, 73(4):97– 110, July 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031000041X.

Pradal:2009:PPB

[PBN^{+}09] C. Pradal, F. Boudon, C. Nouguier, J. Chopard, and C. Godin. PlantGL: a Python-based geometric library for 3D plant modelling at different scales.

Graphical Models, 71(1):1– 21, January 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070308000143.

Popescu:2006:M

[PBSM06] V. Popescu, G. Bahmutov, E. Sacks, and M. Mudure. The ModelCamera. Graphical Models, 68(5–6):385– 401, September/November 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000397.

Plante:2002:CCH

[PCP02] Eric Plante, Marie-Paule Cani, and Pierre Poulin. Capturing the complexity of hair motion. Graphical Models, 64(1):40–58, January 2002. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Peres:2023:NHD

[PCPM23] Victor Peres, Esteban Clua, Thiago Porcino, and Anselmo Montenegro. Nonhomogeneous denoising for virtual reality in real-time path tracing rendering. Graphical Models, 129(??): ??, October 2023. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070323000140.

Ponchio:2019:RCA

[PCS19] Federico Ponchio, Massimiliano Corsini, and Roberto Scopigno. RELIGHT: a compact and accurate RTI representation for the web. Graphical Models, 105(??):Article 101040, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070319300311.

Ponchio:2016:MFD

[PD16] Federico Ponchio and Matteo Dellepiane. Multiresolution and fast decompression for optimal web-based rendering. Graphical Models, 88(??):1–11, November 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316300285.

Picinbono:2003:NLA

[PDA03] Guillaume Picinbono, Hervé Delingette, and Nicholas Ayache. Non-linear anisotropic elasticity for real-time surgery simulation. Graphical Models, 65(5):305–321, September 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Peternell:2000:GPB

[Pet00] Martin Peternell. Geometric properties of bisector surfaces. Graphical Models, 62(3):202–236, May 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.1999.
0521;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.1999.0521/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.1999.0521/ref.

Pasko:2011:PFB

[PFV^{+}11] Alexander Pasko, Oleg Fryazinov, Turlif Vilbrandt, Pierre-Alain Fayolle, and Valery Adzhiev. Procedural function-based modelling of volumetric microstructures. Graphical Models, 73(5):165–181, September 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070311000087.

Paier:2023:ULS

[PHE23] Wolfgang Paier, Anna Hilsmann, and Peter Eisert. Unsupervised learning of style-aware facial animation from real acting performances. Graphical Models, 129(??):??, October 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070323000292.

Peng:2023:GGG

[PKL^{+}23] Tao Peng, Jiewen Kuang, Jinxing Liang, Xinrong Hu, Jiazhe Miao, Ping Zhu, Lijun Li, Feng Yu, and Minghua Jiang. GSNet: Generating 3D garment animation via graph skinning network. Graphical Models, 129(??):??, October 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070323000279.

Park:2012:MST

[PLL12] Min Ki Park, Seung Joo Lee, and Kwan H. Lee. Multi-scale tensor voting for feature extraction from unstructured point clouds. Graphical Models, 74(4):197–208, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000288.

Pickup:2018:ECF

[PLS^{+}18] David Pickup, Juncheng Liu, Xianfang Sun, Paul L. Rosin, Ralph R. Martin, Zhiquan Cheng, Zhouhui Lian, Sipin Nie, Longcun Jin, Gil Shamai, Yusuf Sahillioğlu, and Ladislav Kavan. An evaluation of canonical forms for nonrigid 3D shape retrieval. Graphical Models, 97(??): 17–29, May 2018. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300055.

Pueyo:2019:OGT

[PPP19] O. Pueyo, X. Pueyo, and G. Patow. An overview of generalization techniques for street networks. Graphical Models, 106(??): Article 101049, November 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300402.

Pulli:2000:SRD

[PS00] Kari Pulli and Linda G.

Shapiro. Surface reconstruction and display from range and color data. Graphical Models, 62(3):165–201, May 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/

                                        doi/10.1006/gmod.1999.
                                        0519;   http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.1999.0519/pdf;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.1999.0519/ref.

Paris:2003:RAI

[PS03] Sylvain Paris and François Sillion. Robust acquisition of 3D informations from short image sequences. Graphical Models, 65(4): 222–238, July 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Pasko:2005:SSI

[PS05] Alexander Pasko and Michela Spagnuolo. SMI 2003 special issue. Graphical Models, 67(3):149, May 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Peternell:2007:MSB

[PS07] Martin Peternell and Tibor Steiner. Minkowski sum boundary surfaces of 3Dobjects. Graphical Models, 69(3–4):180–190, May/ July 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070307000021.

Patane:2007:FCG

[PSF07] G. Patanè, M. Spagnuolo, and B. Falcidieno. Families of cut-graphs for bordered meshes with arbitrary genus. Graphical Models, 69(2):119–138, March 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407030600066X.

Page:2002:NVV

[PSK^{+}02] D. L. Page, Y. Sun, A. F.

Koschan, J. Paik, and M. A. Abidi. Normal vector voting: Crease detection and curvature estimation on large, noisy meshes. Graphical Models, 64(3–4): 199–229, May 2002. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Pang:2014:EQD

[PSL14] Xufang Pang, Zhan Song, and Rynson W. H. Lau. An effective quad-dominant meshing method for unorganized point clouds. Graphical Models, 76(2):86– 102, March 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000611.

Pan:2015:CAI

[PT15] Rongjiang Pan and Gabriel Taubin. Color adjustment in image-based texture maps. Graphical Models, 79(??):39–48, May 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000119.

Park:2008:PAU

[PY08] Youngsup Park and Kyunghyun Yoon. Painterly animation using motion maps. Graphical Models, 70(1– 2):1–15, January/March 2008. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407030700015X.

Quynh:2012:IAC

[QHXC12] Dao T. P. Quynh, Ying He, Shi-Qing Xin, and Zhonggui Chen. An intrinsic algorithm for computing geodesic distance fields on triangle meshes with holes. Graphical Models, 74(4):209–220, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031200029X.

Qian:2019:LLT

[QSS^{+}19] Yinling Qian, Jian Shi, Hanqiu Sun, Lei Ma, Yanyun Chen, Qiong Wang, and Pheng-Ann Heng. Layered leaf texturing using structure-guided model. Graphical Models, 103 (??):Article 101029, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300207.

Qin:2002:EPP

[QY02] Xuejie Qin and Yee-Hong Yang. Estimating parameters for procedural texturing by genetic algorithms. Graphical Models, 64(1):19– 39, January 2002. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Qi:2023:GGB

[QZZ^{+}23] Yunjia Qi, Chen Zong, Yunxiao Zhang, Shuangmin Chen, Minfeng Xu, Lingqiang Ran, Jian Xu, Shiqing Xin, and Ying He. GBGVD: Growth-based geodesic Voronoi diagrams. Graphical Models, 129(??): ??, October 2023. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000267.

Rodgman:2006:RVG

[RC06] David Rodgman and Min Chen. Refraction in volume graphics. Graphical Models, 68(5–6):432– 450, September/November 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000580.

Ripolles:2009:RCL

[RCG^{+}09] Oscar Ripolles, Miguel Chover, Jesus Gumbau, Francisco Ramos, and Anna Puig-Centelles. Rendering continuous level-of-detail meshes by Masking Strips. Graphical Models, 71(5):184–195, September 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407030900023X.

Ragragui:2018:RMB

[RCHS18] Anouar Ragragui, Adnane Ouazzani Chahdi, Akram Halli, and Khalid Satori. Revolution mapping with bump mapping support. Graphical Models, 100(??):1–11, November 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300390.

Rodriguez:2011:CCL

[RCVA11] Jorge Ernesto Rodríguez, Irving Cruz, Eduard Vergés, and Dolors Ayala. A connected-component-labeling-based approach to virtual porosimetry. Graphical Models, 73(5):296–310, September 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031100018X.

Richard:2011:DDR

[RFLSA11] A. Richard, L. Fuchs, G. Largeteau-Skapin, and E. Andres. Decomposition of nD-rotations: Classification, properties and algorithm. Graphical Models, 73(6):346–353, November 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070311000221.

Raposo:2012:MAB

[RG12] Adriano N. Raposo and Abel J. P. Gomes. 3D molecular assembling of B-DNA sequences using nucleotides as building blocks. Graphical Models, 74(4):244–254, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000331.

Ramalho:2023:RUT

[RJG23] Bárbara Ramalho, Joaquim Jorge, and Sandra Gama. Representing uncertainty through sentiment and stance visualizations: a survey. Graphical Models, 129(??):??, October 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070323000218.

Robles-Kelly:2005:ESR

[RKH05] Antonio Robles-Kelly and Edwin R. Hancock. Estimating the surface radiance function from single images. Graphical Models, 67(6):518–548, November 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070305000135.

Rajagopalan:2000:LHF

[RKK^{+}00] A. N. Rajagopalan, K. Sunil Kumar, Jayashree Karlekar, R. Manivasakan, M. Milind Patil, U. B. Desai, P. G. Poonacha, and S. Chaudhuri. Locating human faces in a cluttered scene. Graphical Models, 62(5):323–342, September 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.1999.
0511;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.1999.0511/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.1999.0511/ref.

Rosin:2013:AMR

[RL13] Paul L. Rosin and Yu-Kun Lai. Artistic minimal rendering with lines and blocks. Graphical Models, 75(4):208–229, July 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000143.

Rodrigues:2019:GVR

[RMCdST19] Daniele C. Uchoa Maia Rodrigues, Felipe A. Moura, Sergio Augusto Cunha, and Ricardo da S. Torres. Graph visual rhythms in temporal network analyses. Graphical Models, 103(??):Article 101021, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300128.

Rolland-Neviere:2013:RDB

[RNDA13] Xavier Rolland-Nevière, Gwenaël Doërr, and Pierre Alliez. Robust diameterbased thickness estimation of 3D objects. Graphical Models, 75(6):279–296, November 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070313000192.

Rossignac:2010:GCP

[Ros10] Jarek Rossignac. GMOD: Creation, processing, animation, visualization, and dissemination of graphical models. Graphical Models, 72(1):iii–v, January 2010. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070310000032.

Rahbar:2008:ILC

[RP08] Kambiz Rahbar and Hamid Reza Pourreza. Inside looking out camera pose estimation for virtual studio. Graphical Models, 70(4):57–75, July 2008. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070308000027.

Rueda:2004:RCP

[RSFdM04] A. J. Rueda, R. J. Segura, F. R. Feito, and J. Ruiz de Miras. Rasterizing complex polygons without tessellations. Graphical Models, 66(3):127–132, May 2004. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Riaz:2015:MRU

[RTKW15] Qaiser Riaz, Guanhong Tao, Björn Krüger, and Andreas Weber. Motion reconstruction using very few accelerometers and ground contacts. Graphical Models, 79(??):23–38, May 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000107.

Ruminski:2020:LSD

[RW20] Dariusz Rumiński and Krzysztof Walczak. Largescale distributed semantic augmented reality services

— a performance evaluation. Graphical Models, 107 (??):Article 101027, January 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300189.

Shellshear:2013:PPD

[SBA13] Evan Shellshear, Fadi Bitar, and Ulf Assarsson. PDQ: Parallel Distance Queries for deformable meshes. Graphical Models, 75(2):69–78, March 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000027.

Sbert:2000:OAP

[Sbe00]  Mateu Sbert.  Opti-

mal absorption probabilities for random walk radiosity. Graphical Models, 62(1):56–70, January 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
artid/gmod.1999.0513/
production;   http://
www.idealibrary.com/links/
artid/gmod.1999.0513/
production/pdf;   http:
//www.idealibrary.com/
links/artid/gmod.1999.
0513/production/ref.

Schofield:2023:ISS

[SBSG23] Sam Schofield, Andrew Bainbridge-Smith, and Richard Green. An improved semisynthetic approach for creating visual–inertial odometry datasets. Graphical Models, 126(??):??, April 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000036.

Spagnuolo:2009:SSI

[SCOG09] Michela Spagnuolo, Daniel Cohen-Or, and Xianfeng David Gu. SMI 2008 special issue. Graphical Models, 71(2):33, March 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000137.

Sivignon:2004:DIM

[SDC04] Isabelle Sivignon, Florent Dupont, and Jean-Marc Chassery. Digital intersections: minimal carrier, connectivity, and periodicity properties. Graphical Models, 66(4):226–244, July 2004. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Schmitter:2017:CSS

[SFB^{+}17] D. Schmitter, J. Fageot, A. Badoual, P. Garcia-Amorena, and M. Unser. Compactly-supported smooth interpolators for shape modeling with varying resolution. Graphical Models, 94(??):52–64, November 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300620.

Sun:2021:WCD

[SFCD21] Xiaopeng Sun, Jia Fu, Teng Chen, and Yu Dong. Wrinkle and curl distortion of leaves using plant dynamic. Graphical Models, 118(??):??, November 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000230.

Senasli:2000:RVL

[SGHM00] M. Senasli, L. Garnero, A. Herment, and E. Mousseaux. 3D reconstruction of vessel lumen from very few angiograms by dynamic contours using a stochastic approach. Graphical Models, 62(2):105–127, March 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.1999.
0520;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.1999.0520/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.1999.0520/ref.

Sarris:2001:BTD

[SGS01] Nikos Sarris, Nikos Grammalidis, and Michael G. Strintzis. Building three dimensional head models. Graphical Models, 63(5): 333–368, September 2001.

CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Sewall:2009:VSS

[SGTL09] Jason Sewall, Nico Galoppo, Georgi Tsankov, and Ming Lin. Visual simulation of shockwaves. Graphical Models, 71(4):126–138, July 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000150.

Starck:2005:VVS

[SH05] J. Starck and A. Hilton.

Virtual view synthesis of people from multiple view video sequences. Graphical Models, 67(6):600–620, November 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000159.

Sheffer:2003:SM

[She03] Alla Sheffer. Skinning 3D meshes. Graphical Models, 65(5):274–285, September 2003. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Shen:2012:DTA

[SJ12] Jingjing Shen and Xiaogang Jin. Detailed traffic animation for urban road networks. Graphical Models, 74(5):265–282, September 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000227.

Stolte:2001:NTR

[SK01] Nilo Stolte and Arie Kaufman. Novel techniques for robust voxelization and visualization of implicit surfaces. Graphical Models, 63(6):387–412, November 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Sahillioglu:2021:SAI

[SK21] Yusuf Sahillioglu and Ladislav Kavan. Scale-adaptive ICP. Graphical Models, 116(??): ??, July 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000187.

Silva:2024:ECC

[SKMM24] Gabriel Fonseca Silva, Paulo Ricardo Knob, Rubens Halbig Montanha, and Soraia Raupp Musse. Evaluating and comparing crowd simulations: Perspectives from a crowd authoring tool. Graphical Models, 131(??):??, February 2024. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000425.

Shu:2000:EMC

[SLB^{+}00] Huazhong Shu, Limin Luo, Xudong Bao, Wenxue Yu, and Guoniu Han. An efficient method for computation of Legendre moments. Graphical Models, 62(4):237–262, July 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2000.
0523;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0523/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0523/ref.

Shekhawat:2023:AGF

[SLD^{+}23] Krishnendra Shekhawat, Rohit Lohani, Chirag Dasannacharya, Sumit Bisht, and Sujay Rastogi. Automated generation of floorplans with non-rectangular rooms. Graphical Models, 127(??):??, May 2023. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070323000061.

Sheng:2018:ENI

[SLF^{+}18] Bin Sheng, Ping Li, Hongbo Fu, Lizhuang Ma, and Enhua Wu. Efficient non-incremental constructive solid geometry evaluation for triangular meshes. Graphical Models, 97(??):1– 16, May 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070318300067.

Son:2011:SGD

[SLKL11] Minjung Son, Yunjin Lee, Henry Kang, and Seungyong Lee. Structure grid for directional stippling. Graphical Models, 73(3):74– 87, May 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070310000433.

Segall:2014:LAF

[SMKE14] Aviv Segall, Jonathan Mizrahi, Yong Joon Kim, and Gershon Elber. Line accessibility of free form surfaces. Graphical Models, 76(5):301–311, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070314000216.

Seo:2004:EBA

[SMT04] Hyewon Seo and Nadia Magnenat-Thalmann. An example-based approach to human body manipulation. Graphical Models, 66(1):1– 23, January 2004. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Semwal:2001:SFU

[SO01] Sudhanshu K. Semwal and Jun Ohya. Spatial filtering using the active-space indexing method. Graphical Models, 63(3):135–150, May 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2001.
0540;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2001.0540/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.2001.0540/ref.

Steinemann:2009:SMD

[SOG09] Denis Steinemann, Miguel A.

Otaduy, and Markus Gross. Splitting meshless deforming objects with explicit surface tracking. Graphical Models, 71(6):209–220, November 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000034.

Saha:2000:DTS

[SR00] Punam K. Saha and Azriel Rosenfeld. The digital topology of sets of convex voxels. Graphical Models, 62(5):343–352, September 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/
                                        doi/10.1006/gmod.2000.
                                        0527;   http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.2000.0527/pdf;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.2000.0527/ref.

Suzuki:2002:SIN

[SRK02a] Hiromasa Suzuki, Alyn Rockwood, and Leif Kobbelt. Special issue on the Ninth Pacific Graphics Conference (PG 2001). Graphical Models, 64(2):59–60, March 2002. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Szymczak:2002:PRM

[SRK02b] Andrzej Szymczak, Jarek Rossignac, and Davis King. Piecewise regular meshes: Construction and compression. Graphical Models, 64 (3–4):183–198, May 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Sun:2009:NAS

[SRML09] Xianfang Sun, Paul L. Rosin, Ralph R. Martin, and Frank C. Langbein. Noise analysis and synthesis for 3D laser depth scanners. Graphical Models, 71(2):34– 48, March 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070308000337.

Shamir:2006:SBS

[SS06] Ariel Shamir and Amir Shaham. Skeleton based solid representation with topology preservation. Graphical Models, 68(3):307–321, May 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000779.

Sadeghi:2011:SRL

[SS11] Javad Sadeghi and Faramarz F. Samavati. Smooth reverse loop and Catmull– Clark subdivision. Graphical Models, 73(5):202–217, September 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070311000117.

Seylan:2019:STM

[SS19]  Çaglar Seylan and Yusuf

Sahillioglu. 3D skeleton transfer for meshes and clouds. Graphical Models, 105(??):Article 101041, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300323.

Saba:2014:CBB

[SSHS14] Marianna Saba, Teseo Schneider, Kai Hormann, and Riccardo Scateni. Curvature-based blending of closed planar curves. Graphical Models, 76(5): 263–272, September 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070314000319.

Song:2017:ACO

[SSL17] Mofei Song, Zhengxing Sun, and Hongyan Li. Accumulative categorization: Online 3D shape classification for progressive collections. Graphical Models, 89(??):14–27, January 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300012.

Sakkalis:2001:TGP

[SSP01a] T. Sakkalis, G. Shen, and N. M. Patrikalakis. Topological and geometric properties of interval solid models. Graphical Models, 63(3):163–175, May 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2001.
0539;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2001.0539/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.2001.0539/ref.

Shen:2001:BRM

[SSP01b] G. Shen, T. Sakkalis, and N. M. Patrikalakis. Boundary representation model rectification. Graphical Models, 63(3):177–195, May 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2001.
0543;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2001.0543/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.2001.0543/ref.

Shao:2007:AP

[ST07] Wei Shao and Demetri Terzopoulos. Autonomous pedestrians. Graphical Models, 69(5–6):246– 274, September/November 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070307000252.

Stahlhut:2005:ENT

[Sta05] O. Stahlhut. Extending natural textures with multiscale synthesis. Graphical Models, 67(6):496–517, November 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070305000147.

Stojanovic:2019:SOS

[STRD19] Vladeta Stojanovic, Matthias Trapp, Rico Richter, and Jürgen Döllner. Serviceoriented semantic enrichment of indoor point clouds using octree-based multiview classification. Graphical Models, 105 (??):Article 101039, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S152407031930030X.

Shi:2021:LSD

[SWC^{+}21] Min Shi, Yukun Wei, Lan Chen, Dengming Zhu, Tianlu Mao, and Zhaoqi Wang. Learning a shared deformation space for efficient design-preserving garment transfer. Graphical Models, 115(??):??, May 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000114.

Schmeisser:2015:SCB

[SWHH15] Andre Schmeißer, Raimund Wegener, Dietmar Hietel, and Hans Hagen. Smooth convolution-based distance functions. Graphical Models, 82(??):67–76, November 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000235.

Sun:2014:ERU

[SWW^{+}14] Xiaopeng Sun, Guan Wang, Lu Wang, Hongyan Sun, and Xiaopeng Wei. 3D ear recognition using local salience and principal manifold. Graphical Models, 76(5):402–412, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000101.

Shen:2018:HPA

[SWYH18] Wanqiang Shen, Guozhao Wang, Yang Yang, and Aihua Hu. Hexagonal pyramid algorithm over a triangular domain. Graphical Models, 100(??):51–60, November 2018. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070318300419.

Shen:2021:CCH

[SZJ^{+}21] Zeyu Shen, Mingyang Zhao, Xiaohong Jia, Yuan Liang, Lubin Fan, and Dong-Ming Yan. Combining convex hull and directed graph for fast and accurate ellipse detection. Graphical Models, 116(??):??, July 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070321000151.

Taubin:2002:DMR

[Tau02a] Gabriel Taubin. Dual mesh resampling. Graphical Models, 64(2):94–113, March 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Taubin:2002:SIP

[Tau02b] Gabriel Taubin. Special issue on processing of large polygonal meshes. Graphical Models, 64(3–4):145– 146, May 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Teng:2007:CTM

[TCH07] Chin-Hung Teng, Yung-Sheng Chen, and Wen-Hsing Hsu. Constructing a 3D trunk model from two images. Graphical Models, 69(1):33–56, January 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000427.

Tang:2020:PHM

[TCL^{+}20] Jing Tang, Pengcheng Cui, Bin Li, Yaobin Zhang, and Hang Si. Parallel hybrid mesh adaptation by refinement and coarsening. Graphical Models, 111(??): Article 101084, September 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300266.

Theobalt:2004:CFF

[TCMS04] Christian Theobalt, Joel Carranza, Marcus A. Magnor, and Hans-Peter Seidel. Combining 3D flow fields with silhouette-based human motion capture for immersive video. Graphical Models, 66(6):333–351, November 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Tamstorf:2013:DBF

[TG13] Rasmus Tamstorf and Eitan Grinspun. Discrete bending forces and their Jacobians. Graphical Models, 75(6):362–370, November 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000209.

Tolani:2000:RTI

[TGB00] Deepak Tolani, Ambarish Goswami, and Norman I. Badler. Real-time inverse kinematics techniques for anthropomorphic limbs. Graphical Models, 62(5):353–388, September 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

tronic). URL http://www.
idealibrary.com/links/
doi/10.1006/gmod.2000.
0528;   http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0528/pdf;
http://www.idealibrary.
com/links/doi/10.1006/
gmod.2000.0528/ref.

Tsuchie:2012:SMD

[TH12] Shoichi Tsuchie and Masatake Higashi. Surface mesh denoising with normal tensor framework. Graphical Models, 74(4):130–139, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000173.

Taimouri:2014:DSM

[TH14] Vahid Taimouri and Jing Hua. Deformation similarity measurement in quasiconformal shape space. Graphical Models, 76(2):57– 69, March 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000623.

Thurmer:2003:CCD

[Thü03]  Grit Thürmer.  Closed

curves in n-dimensional discrete space. Graphical Models, 65(1–3):43–60, May 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

terHaar:2009:FMF

[tHV09] Frank B. ter Haar and Remco C. Veltkamp. A 3D face matching framework for facial curves. Graphical Models, 71(2):77–91, March 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070308000325.

Takacs:2012:RPS

[TJ12] T. Takacs and B. Jüttler.

H^{2} regularity properties of singular parameterizations in isogeometric analysis. Graphical Models, 74(6):361–372, November 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000380.

Thurey:2009:DPF

[TKPR09] N. Thürey, R. Keiser, M. Pauly, and U. Rüde. Detail-preserving fluid control. Graphical Models, 71(6):221–228, November 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000022.

Tang:2005:OAF

[TL05] Kai Tang and Yong-Jin Liu. An optimization algorithm for free-form surface partitioning based on weighted Gaussian image. Graphical Models, 67(1):17–42, January 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Tarini:2005:AMO

[TLGS05] Marco Tarini, Hendrik P. A. Lensch, Michael Goesele, and Hans-Peter Seidel. 3D acquisition of mirroring objects using striped patterns. Graphical Models, 67(4): 233–259, July 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Terra:2007:PBT

[TM07] Sílvio César Lizana Terra and Ronald Anthony Metoyer. A performance-based technique for timing keyframe animations. Graphical Models, 69(2):89–105, March 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000646.

Tang:2010:MMC

[TMT10] Min Tang, Dinesh Manocha, and Ruofeng Tong. MCCD: Multi-core collision detection between deformable models using front-based decomposition. Graphical Models, 72(2):7–23, March 2010. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070310000020.

Tereshin:2021:HFR

[TPFA21] A. Tereshin, A. Pasko, O. Fryazinov, and V. Adzhiev. Hybrid function representation for heterogeneous objects. Graphical Models, 114(??):Article 101098, March 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000035.

Tukur:2023:SFP

[TPG^{+}23] M. Tukur, G. Pintore, E. Gobbetti, J. Schneider, and M. Agus. SPIDER: a framework for processing, editing and presenting immersive high-resolution spherical indoor scenes. Graphical Models, 128(??): ??, July 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070323000127.

Tobor:2006:RMS

[TRS06] Ireneusz Tobor, Patrick Reuter, and Christophe Schlick. Reconstructing multi-scale variational partition of unity implicit surfaces with attributes. Graphical Models, 68(1):25– 41, January 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070305000573.

Tschinkel:2020:RDU

[Tsc20]  Gerwald Tschinkel.  A

recommendation dashboard utilising micro-filter visualisations: an evaluation and investigation of applicability on mobile devices. Graphical Models, 110(??):Article 101074, July 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300187.

Takahashi:2004:TVS

[TTF04] Shigeo Takahashi, Yuriko Takeshima, and Issei Fujishiro. Topological volume skeletonization and its application to transfer function design. Graphical Models, 66(1):24–49, January 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Untereiner:2013:DMR

[UCB13] Lionel Untereiner, David Cazier, and Dominique Bechmann. n-dimensional multiresolution representation of subdivision meshes with arbitrary topology. Graphical Models, 75(5): 231–246, September 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070313000131.

Unzueta:2008:FBP

[UPBS08] Luis Unzueta, Manuel Peinado, Ronan Boulic, and Ángel Suescun. Fullbody performance animation with Sequential Inverse Kinematics. Graphical Models, 70(5):87–104, September 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070308000040.

Vasa:2011:OMT

[Vás11] Libor Vása. Optimised mesh traversal for dynamic mesh compression. Graphical Models, 73(5):218–230, September 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070311000129.

Varga:2011:DDB

[VBN11] László Varga, Péter Balázs, and Antal Nagy. Directiondependency of binary tomographic reconstruction algorithms. Graphical Models, 73(6):365–375, November 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070311000245.

Vacavant:2009:FDI

[VCT09] Antoine Vacavant, David Coeurjolly, and Laure Tougne. A framework for dynamic implicit curve approximation by an irregular discrete approach. Graphical Models, 71(3):113–124, May 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000149.

Varadhan:2006:AMS

[VM06] Gokul Varadhan and Dinesh Manocha. Accurate Minkowski sum approximation of polyhedral models. Graphical Models, 68(4):343–355, July 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000191.

Vijayalakshmi:2023:SAE

[VN23] D. Vijayalakshmi and Malaya Kumar Nath. A systematic approach for enhancement of homogeneous background images using structural information. Graphical Models, 130(??):??, December 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407032300036X.

Vigo:2012:EAB

[VPAM12] Marc Vigo, Núria Pla, Dolors Ayala, and Jonàs Martínez. Efficient algorithms for boundary extraction of 2D and 3D orthogonal pseudomanifolds. Graphical Models, 74(3):61– 74, May 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000112.

Venkataramanan:2022:DDA

[VRP22] Aishwarya Venkataramanan, Antoine Richard, and Cédric Pradalier. A data driven approach to generate realistic 3D tree barks. Graphical Models, 123(??):??, September 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S152407032200042X.

Vanderhyde:2008:TSI

[VS08] James Vanderhyde and Andrzej Szymczak. Topological simplification of isosurfaces in volumetric data using octrees. Graphical Models, 70(1–2):16–31, January/March 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070307000161.

Varady:2012:TSI

[VSR12] Tamás Várady, Péter Salvi, and Alyn Rockwood. Transfinite surface interpolation with interior control. Graphical Models, 74(6):311–320, November 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000100.

Wendt:2007:FVF

[WBOL07] Jeremy D. Wendt, William Baxter, Ipek Oguz, and Ming C. Lin. Finite volume flow simulations on arbitrary domains. Graphical Models, 69(1):19–32, January 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000415.

Weng:2014:RTF

[WCHZ14] Yanlin Weng, Chen Cao, Qiming Hou, and Kun Zhou. Real-time facial animation on mobile devices. Graphical Models, 76(3):172–179, May 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000295.

Welling:2006:RVF

[WEY06] Joel S. Welling, William F.

Eddy, and Terence K. Young. Rotation of 3D volumes by fourierinterpolated shears. Graphical Models, 68(4):356–370, July 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070305000949.

Wang:2015:QRS

[WG15] Xuhui Wang and Ron Goldman. Quaternion rational surfaces: Rational surfaces generated from the quaternion product of two rational space curves. Graphical Models, 81(??):18–32, September 2015. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070314000289.

Wang:2019:SRM

[WG19] Haohao Wang and Ron Goldman. Surfaces of revolution with moving axes and angles. Graphical Models, 106(??):Article 101047, November 2019. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070319300384.

Weber:2012:SFP

[WHHB12] Christopher Weber, Stefanie Hahmann, Hans Hagen, and Georges-Pierre Bonneau. Sharp feature preserving MLS surface reconstruction based on local feature line approximations. Graphical Models, 74(6):335–345, November 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031200032X.

Wang:2002:CQS

[WJG02] Wenping Wang, Barry Joe, and Ronald Goldman. Computing quadric surface intersections based on an analysis of plane cubic curves. Graphical Models, 64(6):335–367, November 2002. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Wang:2014:SMS

[WLAT14] Hao Wang, Tong Lu, Oscar Kin-Chung Au, and Chiew-Lan Tai. Spectral 3D mesh segmentation with a novel single segmentation field. Graphical Models, 76(5):440–456, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000356.

Wang:2014:RPS

[WLL14] Chunxue Wang, Zheng Liu, and Ligang Liu. Asrigid-as-possible spherical parametrization. Graphical Models, 76(5):457–467, September 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S152407031400023X.

Wu:2022:PBM

[WLLL22] Xiaoqun Wu, Xiaoyun Lin, Nan Li, and Haisheng Li. Patch-based mesh inpainting via low rank recovery. Graphical Models, 122(??): ??, July 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070322000169.

Wang:2014:NSB

[WLT14] Feng Wang, Lanfen Lin, and Min Tang. A new sketch-based 3D model retrieval approach by using global and local features. Graphical Models, 76(3):128–139, May 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070313000593.

Wuhrer:2015:FEB

[WLTS15] Stefanie Wuhrer, Jochen Lang, Motahareh Tekieh, and Chang Shu. Finite element based tracking of deforming surfaces. Graphical Models, 77(??):1–17, January 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000538.

Wang:2006:STI

[WLW06] Wencheng Wang, Kuiyu Li, and Enhua Wu. Stick textures for image-based rendering. Graphical Models, 68(3):294–306, May 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000154.

Wang:2022:DFM

[WMC^{+}22] Hui Wang, Bitao Ma, Junjie Cao, Xiuping Liu, and Hui Huang. Deep functional maps for simultaneously computing direct and symmetric correspondences of 3D shapes. Graphical Models, 123(??):??, September 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407032200039X.

Wu:2014:UNP

[WMR^{+}14] J. Wu, R. R. Martin, P. L. Rosin, X.-F. Sun, Y.-K. Lai, Y.-H. Liu, and C. Wallraven. Use of non-photorealistic rendering and photometric stereo in making bas-reliefs from photographs. Graphical Models, 76(4):202–213, July 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S152407031400006X.

Wolter:2000:SIC

[WP00] Franz-Erich Wolter and Nicholas M. Patrikalakis. Special issue for CGI ’98. Graphical Models, 62(1):1, January 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/artid/gmod.1999.
                                        0516/production;   http:
                                        //www.idealibrary.com/
                                        links/artid/gmod.1999.
                                        0516/production/pdf.

Williams:2005:MMS

[WR05] Jason Williams and Jarek Rossignac. Mason: morphological simplification. Graphical Models, 67(4): 285–303, July 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Wang:2003:EMD

[WS03] Yang Wang and Dimitris Samaras. Estimation of multiple directional light sources for synthesis of augmented reality images. Graphical Models, 65(4): 185–205, July 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Wang:2012:EMD

[WSC^{+}12] Hui Wang, Zhixun Su, Junjie Cao, Ye Wang, and Hao Zhang. Empirical mode decomposition on surfaces. Graphical Models, 74(4):173–183, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000252.

Wan:2012:SUP

[WSCO^{+}12] Guowei Wan, Noah Snavely, Daniel Cohen-Or, Qian Zheng, Baoquan Chen, and Sikun Li. Sorting unorganized photo sets for urban reconstruction. Graphical Models, 74(1):14–28, January 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070311000658.

Wu:2013:MSG

[WSZL13] Jinliang Wu, Xiaoyong Shen, Wei Zhu, and Ligang Liu. Mesh saliency with global rarity. Graphical Models, 75(5):255–264, September 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000180.

Wen:2017:TBE

[WTS17] Rui Wen, Weiqing Tang, and Zhiyong Su. Topology based 2D engineering drawing and 3D model matching for process plant. Graphical Models, 92(??):1–15, July 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070317300504.

Wu:2002:DMD

[Wu02] Wen-Yen Wu. A dynamic method for dominant point detection. Graphical Models, 64(5):304–315, September 2002. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Wang:2012:RCS

[WWWM12] Sen Wang, Jianhuang Wu, Mingqiang Wei, and Xin Ma. Robust curve skeleton extraction for vascular structures. Graphical Models, 74(4):109–120, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S152407031200015X.

Wang:2020:SSM

[WXGZ20] Yuting Wang, Shiqing Xin, Shanshan Gao, and Yuanfeng Zhou. Skeletal saliency map computation based on projection symmetry analysis. Graphical Models, 109(??):Article 101070, May 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407032030014X.

Wang:2018:LES

[WXN18] Weiwen Wang, Yong Xu, and Edward Ng. Largeeddy simulations of pedestrianlevel ventilation for assessing a satellite-based approach to urban geometry generation. Graphical Models, 95(??):29–41, January 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300528.

Wang:2007:VNF

[WXRA07] Hongcheng Wang, Ning Xu, Ramesh Raskar, and Narendra Ahuja. Videoshop: a new framework for spatiotemporal video editing in gradient domain. Graphical Models, 69(1):57–70, January 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000439.

Wu:2016:SGS

[WXZ^{+}16] Bo Wu, Kai Xu, Yang Zhou, Yueshan Xiong, and Hui Huang. Skeletonguided 3D shape distance field metamorphosis. Graphical Models, 85(??): 37–45, May 2016. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070316300017.

Wang:2011:QMS

[WY11] Jun Wang and Zeyun Yu. Quality mesh smoothing via local surface fitting and optimum projection. Graphical Models, 73(4):127–139, July 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070311000051.

Wyvill:2003:SII

[Wyv03] Brian Wyvill. Special issue on the International Conference of Shape Modeling (SMI) 2002. Graphical Models, 65(5):259, September 2003. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Wang:2014:SWC

[WZ14] Xiangrong Wang and Jieyu Zhao. Swendsen–Wang Cuts sampling for spatially constrained Dirichlet process mixture models. Graphical Models, 76(5):496–506, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000150.

Wang:2003:LKB

[WZL^{+}03] Tian-Shu Wang, Nan-Ning Zheng, Yan Li, Ying-Qing Xu, and Heung-Yung Shum. Learning kernelbased HMMs for dynamic sequence synthesis. Graphical Models, 65(4):206–221, July 2003. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Wen:2014:MSM

[WZtW^{+}14] Yan Wen, Yan Zhang, Wen tao Wu, Mo fei Song, and Zheng xing Sun. A model synthesis method based on single building facade. Graphical Models, 76(5):363–375, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000253.

Xing:2020:DAS

[XHH^{+}20] Yan Xing, Yeyuan He, Lei He, Wenshu Zha, and Jieqing Tan. A dynamic and adaptive scheme for feature-preserving mesh denoising. Graphical Models, 110(??):Article 101065, July 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300102.

Xu:2023:RSF

[XL23] Yixin Xu and Shiguang Liu. Realistic simulation of fruit mildew diseases: Skin discoloration, fungus growth and volume shrinkage. Graphical Models, 129(??):??, October 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000243.

Xu:2019:SBR

[XLX^{+}19] Jinlan Xu, Chengnan Ling, Gang Xu, Zhongping Ji, and Timon Rabczuk. Spline bas-relief modeling from sketches by isogeometric analysis approach. Graphical Models, 103(??):Article 101025, ???? 2019. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070319300165.

Xu:2004:VHB

[XTLP04] Songhua Xu, Min Tang, Francis C. M. Lau, and Yunhe Pan. Virtual hairy brush for painterly rendering. Graphical Models, 66(5):263–302, September 2004. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Xu:2016:BBO

[XTW16] Minfeng Xu, Changhe Tu, and Wenping Wang. Building binary orientation octree for an arbitrary scattered point set. Graphical Models, 85(??):30–36, May 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070316300029.

Xumin:2010:HPU

[XWYY10] Liu Xumin, Xu Weixiang, Guan Yong, and Shang Yuanyuan. Hyperbolic polynomial uniform B-spline curves and surfaces with shape parameter. Graphical Models, 72(1):1– 6, January 2010. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000265.

Xu:2015:SSG

[XWZ^{+}15] Linlin Xu, Ruimin Wang, Juyong Zhang, Zhouwang Yang, Jiansong Deng, Falai Chen, and Ligang Liu. Survey on sparsity in geometric modeling and processing. Graphical Models, 82(??):160–180, November 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000314.

Xin:2023:JED

[XXZL23] Hanggao Xin, Chenzhong Xiang, Wenyang Zhou, and Dun Liang. Jrender: an efficient differentiable rendering library based on Jittor. Graphical Models, 130(??):??, December 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000322.

Xie:2023:HMM

[XYZ^{+}23] Wei Xie, Zhipeng Yu, Zimeng Zhao, Binghui Zuo, and Yangang Wang. HMDO: Markerless multiview hand manipulation capture with deformable objects. Graphical Models, 127(??):??, May 2023. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070323000085.

Xu:2006:PSI

[XZWB06] Dong Xu, Hongxin Zhang, Qing Wang, and Hujun Bao. Poisson shape interpolation. Graphical Models, 68(3):268–281, May 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070306000208.

Yun:2017:RMP

[YCKK17] Dong Ho Yun, Sung In Choi, Sung Han Kim, and Kwang Hee Ko. Registration of multiview point clouds for application to ship fabrication. Graphical Models, 90(??):1–12, March 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070317300036.

Yilmaz:2007:COC

[YG07] Türker Yilmaz and Ugur Güdükbay. Conservative occlusion culling for urban visualization using a slice-wise data structure. Graphical Models, 69(3–4):191–210, May/ July 2007. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070307000033.

Yang:2018:BDT

[YGL^{+}18] Jie Yang, Lin Gao, Yu-Kun Lai, Paul L. Rosin, and Shihong Xia. Biharmonic deformation transfer with automatic key point selection. Graphical Models, 98(??):1– 13, July 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070318300304.

Kwon:2011:ABF

[yKL11] Ji yong Kwon and In-Kwon Lee. An animation bilateral filter for slow-in and slowout effects. Graphical Models, 73(5):141–150, September 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070311000075.

Yoon:2008:SCN

[YL08] Jong-Chul Yoon and In-Kwon Lee. Stable and controllable noise. Graphical Models, 70(5):105–115, September 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070308000052.

Yoon:2014:VGD

[YL14] Jong-Chul Yoon and In-Kwon Lee. Visualization of graphical data in a user-specified 2D space using a weighted Isomap method. Graphical Models, 76(2):103–114, March 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000034.

Yang:2020:RCS

[YLG^{+}20] Shanwen Yang, Tianrui Li, Xun Gong, Bo Peng, and Jie Hu. A review on crowd simulation and modeling. Graphical Models, 111(??): Article 101081, September 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300242.

Yu:2012:FSM

[YLL12] Wei Yu, Maoqing Li, and Xin Li. Fragmented skull modeling using heat kernels. Graphical Models, 74(4):140–151, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000185.

Yu:2014:GEO

[YLW^{+}14] Cheng-Chi Yu, Yong-Jin Liu, Matt Tianfu Wu, Kai-Yun Li, and Xiaolan Fu. A global energy optimization framework for 2.1D sketch extraction from monocular images. Graphical Models, 76(5):507–521, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000228.

Yuan:2019:DDW

[YLW^{+}19] Yu-Jie Yuan, Yu-Kun Lai, Tong Wu, Shihong Xia, and Lin Gao. Datadriven weight optimization for real-time mesh deformation. Graphical Models, 104(??):Article 101037, ???? 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300281.

Yang:2018:IPR

[YRZ18] Yi-Jun Yang, Muhammad Razib, and Wei Zeng. Intrinsic parameterization and registration of graph constrained surfaces. Graphical Models, 97(??): 30–39, May 2018. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300079.

Yang:2021:LFR

[YWH^{+}21] Li Yang, Jing Wu, Jing Huo, Yu-Kun Lai, and Yang Gao. Learning 3D face reconstruction from a single sketch. Graphical Models, 115(??):??, May 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000072.

Yang:2019:RVC

[YWL^{+}19] Xin Yang, Yuliang Wang, Shuai Li, Xinglin Piao, Baocai Yin, Qiang Zhang, Dongsheng Zhou, and Xiaopeng Wei. Real–virtual consistent traffic flow interaction. Graphical Models, 106(??):Article 101048, November 2019. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300396.

Yang:2017:EIT

[YWZB17] Yinhui Yang, Rui Wang, Hongxin Zhang, and Hujun Bao. ExploreTree: Interactive tree modeling in semantic trait space with online intent learning. Graphical Models, 91(??):39–51, May 2017. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070317300103.

Yang:2014:MSG

[YXF14] Long Yang, Chunxia Xiao, and Jun Fang. Multiscale geometric detail enhancement for time-varying surfaces. Graphical Models, 76(5):413–425, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070314000174.

Yang:2000:CHM

[YXYW00] Xue Dong Yang, Zhan Xu, Jun Yang, and Tao Wang. The cluster hair model. Graphical Models, 62(2):85–103, March 2000. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/
                                        doi/10.1006/gmod.1999.
                                        0518;   http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.1999.0518/pdf;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.1999.0518/ref.

Yang:2020:QRM

[YZ20] Yi-Jun Yang and Wei Zeng. Quasiconformal rectilinear map. Graphical Models, 107 (??):Article 101057, January 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070319300475.

Yang:2014:EPN

[YZC14] Yi-Jun Yang, Wei Zeng, and Jian-Feng Chen. Equiareal parameterizations of NURBS surfaces. Graphical Models, 76(1):43–55, January 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031300057X.

Yang:2010:PCB

[YZZ^{+}10] Yi-Jun Yang, Wei Zeng, Hui Zhang, Jun-Hai Yong, and Jean-Claude Paul. Projection of curves on B-spline surfaces using quadratic reparameterization. Graphical Models, 72(5):47–59, September 2010. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070310000123.

Zhao:2005:AVM

[ZB05] Liwei Zhao and Norman I.

Badler. Acquiring and validating motion qualities from live limb gestures. Graphical Models, 67(1):1– 16, January 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Zhang:2019:PRG

[ZbQC^{+}19] Yu-Wei Zhang, Bei bei Qin, Yanzhao Chen, Zhongping Ji, and Caiming Zhang. Portrait relief generation from 3D object. Graphical Models, 102(??):10–18, March 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070319300025.

Zhong:2018:CMA

[ZC18] Yanjun Zhong and Falai Chen. Computing medial axis transformations of 2D point clouds. Graphical Models, 97(??):50–63, May 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070318300092.

Zordan:2006:BEM

[ZCCD06] Victor B. Zordan, Bhrigu Celly, Bill Chiu, and Paul C. DiLorenzo. Breathe easy: Model and control of human respiration for computer animation. Graphical Models, 68(2):113–132, March 2006. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070305000287.

Zhang:2015:EVM

[ZCS^{+}15] Ran Zhang, Xuejin Chen, Takaaki Shiratori, Xin Tong, and Ligang Liu. An efficient volumetric method for non-rigid registration. Graphical Models, 79(??):1– 11, May 2015. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000053.

Zhang:2022:GPG

[ZCW^{+}22] Cheng Zhang, Hao Chen, Haocheng Wan, Ping Yang, and Zizhao Wu. Graph-PBN: Graph-based parallel branch network for efficient point cloud learning. Graphical Models, 119(??): ??, January 2022. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000254.

Zhang:2011:SGS

[ZDL^{+}11] Guo-Xin Zhang, Song-Pei Du, Yu-Kun Lai, Tianyun Ni, and Shi-Min Hu. Sketch guided solid texturing. Graphical Models, 73(3):59– 73, May 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070310000226.

Zhang:2023:CAG

[ZFLL23] Fan Zhang, Qiang Fu, Yang Liu, and Xueming Li. Component-aware generative autoencoder for structure hybrid and shape completion. Graphical Models, 129(??):??, October 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070323000152.

Zeng:2012:DHK

[ZGLG12] Wei Zeng, Ren Guo, Feng Luo, and Xianfeng Gu. Discrete heat kernel determines discrete Riemannian metric. Graphical Models, 74(4):121–129, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000161.

Zhu:2012:ATR

[ZGLP12] Hua Zhu, Shuming Gao, Ming Li, and Wanbin Pan. Adaptive tetrahedral remeshing for modified solid models. Graphical Models, 74(4):76–86, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000124.

Zhang:2018:EVU

[ZGX^{+}18] Yumin Zhang, Steven Garcia, Weiwei Xu, Tianjia Shao, and Yin Yang. Efficient voxelization using projected optimal scanline. Graphical Models, 100(??):61–70, November 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407031730053X.

Zhang:2014:UDS

[ZGZ^{+}14] Min Zhang, Ren Guo, Wei Zeng, Feng Luo, Shing-Tung Yau, and Xianfeng Gu. The unified discrete surface Ricci flow. Graphical Models, 76(5):321–339, September 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000344.

Zhang:2011:RME

[ZHM11] Shaoting Zhang, Junzhou Huang, and Dimitris N. Metaxas. Robust mesh editing using Laplacian coordinates. Graphical Models, 73(1):10–19, January 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070310000160.

Zou:2022:OAO

[ZHMW22] Zi-Xin Zou, Shi-Sheng Huang, Tai-Jiang Mu, and Yu-Ping Wang. Object-Fusion: Accurate objectlevel SLAM with neural object priors. Graphical Models, 123(??):??, September 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070322000418.

Zhou:2001:ESE

[ZK01] Lin Zhou and Chandra Kambhamettu. Extending superquadrics with exponent functions: Modeling and reconstruction. Graphical Models, 63(1):1–20, January 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/
                                        doi/10.1006/gmod.2000.
                                        0529;   http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.2000.0529/pdf;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        gmod.2000.0529/ref.

Zhang:2005:ECU

[ZK05] Jiwen Zhang and Frank L. Krause. Extending cubic uniform B-splines by unified trigonometric and hyperbolic basis. Graphical Models, 67(2):20, March 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Zhang:2008:ETS

[ZK08] Xinyu Zhang and Young J.

Kim. Efficient texture synthesis using strict Wang Tiles. Graphical Models, 70(3):43–56, May 2008. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070307000276.

Zube:2015:RDC

[ZK15] Severinas Zube and Rimvydas Krasauskas. Representation of Dupin cyclides using quaternions. Graphical Models, 82(??):110–122, November 2015. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070315000272.

Zhang:2014:GBO

[ZL14] Kang Zhang and Xin Li.

A graph-based optimization algorithm for fragmented image reassembly. Graphical Models, 76(5):484–495, September 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070314000083.

Zhang:2015:FAN

[ZL15] Xinyu Zhang and Yao Liu. A fast algebraic nonpenetration filter for continuous collision detection. Graphical Models, 80(??): 31–40, July 2015. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S152407031500020X.

Zimmer:2014:ZSA

[ZLAK14] Henrik Zimmer, Florent Lafarge, Pierre Alliez, and Leif Kobbelt. Zometool shape approximation. Graphical Models, 76(5):390–401, September 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070314000162.

Zhang:2013:ESG

[ZLH13] Guo-Xin Zhang, Yu-Kun Lai, and Shi-Min Hu. Efficient synthesis of gradient solid textures. Graphical Models, 75(3):104–117, May 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000744.

Zhou:2019:GTL

[ZLRL19] Yong Zhou, Han Lu, Qingrong Ren, and Yang Li. Generation of a tree-like support structure for fused deposition modelling based on the L-system and an octree. Graphical Models, 101(??):8–16, January 2019. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070318300468.

Zhang:2020:ATS

[ZLS^{+}20] Jianda Zhang, Chunpeng Li, Qiang Song, Lin Gao, and Yu-Kun Lai. Automatic 3D tooth segmentation using convolutional neural networks in harmonic parameter space. Graphical Models, 109(??):Article 101071, May 2020. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300151.

Zhang:2014:ARV

[ZLW^{+}14] Minqi Zhang, Fang Li, Xingce Wang, Zhongke Wu, Shi-Qing Xin, Lok-Ming Lui, Lin Shi, Defeng Wang, and Ying He. Automatic registration of vestibular systems with exact landmark correspondence. Graphical Models, 76(5):532–541, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070314000368.

Zhaohui:2018:DAS

[ZLX^{+}18] Wu Zhaohui, Wang Longge, Wu Xiaobo, Sun Hong, Li Ying, and Shi Ke. Data acquisition and simulation of dynamic flame with temperature distribution. Graphical Models, 98(??): 24–32, July 2018. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070318300316.

Zhang:2022:AMS

[ZLZ^{+}22] Jian Zhang, Chen Li, Peichi Zhou, Changbo Wang, Gaoqi He, and Hong Qin. Authoring multistyle terrain with globalto-local control. Graphical Models, 119(??):??, January 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070321000278.

Zhang:2020:CCF

[ZML^{+}20] Zili Zhang, Yue Ma, Yunfei Li, Frederick W. B. Li, Hubert P. H. Shum, Bailin Yang, Jing Guo, and Xiaohui Liang. Cumuliform cloud formation control using parameter-predicting convolutional neural network. Graphical Models, 111(??):Article 101083, September 2020. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300254.

Zeng:2014:RBB

[ZMW^{+}14] Qiong Zeng, Ralph R. Martin, Lu Wang, Jonathan A. Quinn, Yuhong Sun, and Changhe Tu. Region-based bas-relief generation from a single image. Graphical Models, 76(3):140–151, May 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000283.

Zhou:2013:CRU

[ZN13] Qian-Yi Zhou and Ulrich Neumann. Complete residential urban area reconstruction from dense aerial LiDAR point clouds. Graphical Models, 75(3):118–125, May 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000689.

Zhang:2022:SDL

[ZNYL22] Zhaochen Zhang, Jianhui Nie, Mengjuan Yu, and Xiao Liu. SharpNet: a deep learning method for normal vector estimation of point cloud with sharp features. Graphical Models, 124(??):??, November 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070322000431.

Zhang:2018:BSD

[ZPG18] Yang Zhang, Visit Pataranutaporn, and Ron Goldman. de Boor-suitable (DS) T -splines. Graphical Models, 97(??):40–49, May 2018. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070318300080.

Zhou:2011:BMP

[ZQ11] Pei Zhou and Wen-Han Qian. Blending multiple parametric normal ringed surfaces using implicit functional splines and auxiliary spheres. Graphical Models, 73(4):87–96, July 2011. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070310000457.

Zhou:2021:AGS

[ZRZ21] Hongyang Zhou, Zhong Ren, and Kun Zhou. Adaptive geometric sound propagation based on Aweighting variance measure. Graphical Models, 116 (??):??, July 2021. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S152407032100014X.

Zhao:2009:AGC

[ZS09] Liming Zhao and Alla Safonova. Achieving good connectivity in motion graphs. Graphical Models, 71(4):139–152, July 2009. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000174.

Zhu:2014:CCS

[ZSC^{+}14] Yanshu Zhu, Feng Sun, Yi-King Choi, Bert Jüttler, and Wenping Wang. Computing a compact spline representation of the medial axis transform of a 2D shape. Graphical Models, 76(5):252–262, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070314000149.

Zhong:2014:ASM

[ZSJG14] Zichun Zhong, Liang Shuai, Miao Jin, and Xiaohu Guo. Anisotropic surface meshing with conformal embedding. Graphical Models, 76(5):468–483, September 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070314000186.

Zhou:2022:PCD

[ZSL^{+}22] Lang Zhou, Guoxing Sun, Yong Li, Weiqing Li, and Zhiyong Su. Point cloud denoising review: from classical to deep learningbased approaches. Graphical Models, 121(??):??, May 2022. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070322000170.

Zhu:2020:BCS

[ZSZ^{+}20] Xiaoqiang Zhu, Chenze Song, Mengyao Zhu, Xiangyang Wang, Lihua You, and Xiaogang Jin. Barycentric convolution surfaces based on general planar polygon skeletons. Graphical Models, 109(??):Article 101069, May 2020. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070320300138.

Zheng:2003:PBC

[ZW03] Jianmin Zheng and Guozhao Wang. Perturbing Bézier coefficients for best constrained degree reduction in the L_{2}-norm. Graphical Models, 65(6):351–368, November 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Zhang:2014:PBM

[ZWSH14] Minying Zhang, Wencheng Wang, Hanqiu Sun, and Honglei Han. Perceptionbased model simplification for motion blur rendering. Graphical Models, 76(3):116–127, May 2014. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070313000301.

Zhang:2012:ERL

[ZXY^{+}12] Long Zhang, Jiazhi Xia, Xiang Ying, Ying He, Wolfgang Mueller-Wittig, and Hock-Soon Seah. Efficient and robust 3D line drawings using differenceof-Gaussian. Graphical Models, 74(4):87–98, July 2012. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070312000136.

Zhang:2021:GBL

[ZXZ21] Shao-Kui Zhang, Wei-Yu Xie, and Song-Hai Zhang. Geometry-based layout generation with hyper-relations AMONG objects. Graphical Models, 116(??):??, July 2021. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1524070321000096.

Zhang:2023:MMS

[ZYL^{+}23] Yu-Wei Zhang, Hongguang Yang, Ping Luo, Zhi Li, Hui Liu, Zhongping Ji, and Caiming Zhang. Modeling multi-style portrait relief from a single photograph. Graphical Models, 130(??):??, December 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070323000401.

Zheng:2009:SB

[ZYP09] Wen Zheng, Jun-Hai Yong, and Jean-Claude Paul. Simulation of bubbles. Graphical Models, 71(6):229–239, November 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070309000253.

Zhu:2014:SIR

[ZZ14] Chun-Gang Zhu and Xuan-Yi Zhao. Self-intersections of rational Bézier curves. Graphical Models, 76(5): 312–320, September 2014. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1524070314000277.

Zheng:2021:VSM

[ZZC^{+}21] Xinyu Zheng, Chen Zong, Jingliang Cheng, Jian Xu, Shiqing Xin, Changhe Tu, Shuangmin Chen, and Wenping Wang. Visually smooth multi-UAV formation transformation. Graphical Models, 116(??): ??, July 2021. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070321000163.

Zhang:2023:VPE

[ZZD^{+}23] Xiangjun Zhang, Yinglin Zheng, Wenjin Deng, Qifeng Dai, Yuxin Lin, Wangzheng Shi, and Ming Zeng. Vertex position estimation with spatial-temporal transformer for 3D human reconstruction. Graphical Models, 130(??):??, December 2023. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070323000371.

Zhang:2013:LBS

[ZZLZ13] Yu-Wei Zhang, Yi-Qi Zhou, Xue-Lin Li, and Li-Li Zhang. Line-based sunken relief generation from a 3D mesh. Graphical Models, 75(6):297–304, November 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S1524070313000222.

Zhang:2019:NSC

[ZZS19] Baoxing Zhang, Hongchan Zheng, and Weijie Song. A non-stationary Catmull– Clark subdivision scheme with shape control. Graphical Models, 106(??):Article 101046, November 2019. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1524070319300372.

Zeng:2013:OBF

[ZZZL13] Ming Zeng, Fukai Zhao, Jiaxiang Zheng, and Xinguo Liu. Octree-based fusion for realtime 3D reconstruction. Graphical Models, 75(3):126–136, May 2013. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000768.

Zhang:2013:RTB

[ZZZY13] Yu-Wei Zhang, Yi-Qi Zhou, Xiao-Feng Zhao, and Gang Yu. Real-time bas-relief generation from a 3D mesh. Graphical Models, 75(1):2– 9, January 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1524070312000719.